### v6

# Forecasting Response to Treatment with Global Deep Learning and Patient-Specific Pharmacokinetic Priors

Abstract
Forecasting healthcare time series is crucial for early detection of adverse outcomes and for patient monitoring. Forecasting, however, can be difficult in practice due to noisy and intermittent data. The challenges are often exacerbated by change points induced via extrinsic factors, such as the administration of medication. To address these challenges, we propose a novel hybrid global-local architecture and a pharmacokinetic encoder that informs deep learning models of patient-specific treatment effects. We showcase the efficacy of our approach in achieving significant accuracy gains for a blood glucose forecasting task using both realistically simulated and real-world data. Our global-local architecture improves over patient-specific models by 9.2-14.6%. Additionally, our pharmacokinetic encoder improves over alternative encoding techniques by 4.4% on simulated data and 2.1% on real-world data. The proposed approach can have multiple beneficial applications in clinical practice, such as issuing early warnings about unexpected treatment responses, or helping to characterize patient-specific treatment effects in terms of drug absorption and elimination characteristics.
keywords:
Multi-Input Forecasting, Glucose, Pharmacokinetics, Deep Learning
1 Introduction
Forecasting healthcare time series is crucial for early detection of adverse outcomes and patient monitoring (Churpek et al., 2016; Gerry et al., 2020). Notably, healthcare time series data are often affected by exogenous factors, such as medication, exercise, or diet. Accurate forecasting systems should strive to harness the predictive potential of these factors. While the inclusion of exogenous variables can significantly enhance prediction accuracy (Olivares et al., 2022a), there are instances in which their integration yields minimal to no improvement (McShinsky and Marshall, 2020; Rubin-Falcone et al., 2020).
Figure 1: Insulin dose is usually known and recorded in data, whereas plasma insulin concentration is unobservable and subject to changes over time.
Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where patient signals and exogenous factors, such as medication dose, often exhibit a temporal resolution mismatch. Furthermore, while the medication dose is usually known, we cannot directly observe the drug absorption and elimination parameters or plasma concentration of the drug, as illustrated in Fig. 1. Consequently, their impact on the predictive signal remains hidden.
We consider the blood glucose forecasting task while taking into account hidden medication effects. In our data, blood glucose measurements are taken every five minutes, and exogenous variables such as carbohydrate ingestion or application of insulin are recorded as instantaneous points in time. To model hidden temporal effects of such exogenous events, we propose to modify contemporary deep learning forecasting models by adding a pharmacokinetic encoder. The encoder leverages pharamacokinetic knowledge to generate patient-specific plasma insulin concentration profiles. In our framework, global parameters of a forecasting architecture are shared across multiple patients (Semenoglou et al., 2021), but the parameters of the encoder are learned from the individual patent’s historical data to address inter-subject variability in response to treatment.
The main contributions of this paper are:
  • Pharmacokinetic Encoder. We introduce a novel architecture module that generates plasma drug concentration profiles to capture time-dependent medication treatment effects.
  • Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific pharmacokinetic encoder.
  • State-of-the-Art Results. We obtain significant accuracy improvements on a large scale simulated and real-world blood glucose forecasting datasets.
1.1 Glucose Forecast Background
Type I diabetes (T1DM) patients must regularly monitor and regulate their blood glucose values to prevent hypoglycemic and hyperglycemic events, which can drastically impact organs or even be fatal. Insulin dosages, T1DM’s most crucial treatment, are represented as instantaneous events in data, which can hide their long-lasting effects. Effectively capturing the dynamic effects of medication over time is crucial to accurately predict the impact of insulin on blood glucose.
Pharmacokinetic modeling is a natural solution to the aforementioned problem, as it can unveil medication effects through knowledge of drug movement through the body. For example, pharmacokinetics provides a well-established understanding of plasma insulin concentration changes over time based on drug absorption and elimination characteristics  (Binder et al., 1984).
For extravascular injections, the plasma concentration-time profile exhibits right-skewed characteristics (Rosenbaum, 2017). Fig. 2 depicts pharmacokinetic modeling of plasma insulin concentration over time for different bolus and basal insulin types. Generally, bolus doses are quick-acting and, thus, often taken before meals compared to basal insulin, which is longer-acting and serves to regulate blood glucose levels throughout the day.
In Fig. 2, the area under the concentration curve measures drug bioavailability, the fraction of the absorbed dose that reaches the intended site of systemic circulation intact. Bioavailability can vary based on many factors, including type of medication, dosage, and patient-specific characteristics. For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent (Søeborg et al., 2012). However, for subcutaneous injections, bioavailability is lower due to unintended drug absorption or elimination.
Figure 2: Pharmacokinetic modeling of plasma insulin concentration over time.
1.2 Previous Work
Neural Forecasting. In recent years, neural networks have demonstrated significant improvements compared to traditional forecasting algorithms, following the adoption of the cross-learning technique (Smyl, 2019; Semenoglou et al., 2021), which involves utilizing expressive global models to leverage data from vast collections of related time series.
The success of this global modeling approach has seen deep forecasting methods become ubiquitous in industrial forecasting systems (Wen et al., 2017; Lim et al., 2021), and in academia (Oreshkin et al., 2020).
Blood Glucose Forecasting. Neural Forecasting literature has permeated into the field of blood glucose level prediction, outperforming the accuracy of classic statistical methods (Xie and Wang, 2020). Past research includes various neural network designs ranging from Multi Layer Perceptrons (MLP; Jahangir et al. 2017), Recurrent Neural Networks (RNN; Fox et al. 2018; Rabby et al. 2021; Rubin-Falcone et al. 2022) to Temporal Convolutional Networks (TCN; Li et al. 2019).
Despite recent advancements in blood glucose forecasting, two significant challenges remain. Firstly, existing architectures struggle to effectively utilize sparse exogenous variables with the exception of Rubin-Falcone et al. (2022). Medication and carbohydrate values are recorded as sparse values in data, which may contribute to model difficulty in learning the effects of treatment on blood glucose levels. Early solutions propose summing up exogenous values to the current point within an input window to replace sparse features as shown in Fig. 3. However, this “sum-total” preprocessing approach fails to capture time-dependent changes in medication, which can be achieved through pharmacokinetic modeling of plasma insulin concentration. Despite this, deep learning methods used to forecast blood glucose do not do not consider pharmacokinetics to inform model training. Secondly, many studies (Xie and Wang, 2020; Rubin-Falcone et al., 2022; Zaidi et al., 2021) rely on individualized networks for each patient, with limited adoption of cross-learning approaches that could enhance forecasting models by leveraging shared knowledge across patients. Our proposed approach address both challenges in a novel hybrid global-local modeling approach that learns patient-specific pharmacokinetics to improve deep learning model capabilities in forecasting blood glucose levels.
2 Cohort
Simulated data
An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes. Clinical parameters include information on age and weight, and biological parameters include information on insulin-glucose kinetics, such as absorption constants. We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set. In a similar approach to (Rubin-Falcone et al., 2022), the meal schedule used to generate simulated data was based on the Harrison-Benedict equation (Arthur Harris and Gano Benedict, 21919) with approximately 3 meals per day and no additional snacks. Height data was not provided in the patient parameters, and so was estimated at 140cm, 170cm, and 175cm for adults, adolescents, and children, respectively.
The ‘Dexcom’ continuous glucose monitor (CGM) option was used to obtain glucose readings at 5-minute intervals. The default basal-bolus controller provided in the simulation was used to administer insulin at the time of each meal, where the bolus amount is computed based on the current glucose level, the target glucose level, the patient’s correction factor, and the patient’s carbohydrate ratio. However, we adapt the controller to deliver basal doses once per hour, consistent with the OhioT1DM dataset. The resulting simulated data consists of CGM (in mg/dL) values, insulin rate (in units per minute), and carbohydrates (in grams), referred to as CHO. We extract the bolus insulin rate from the overall insulin rate as the difference in the continuous basal insulin rate, which is consistent hourly doses. Blood glucose time series for 6 simulated patients are shown in Fig. 7 in Appendix A.1.
Figure 3: Instantaneous medication administrations are represented as sparse variables in time series data. We propose a pharmacokinetic encoder to effectively capture time-dependent plasma drug concentration.
OhioT1DM
12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020). Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes. The dataset consists of both female (n=5) and male (n=8) subjects with unspecified ages within the ranges of 20-40, 40-60, and 60-80. Information on insulin pump model and type of insulin is also provided.
Sparse exogenous features in the data include basal rate (in units per hour), temporary basal rate (units per hour), and bolus insulin (units), and CHO values. Bolus insulin consists of both ‘normal’ insulin, which is recorded as an instantaneous administration, and ‘square dual’, where the insulin dose is stretched over a defined period. ‘Square dual‘ insulin doses were divided evenly across the specified administration window with the assumption that insulin units are administered consistently across specified timestamps. If multiple bolus administrations occurred within the same minute, the insulin units were summed. Basal insulin was included as hourly doses. For timestamps with overlapping basal rates, the last basal rate value was considered. When specified, the temporary basal rate superseded the basal rate.
We use the specified train and test partitions provided in the dataset. The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days. The sample count and percent missing data for each subject are included in Table 2 in Appendix A.1. Missing values were forward filled to prevent data leakage. Blood glucose time series plots for 6 subjects are shown in Fig. 8 in Appendix A.1.
3 Methods
We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), based on a 10 hour history of blood glucose, carbohydrate, insulin basal, and insulin bolus values.
3.1 Pharmacokinetic Encoder
We propose a novel pharamacokinetic encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes. The encoder consists of a function, $C$, that takes in sparse exogenous medication variables and generates a concentration-time profile for a specified dose, $d$, using the following equation:
$C(t,d,\textbf{k}^{(i)})=\frac{d}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
where $t$ is the time and $\textbf{k}^{(i)}$ is the absorption rate constant specific to patient, $i$. Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, $d$. Concentration profiles are generated under the assumption that all insulin administrations adhere to 100% bioavailability. As bioavailability equates to the area under the concentration curve, the scaled curve thus integrates to the specified dose, $d$.
Let $N$ be the number of subjects in the data. The absorption rate constant, $\textbf{k}_{\text{bolus}}\in[0,1]^{N\times 1}$ and $\textbf{k}_{\text{basal}}\in[0,1]^{N\times 1}$ are represented as network embedding weights that are initialized between values of 0 and 1 prior to model training. Initial $\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{bolus}}$ values are tuned as model hyperparameters. During model training, sparse basal and bolus features are passed into the pharmacokinetic encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig. 4. The encoder than replaces sparse medication features in the training window with the corresponding concentration curve. Non-medication features, such as CHO, are unaffected. For each training step, $\textbf{k}^{(i)}_{\text{bolus}}$ and $\textbf{k}^{(i)}_{\text{basal}}$ are updated for each patient, $i$, using gradient descent optimization.
For a single dose, $d$, of bolus insulin, $d_{\text{bolus}}$, or basal insulin, $d_{\text{basal}}$, administered at time, $t$, we obtain the concentration, $c$, at that time:
$c^{(i)}_{\text{bolus}}$ $=C(t,d^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)})$ (2)
$c^{(i)}_{\text{basal}}$ $=C(t,d^{(i)}_{\text{basal}},\textbf{k}_{\text{basal}}^{(i)})$ (3)
Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig. 3. The encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time, to perform vectorized operations that enable generating concentration curves for multiple doses in an input window in $\mathcal{O}(1)$ time. Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row. Concentration curves are generated for each row of w using the dose at the corresponding time index. The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps.
w $=\begin{bmatrix}0&1&2&3&\cdots&L-1\\ 0&0&1&2&\cdots&L-2\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ 0&0&0&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{L\times L}$ (8)
$\textbf{c}^{(i)}$ $=\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}^{(i)}_{t},\textbf{k}^{(i)})$ (9)
Here, $f$ is the frequency of the data, $\textbf{d}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of doses, and $\textbf{c}^{(i)}\in\mathbb{R}^{1\times L}$ is the new plasma insulin concentration feature for patient, $i$.
For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of concentration curves, c, other non-medication time series, x, and static features, s,
$\hat{\textbf{y}}^{(i)}_{t:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\textbf{c}^{(i)}_{t-L:t},\textbf{x}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (10)
In the proposed context, $\textbf{c}_{t-L:t}$ includes concentration curves for bolus insulin, $\textbf{c}_{\text{bolus}}$, and basal insulin $\textbf{c}_{\text{basal}}$, x includes CHO intake, and s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type, if applicable.
Figure 4: Hybrid Global-Local architecture. Global architectures with parameters shared across patients’ time series with local parameters from a patient-specific pharmacokinetic (Ph-K.) encoder. Sparse medication dose time series are passed into the pharmacokinetic encoder along with their respective $k_{a}$ parameters to generate concentration curves. Concentration curves are then passed into a deep learning model which learns relevant hidden states (hidden) and output parameters, $\hat{\theta}$.
3.2 Models
To assess forecasting performance across various model architectures, we trained models with six different algorithms as baselines. The implementations of these models were obtained from the Neuralforecast (Olivares et al., 2022b) and StatsForecast (Garza et al., 2022) libraries. More information on these models is provided in Appendix A.2.
Statistical Model
We employed Exponential Smoothing (ETS; Hyndman 2008) as our statistical baseline model.
Deep Learning Models
Our deep learning baselines included the following models: Multi-Layer Perceptron (MLP; Rosenblatt 1958), NBEATS/NHITS Oreshkin et al. (2020); Challu et al. (2022), Recurrent Neural Network (RNN; Elman (1990); Cho et al. (2014)), Long-Short Term Memory (LSTM; Sak et al. (2014)), Temporal Convolution Network with an MLP Decoder (TCN; Bai et al. (2018); van den Oord et al. (2016)), and Temporal Fusion Transformer (TFT; Lim et al. (2021))
In contrast to previous efforts that focused on training models for individual subjects, or local models, we employ global architectures with parameters shared across subjects Rubin-Falcone et al. (2022); Xie and Wang (2020). For each dataset, models were trained using all patient data to generate one global model that was applied across patients. We showcase the efficacy of global architectures in producing improved forecasts by training local models for comparison against our global model. Moreover, apart from reducing computational and storage costs, global models could prove more applicable in real-world scenarios since they can be utilized for new patients with limited or no available training data.
Models were trained using Adam optimizer (Kingma and Ba, 2017). Optimal hyperparameters were selected based on a validation set with cross-validation. Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected. Huber estimator was selected as a robust loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)$ =$=\begin{cases}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta\cdot(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{cases}$ (11)
Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set. The deep learning models were trained using an NVIDIA A100 Tensor Core GPU.
Figure 5: The boxplots show average MAE computed across 8 trials for global (orange) and local (blue) models for OhioT1DM subjects. Models trained on all subject data, or global models, have lower MAE, on average, than models trained on individual subjects, or local models.
3.3 Baselines
We implement three baselines against which we compare our proposed hybrid global-local architexture and pharmacokinetic encoder approach:
Univariate signal
We train models with the univariate blood glucose signal.
Univariate signal with exogenous features
We train models with the univariate blood glucose signal, static features, and exogenous features, consisting of sparse basal and bolus insulin dose values.
Univariate signal with sum-total feature engineering
We train models with the univariate blood glucose signal, static features, and exogenous features, consisting of insulin dose values. The sum total feature preprocessing approach was used to transform sparse features into non-sparse features of cumulative dose across an input window, as shown in Fig. 3.
3.4 Evaluation
Trained models were used to generate rolling window forecasts at one-step intervals. In accordance with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate the blood glucose forecasts using two metrics: mean absolute error (MAE) and root mean squared error (RMSE) computed across 8 trials of individually trained models. Mathematical formulations for these metrics are provided in Appendix A.4. We evaluate model performance across all points within the specified time horizon. Furthermore, given the critical importance of healthcare outcomes related to blood glucose levels exceeding or falling below safe thresholds, we also assess the model’s performance exclusively at time points where blood glucose levels reach or fall below certain thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL). For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation.
Our code implementation is open-source and can be found on Github. [^1] https://github.com/Nixtla/neuralforecast/tree/treat_module
4 Results
The pharmacokinetic encoder significantly improves forecasting accuracy for the top performing deep learning model by approximately 4.4% and 2.1%, for simulated and OhioT1DM datasets, respectively, as shown in Table 1. For the OhioT1DM dataset, the NHITS-Pharamacokinetic model has significantly smaller MAE than the NHITS-Sparse Exogenous model for forecasts evaluated at all values (t-test p-value: 4.44e-05) and critical values (t-test p-value: 1.6e-2). Fig. 6 demonstrates improved forecasting performance using the the NHITS-Pharamacokinetic model during hyperglycemic and hypoglycemic events for three randomly selected OhioT1DM subjects. Mean MAE and RMSE results with standard deviations computed across 8 trials are shown in Tables 7 and 8 in Appendix B.3.
We showcase the efficacy of our novel hybrid global-local model in producing more accuracy forecasts than models trained per patient. Global models consistently outperform local models, on average, for each OhioT1DM subject as shown in Table 4 and Fig. 9 in Appendix B.1. For the OhioT1DM dataset, the hybrid global-local pharmacokinetic model achieves a significantly lower (t-test p-value: 1.29e-2) MAE by 14.6% over patient-specific pharmacokinetic models. Fig. 5 shows that the global-local pharmacokinetic model has lower MAE, on average, across all OhioT1DM patients compared to the local models. Furthermore, the hybrid global-local pharmacokinetic model also achieves a significantly lower (t-test p-value: 3.83e-2) MAE by 9.2% over the patient-specific sparse exogenous models.
Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
Simulated Dataset
Model All Values Critical Values
Baseline Exogenous Baseline Exogenous
ETS 10.040 13.531
MLP 9.734 8.653 12.189 9.887
LSTM 8.731 8.372 10.863 9.996
TCN 16.772 12.234 28.913 18.122
TFT 7.788 6.762 9.664 7.828
NHITS 8.333 7.282 10.184 8.353
TFT- Sum-Total 6.698 7.634
NHITS- Pharmacokinetic 7.027 7.917
TFT- Pharmacokinetic 6.466 7.514
OhioT1DM Dataset
ETS 10.126 11.781
MLP 10.568 10.275 11.805 11.434
LSTM 10.604 9.606 12.999 11.144
TCN 11.011 11.090 12.693 13.212
TFT 9.605 9.515 10.686 10.555
NHITS 9.234 8.954 10.436 10.117
NHITS- Sum-Total 8.988 10.092
NHITS- Pharmacokinetic 8.769 10.007
TFT- Pharmacokinetic 9.492 10.484
Figure 6: Blood glucose forecasts of the final horizon point for three OhioT1DM subjects. The NHITS -Pharmacokinetic model (orange) produces noticeably improved forecasts of the ground truth (blue) compared to the NHITS model with sparse exogenous features (yellow).
5 Discussion
Our pharmacokinetic encoder enables deep learning models to achieve significantly lower forecasting errors. As such, models that only include sparse features may be under-leveraging the potential utility of exogenous information. Our pharmacokinetic encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings about unexpected treatment response, or helping to characterize patient-specific treatment effects in terms of drug absorption and elimination characteristics. The characterization of patient-specific pharmacokinetic parameters holds substantial potential for advancing the development of more effective treatments by tailoring medication to individual patients.
Assessing a forecasting model’s ability to effectively capture the impact of variables on the target signal is essential for ensuring the reliability of forecasts. We examine whether the pharmacokinetic model can accurately capture the treatment effects of insulin on blood glucose by generating forecasts during time periods for three scenarios: with the original bolus insulin doses, with bolus insulin doses removed, and with bolus insulin doses increased by a factor of 10. When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig. 10 in Appendix B.2. Our model thus effectively learns the inhibitory treatment effects of insulin on blood glucose.
Bolus insulin is quick-acting and often taken before meals compared to basal insulin, which is longer-acting. Our pharamcokinetic encoder learns $k_{a}$ values that align with this established knowledge. For OhioT1DM data, the average $k_{a_{bolus}}$ and $k_{a_{basal}}$ were 1.622 and 1.486, respectively. For our approach, larger $k_{a}$ values correspond to concentration curves with shorter duration, or steeper curves. The average $k_{a}$ values for bolus insulin were higher than those for basal insulin, on average, resulting in shorter durations for bolus insulin concentration profiles.
To evaluate whether the difference in $k_{a_{bolus}}$ and $k_{a_{basal}}$ values was statistically significant, we compute a one-sided paired t-test of model-inferred $k_{a}$ values for patients across the 8 trials. The alternative hypothesis claims that $k_{a}{{}_{bolus}}$ values are larger than $k_{a_{basal}}$. The test yielded a p-value of 0.043 indicating that the $k_{a}$ parameter for bolus insulin is significantly higher than that for basal insulin when considering a significance level of 0.05. Thus, our approach can learn the relative differences in pharamcokinetic characteristics between basal and bolus insulin.
Limitations. We equate $\sigma$ to our $k_{a}$ parameter in the log-normal function and hold $\mu$ constant at 1. This simplification reduces the number of potential concentration curve shapes the model can learn. Tuning both $\sigma$ and $\mu$ would enable more flexible concentration curves with potentially more accurate duration and peak characteristics, which could further improve forecasts.
Future Work. Future work will consider adaptations to the concentrator log-normal function to enable more flexible concentration curves. Future work will also consider expanding our pharmacokinetic encoder with functions relevant to other medication administration methods, such as intravenous injections. Finally, another interesting line of research is determining how to leverage the advantages of the global forecasting model with patient-specific $k_{a}$ values for new patients.
[…]
This work was partially supported by the NIH award R01HL141916. The authors wish to thank Dr. Frank Guyette, Dr. Leonard Weiss and Dr. Gilles Clermont for their guidance and informative conversations.
Appendix A First Appendix
A.1 Data
Fig. 7 shows the first 10 days of data for 6 subjects from the simulated dataset.
Fig. 8 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset.
Table 2 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
Table 2: Sample count (miss- ing samples) for the original OhioT1DM datasets and prepro- cessed test set
ID Original Preprocessed
Train Test Test
540 13109 3066 2691
(8.87%) (6.43%) (6.91%)
544 12671 3137 2691
(16.16%) (13.42%) (14.08%)
552 11097 3950 2691
(18.80%) (40.15%) (53.36%)
559 12081 2876 2691
(10.64%) (12.59%) (13.45%)
563 13098 2691 2691
(7.44%) (4.50%) (4.50%)
567 13536 2871 2691
(19.78%) (16.79%) (16.91%)
570 11611 2880 2691
(5.42%) (4.69%) (4.91%)
575 13103 2719 269
(9.44%) (4.74%) (4.79%)
584 13248 2995 2691
(8.29%) (11.02%) (11.93%)
588 13105 2881 2691
(3.55%) (3.12%) (3.34%)
591 12755 2847 2691
(14.96%) (3.06%) (3.23%)
596 13630 3003 2691
(20.20%) (8.66%) (9.66%)
Figure 7: Time series examples for the first 10 days of 6 subjects from the simulated dataset.
Figure 8: Time series for the first 10 days of 6 subjects from the OhioT1DM dataset.
A.2 Models
Exponential Smoothing (ETS) The ETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. (Hyndman, 2008).
Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
Neural Hierarchical Interpolation for Time Series (NHITS) - A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands. The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation (Challu et al., 2022).
Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states. The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts.(Elman, 1990; Cho et al., 2014)
Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.(Sak et al., 2014)
Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts. To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions (Bai et al., 2018; van den Oord et al., 2016).
Temporal Fusion Transformer (TFT) An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies. TFT also leverages variable selection networks as well as a series of gating layers to suppress unnecessary parts of the architecture (Lim et al., 2021).
A.3 Hyperparameters
For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 3. Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 3.
Hyperparameter Considered Values
Learning rate Loguniform(1e-4, 1e-2)
Batch size 4
Dropout {0.0, 0.1, 0.2, 0.3, 0.4, 0.5}
Training steps {1000, 2000}
Random seed DiscreteRange(1, 10)
Basal dose initial $k_{a}$ Uniform(1, 2, 11)
Bolus dose initial $k_{a}$ Uniform(1, 2, 11)
Table 3: Common hyperparameter search space
We tune additional specific hyperparameters for each architecture. For the MLP  we tune the number of hidden units per layer ([128,256]); for the LSTM  the hidden size ([128,256]); for the TCN the hidden size ([128,256]) and number of dilations ([2,4,8]); for the TFT the hidden size ([128,256]); for the NHITS we did not tune additional hyperparameters. The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
A.4 Metrics
Mean Absolute Error (MAE)
$MAE(y_{t+1,t+H},\hat{y}_{t+1,t+H})=\frac{1}{H}\sum_{i=t+1}^{t+H}|y_{i}-\hat{y}_{i}|$
Root Mean Square Error (RMSE)
$RMSE(y_{t+1,t+H},\hat{y}_{t+1,t+H})=\sqrt{\frac{1}{H}\sum_{i=t+1}^{t+H}(y_{i}-\hat{y}_{i})^{2}}$
Appendix B Second Appendix
B.1 Patient-specific (local) models
We showcase the effectiveness of global models, with parameters shared across subjects, in generating more accuracy forecasts compared to local models trained for individual subjects. Global models consistently outperform local models, on average across 8 trials, for each OhioT1DM subject as shown in Table. 4 and Fig 9b. Furthermore, the global NHITS-Pharamcokinetic model consistently outperforms the global NHITS-Exogenous model, on average, for each OhioT1DM subject as shown in Fig 9a.
Table 4: MAE computed across forecast horizon values using local and global models.
ID Exogenous Pharmacokinetic
Local Global Local Global
540 11.358 10.325 12.059 10.050
(0.236) (0.168) (0.414) (0.150)
544 7.513 7.312 7.839 7.028
(0.207) (0.150) (0.239) (0.127)
552 8.400 7.547 8.854 7.308
(0.346) (0.105) (0.282) (0.150)
559 9.888 9.093 10.486 9.000
(0.379) (0.094) (0.579) (0.065)
563 9.247 8.610 9.172 8.580
(0.102) (0.071 (0.302) (0.107)
567 11.254 9.835 11.455 9.662
(0.534) (0.142) (0.349) (0.181)
570 8.160 7.470 8.259 7.280
(0.318) (0.134) (0.511) (0.160)
575 10.500 9.734 11.149 9.539
(0.166) (0.086) (0.377) (0.136)
584 10.605 9.958 13.706 9.787
(0.332) (0.146) (6.499) (0.135)
588 8.966 8.295 9.454 8.158
(0.258) (0.041) (0.297) (0.077)
591 11.237 10.344 11.275 10.103
(0.321) (0.080) (0.102) (0.075)
596 8.797 8.230 9.471 8.008
(0.125) (0.073) (0.624) (0.136)
All 9.661 8.954 10.265 8.769
(1.272) (0.058) (1.651) (0.069)
Figure 9: The pharmacokinetic NHITS  model outperforms the NHITS baseline exogenous model across all patients, on average (a). Additionally, the pharmacokinetic model trained on data from subjects, or the global model, achieves lower MAE than models trained on individual subjects, or local models (b).
B.2 Forecasting treatment effects
In this appendix, we explore the effect of changing the insulin dosage on the forecasts produced by our approach. Fig. 10, demonstrates that the pharmacokinetic model appropriately learns the treatment effect of insulin on blood glucose for OhioT1DM subjects, decreasing the forecast when the dosage increases.
Figure 10: Treatment effect of insulin on the NHITS pharmacokinetic model. Each point in the plot corresponds to the final horizon point. Forecasts with original bolus doses (orange), removing bolus insulin doses results in higher forecasts (blue), and augmenting doses results in lower blood glucose forecasts (yellow).
B.3 Model MAE and RMSE results
This appendix presents the complete results for all models, metrics of evaluation, and both simulated data and OhioT1DM. Standard deviation across 8 trials are provided in parenthesis.
Table 5: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated data. Average results are computed across 8 trials.
Model All Values Critical Values
Baseline Exogenous Baseline Exogenous
ETS 10.040 13.531
(0.00) (0.00)
MLP 9.734 8.653 12.189 9.887
(1.125) (1.500) (1.198) (1.681)
LSTM 8.731 8.372 10.863 9.996
(0.573) (1.553) (1.555) (2.315)
TCN 16.772 12.234 28.913 18.122
(9.795) (7.866) (23.303) (19.911)
TFT 7.788 6.762 9.664 7.828
(1.599) (0.570) (1.999) (0.752)
NHITS 8.333 7.282 10.184 8.353
(0.166) (0.184) (0.303) (0.363)
TFT-Sum Total 6.698 7.634
(0.508) (0.546)
NHITS-Pharmacokinetic 7.027 7.917
(0.091) (0.268)
TFT-Pharmacokinetic 6.466 7.514
(0.570) (0.985)
Table 6: Root Mean Square error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated data.
Model All Values Critical Values
Baseline Exogenous Baseline Exogenous
ETS 15.754 23.151
(0.00) (0.00)
MLP 14.285 11.607 19.443 13.731
(0.981) (1.756) (0.737) (2.321)
LSTM 12.802 11.273 16.903 13.595
(0.759) (1.918) (2.100) (3.260)
TCN 22.874 16.194 36.468 22.533
(11.937) (10.309) (22.454) (20.991)
TFT 11.930 9.801 15.736 11.708
(1.424) (0.690) (1.676) (0.948)
NHITS 13.058 10.479 18.036 12.822
(0.255) (0.372) (0.644) (0.855)
TFT-Sum Total 9.711 11.283
(0.599) (0.748)
NHITS-Pharmacokinetic 10.024 11.931
(0.204) (0.549)
TFT-Pharmacokinetic 9.407 11.224
(0.654) (1.152)
Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data. Average results are computed across 8 trials.
Model All Values Critical Values
Baseline Exogenous Baseline Exogenous
ETS 10.126 11.781
(0.00) (0.00)
MLP 10.568 10.275 11.805 11.434
(0.628) (0.183) (0.731) (0.186)
LSTM 10.604 9.606 12.999 11.144
(0.929) (0.372) (2.294) (0.708)
TCN 11.011 11.090 12.693 13.212
(1.066) (0.708) (1.334) (0.886)
TFT 9.605 9.515 10.686 10.555
(0.735) (0.310) (0.986) (0.495)
NHITS 9.234 8.954 10.436 10.117
(0.056) (0.058) (0.125) (0.086)
NHITS-Sum Total 8.988 10.092
(0.023) (0.082)
NHITS-Pharmacokinetic 8.769 10.007
(0.069) (0.087)
TFT-Pharmacokinetic 9.492 10.484
(0.399) (0.374)
Table 8: Root mean square error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data.
Model All Values Critical Values
Baseline Exogenous Baseline Exogenous
ETS 15.351 17.478
(0.00) (0.00)
MLP 16.128 15.625 17.942 17.482
(0.580) (0.131) (0.716) (0.232)
LSTM 15.766 14.704 18.565 16.784
(0.930) (0.297) (1.818) (0.688)
TCN 16.173 16.192 18.432 18.978
(0.932) (0.545) (1.097) (0.806)
TFT 15.031 14.820 16.720 16.455
(0.576) (0.186) (0.954) (0.413)
NHITS 15.106 14.596 16.901 16.421
(0.075) (0.080) (0.168) (0.133)
NHITS-Sum Total 14.714 16.438
(0.043) (0.135)
NHITS-Pharmacokinetic 14.375 16.287
(0.126) (0.154)
TFT-Pharmacokinetic 14.846 16.351
(0.366) (0.284)

### v7

@@
-Forecasting Response to Treatment with Global Deep Learning and Patient-Specific Pharmacokinetic Priors
+Forecasting Treatment Response with Deep Pharmacokinetic Encoders
 Abstract
-Forecasting healthcare time series is crucial for early detection of adverse outcomes and for patient monitoring.
-Forecasting, however, can be difficult in practice due to noisy and intermittent data.
-The challenges are often exacerbated by change points induced via extrinsic factors, such as the administration of medication.
-To address these challenges, we propose a novel hybrid global-local architecture and a pharmacokinetic encoder that informs deep learning models of patient-specific treatment effects.
+Forecasting healthcare time series data is vital for early detection of adverse outcomes and patient monitoring.
+However, forecasting is challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties for each patient.
+To address these challenges, we propose a novel hybrid global-local architecture and a PK encoder that informs deep learning models of patient-specific treatment effects.
 We showcase the efficacy of our approach in achieving significant accuracy gains for a blood glucose forecasting task using both realistically simulated and real-world data.
-Our global-local architecture improves over patient-specific models by 9.2-14.6%.
-Additionally, our pharmacokinetic encoder improves over alternative encoding techniques by 4.4% on simulated data and 2.1% on real-world data.
-The proposed approach can have multiple beneficial applications in clinical practice, such as issuing early warnings about unexpected treatment responses, or helping to characterize patient-specific treatment effects in terms of drug absorption and elimination characteristics.
+Our hybrid global-local architecture improves over patient-specific models by 15.8%  on average.
+Additionally, our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels.
+Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
 keywords:
-Multi-Input Forecasting, Glucose, Pharmacokinetics, Deep Learning
+Multi-Input Forecasting, Healthcare, Pharmacokinetics, Deep Learning
 1 Introduction
-Forecasting healthcare time series is crucial for early detection of adverse outcomes and patient monitoring (Churpek et al., 2016; Gerry et al., 2020).
-Notably, healthcare time series data are often affected by exogenous factors, such as medication, exercise, or diet.
-Accurate forecasting systems should strive to harness the predictive potential of these factors.
-While the inclusion of exogenous variables can significantly enhance prediction accuracy (Olivares et al., 2022a), there are instances in which their integration yields minimal to no improvement (McShinsky and Marshall, 2020; Rubin-Falcone et al., 2020).
-Figure 1: Insulin dose is usually known and recorded in data, whereas plasma insulin concentration is unobservable and subject to changes over time.
-Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where patient signals and exogenous factors, such as medication dose, often exhibit a temporal resolution mismatch.
-Furthermore, while the medication dose is usually known, we cannot directly observe the drug absorption and elimination parameters or plasma concentration of the drug, as illustrated in Fig.
+Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution.
+Furthermore, while the medication dose is usually known, one cannot directly observe the plasma concentration of the drug, as illustrated in Fig.
 1.
-Consequently, their impact on the predictive signal remains hidden.
-We consider the blood glucose forecasting task while taking into account hidden medication effects.
-In our data, blood glucose measurements are taken every five minutes, and exogenous variables such as carbohydrate ingestion or application of insulin are recorded as instantaneous points in time.
-To model hidden temporal effects of such exogenous events, we propose to modify contemporary deep learning forecasting models by adding a pharmacokinetic encoder.
-The encoder leverages pharamacokinetic knowledge to generate patient-specific plasma insulin concentration profiles.
-In our framework, global parameters of a forecasting architecture are shared across multiple patients (Semenoglou et al., 2021), but the parameters of the encoder are learned from the individual patent’s historical data to address inter-subject variability in response to treatment.
+Consequently, the impact of dynamic changes in treatment resulting from drug absorption and elimination remains hidden.
+Figure 1: Insulin dose is usually known and recorded in data, whereas plasma insulin concentration is unobservable and varies over time.
+We consider the blood glucose forecasting task while taking into account latent, time-dependent treatment effects.
+Patients with diabetes must regularly monitor and regulate their blood glucose values to prevent hypoglycemic and hyperglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal.
+Hyperglycemic and hypoglycemic episodes are responsible for numerous hospital admissions each year.
+A 2009 report on hospital admissions for hyperglycemia indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009).
+Another report on hypoglycemic hospital admissions suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021).
+With over 38 million patients with Diabetes Mellitus in the United States alone, this equates to a conservative estimate of 212,800 patients hospitalized for hypoglycemic episodes each year.
+Capturing the dynamic effects of insulin is crucial for accurately predicting its impact on blood glucose levels to anticipate and proactively address abnormal levels to prevent adverse patient events.
 The main contributions of this paper are:
-  • Pharmacokinetic Encoder. We introduce a novel architecture module that generates plasma drug concentration profiles to capture time-dependent medication treatment effects.
-  • Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific pharmacokinetic encoder.
-  • State-of-the-Art Results. We obtain significant accuracy improvements on a large scale simulated and real-world blood glucose forecasting datasets.
-1.1 Glucose Forecast Background
-Type I diabetes (T1DM) patients must regularly monitor and regulate their blood glucose values to prevent hypoglycemic and hyperglycemic events, which can drastically impact organs or even be fatal.
-Insulin dosages, T1DM’s most crucial treatment, are represented as instantaneous events in data, which can hide their long-lasting effects.
-Effectively capturing the dynamic effects of medication over time is crucial to accurately predict the impact of insulin on blood glucose.
-Pharmacokinetic modeling is a natural solution to the aforementioned problem, as it can unveil medication effects through knowledge of drug movement through the body.
-For example, pharmacokinetics provides a well-established understanding of plasma insulin concentration changes over time based on drug absorption and elimination characteristics  (Binder et al., 1984).
-For extravascular injections, the plasma concentration-time profile exhibits right-skewed characteristics (Rosenbaum, 2017).
-Fig.
-2 depicts pharmacokinetic modeling of plasma insulin concentration over time for different bolus and basal insulin types.
-Generally, bolus doses are quick-acting and, thus, often taken before meals compared to basal insulin, which is longer-acting and serves to regulate blood glucose levels throughout the day.
-In Fig.
-2, the area under the concentration curve measures drug bioavailability, the fraction of the absorbed dose that reaches the intended site of systemic circulation intact.
-Bioavailability can vary based on many factors, including type of medication, dosage, and patient-specific characteristics.
-For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent (Søeborg et al., 2012).
-However, for subcutaneous injections, bioavailability is lower due to unintended drug absorption or elimination.
-Figure 2: Pharmacokinetic modeling of plasma insulin concentration over time.
-1.2 Previous Work
-Neural Forecasting.
-In recent years, neural networks have demonstrated significant improvements compared to traditional forecasting algorithms, following the adoption of the cross-learning technique (Smyl, 2019; Semenoglou et al., 2021), which involves utilizing expressive global models to leverage data from vast collections of related time series.
-The success of this global modeling approach has seen deep forecasting methods become ubiquitous in industrial forecasting systems (Wen et al., 2017; Lim et al., 2021), and in academia (Oreshkin et al., 2020).
+  • Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific encoder.
+  • Pharmacokinetic (PK) Encoder. We introduce a novel deep learning model-agnostic module that generates plasma concentration profiles to capture time-dependent treatment effects.
+  • Multiple Dose Effects Framework.
+We provide a computationally efficient framework for encoding the effects of multiple doses.
+We verify this approach mathematically under the assumption of linear PK.
+  • State-of-the-Art Results. We obtain significant accuracy improvements on large scale simulated and real-world patient blood glucose forecasting datasets.
+1.1 Pharmacokinetics Background
+PK modeling is a natural solution to the aforementioned problem, as it can unveil medication effects through knowledge of drug movement within the body.
+For example, PK provides a well-established understanding of plasma insulin concentration changes over time based on drug absorption and elimination characteristics  (Binder et al., 1984).
+Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves.
+The degree of skew varies with insulin type and patient-specific factors, as shown in Fig.
+2, where the area under the concentration curve measures drug bioavailability, or the fraction of the absorbed dose that reaches the intended site of systemic circulation intact.
+Most drugs used in clinical practice are assumed to follow linear PK (Rosenbaum, 2017).
+Under this model, the area under the concentration curve is directly proportional to the drug dose.
+A complete reference for supporting PK information regarding concentration-time profiles, insulin analogs, bioavailability, first-order kinetics, and linear PK can be found in Appendix B.
+Figure 2: Pharmacokinetic modeling of plasma insulin concentration over time for various insulin types.
+2 Related Work
 Blood Glucose Forecasting.
-Neural Forecasting literature has permeated into the field of blood glucose level prediction, outperforming the accuracy of classic statistical methods (Xie and Wang, 2020).
-Past research includes various neural network designs ranging from Multi Layer Perceptrons (MLP; Jahangir et al.
-2017), Recurrent Neural Networks (RNN; Fox et al.
-2018; Rabby et al.
-2021; Rubin-Falcone et al.
-2022) to Temporal Convolutional Networks (TCN; Li et al.
-2019).
-Despite recent advancements in blood glucose forecasting, two significant challenges remain.
-Firstly, existing architectures struggle to effectively utilize sparse exogenous variables with the exception of Rubin-Falcone et al.
-(2022).
-Medication and carbohydrate values are recorded as sparse values in data, which may contribute to model difficulty in learning the effects of treatment on blood glucose levels.
-Early solutions propose summing up exogenous values to the current point within an input window to replace sparse features as shown in Fig.
+Neural Forecasting has previously been applied to the field of blood glucose forecasting, outperforming the accuracy of classic statistical methods.
+RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020).
+Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020; Jahangir et al., 2017) have been utilized in proposed models and as baselines.
+Blood Glucose Treatment Effect Modeling.
+Previous studies have successfully modeled treatment effects on patient blood glucose levels.
+Miller et al.
+present one of the first hybrid statistical-and-physiological model that combines glucose-insulin dynamics and deep learning (Miller et al., 2020).
+They present the Deep T1D Simulator (DTD-Sim) model, which uses a deep state-space model to describe parameters representing insulin and carbohydrate absorption rates and sensitivities of the UVA/Padova simulator that simulates blood glucose levels.
+Odnoblyudova et al.
+demonstrated enhanced accuracy in predicting blood glucose levels using Multi-output Gaussian Processes (MOGPs) to model composite effects of carbohydrate and fat components from meals.
+(Odnoblyudova et al., 2023).
+Despite recent advancements in forecasting treatment effects, significant challenges remain, including:
+Sparse data.
+Limited treatment data can hinder deep learning models from accurately learning treatment responses (Rubin-Falcone et al., 2022; Hussain et al., 2021).
+To address this challenge, we integrate domain knowledge into the deep learning models, allowing them to more effectively encode the time-dependent effects of these variables and improve forecasting performance.
+Composite treatment effects.
+Unlike meals that consist of various components such as carbohydrates and fats, insulin is a homogenous protein hormone composed of a single type of molecule.
+Therefore, modeling its component effects is not applicable (Odnoblyudova et al., 2023).
+Instead, we aim to model patient-specific holistic effects of the drug as well as composite effects of multiple doses.
+Inflexible models.
+Previous glucose forecasting models, like the UVA/Padova simulator, are specific to blood glucose dynamics and inflexible to other forecasting targets and variables (Visentin et al., 2016; Miller et al., 2020).
+Our approach leverages a deep learning model to generate predictions.
+This distinction provides greater flexibility, as our forecasting model is not constrained by a specific physiological model and can accommodate various deep learning models and forecasting targets and inputs.
+Model application scope.
+Prior work that integrates PK domain knowledge in deep learning models has been used in the context of predicting disease progression for cancer and COVID-19 with long-term horizons of interest (Hussain et al., 2021; Qian et al., 2021).
+In contrast, we leverage linear PK assumptions to create a framework that can be directly applied to various medications used in clinical practice and is easily adapted to other treatments.
+Individualized networks.
+Studies often use individualized networks, limiting cross-learning. Our novel global-local architecture allows efficient training of a more accurate global model applicable to multiple patients.
+3 Methods
+We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10 hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values.
+We choose a 10-hour duration to balance sufficient historical information with GPU memory constraints.
+the duration of the rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (ins, 2018, ).
+Therefore, the selection of a 10-hour lag time can fully cover multiple dose durations, allowing the model to learn treatment effects across various dosages and durations.
+3.1 PK Encoder
+Medication administrations are represented as instantaneous, sparse variables in time series data.
+We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes as shown in Fig.
 3.
-However, this “sum-total” preprocessing approach fails to capture time-dependent changes in medication, which can be achieved through pharmacokinetic modeling of plasma insulin concentration.
-Despite this, deep learning methods used to forecast blood glucose do not do not consider pharmacokinetics to inform model training.
-Secondly, many studies (Xie and Wang, 2020; Rubin-Falcone et al., 2022; Zaidi et al., 2021) rely on individualized networks for each patient, with limited adoption of cross-learning approaches that could enhance forecasting models by leveraging shared knowledge across patients.
-Our proposed approach address both challenges in a novel hybrid global-local modeling approach that learns patient-specific pharmacokinetics to improve deep learning model capabilities in forecasting blood glucose levels.
+The encoder consists of a function, $C$, that takes in sparse exogenous variables and generates a concentration-time profile for a specified dose, $\textbf{x}^{(i)}$, using the following equation:
+$C(t,\textbf{x}^{(i)},\textbf{k}^{(i)})=\frac{\textbf{x}^{(i)}}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
+where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate that specific to patient, $i$.
+Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, x.
+The intuition for the log-normal shape of the concentration curves stems from pharmacokinetics literature as outlined in Section 1.1 and Appendix B.
+Figure 3: Medication administrations are represented as sparse variables in time series data.
+We propose a pharmacokinetic (PK) encoder to effectively capture time-dependent plasma drug concentration (Conc.).
+Under linear PK assumptions, PK parameters, including bioavailability, are assumed to be constant with respect to the drug.
+Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.5.
+To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose, x, as a scaling factor in $C$.
+This ensures that the area under the curve integrates to the dose value, under the simplifying assumption that all insulin administrations adhere to 100% bioavailability, which we show is upper bounded by the dose.
+Scaling the concentration curve changes its magnitude but does not alter the PK parameters or shape of the curve.
+Let $N$ be the number of subjects in the data.
+The PK parameter $\textbf{k}\in[0,1]^{N\times 1}$ that controls the shape of the concentration curve, is represented as network embedding weight that is initialized between values of 1 and 2 prior to model training.
+The initial k values are tuned as model hyperparameters.
+The encoder is trained end-to-end with the model.
+During model training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig.
+4.
+The encoder than replaces sparse medication features in the training window with the corresponding concentration curve.
+For a single dose, x, of insulin administered at time, $t$, we obtain the concentration, $\widecheck{\textbf{x}}^{(i)}$, at that time for patient, $i$:
+$\widecheck{\textbf{x}}^{(i)}$ $=C(t,\textbf{x}^{(i)},\textbf{k}^{(i)}).$ (2)
+For each training step, $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
+For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded in the time prior to $t$ based on the lag time, $L$,
+$\hat{\textbf{y}}^{(i)}_{t:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
+In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type.
+$\widecheck{\textbf{x}}_{t-L:t}$ includes concentration curves for basal insulin, $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin, $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO, $\widecheck{\textbf{x}}_{\text{CHO}}$.
+We include plasma treatment effects of CHO because, similar to drugs, it undergoes absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
+The PK encoder training framework is outlined in Appendix A.6, which can be integrated into any deep learning architecture capable of handling exogenous variables alongside modeling the target variable.
+3.2 Modeling Multiple Dose Effects
+Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig.
+11.
+To model PK effects for multiple doses in an input window, we generate a concentration curve for each dose event and sum concentration curves across the windows.
+Under the assumption of linear pharmacokinetics (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.5.
+To perform vectorized operations that enable generating concentration curves for multiple doses in an input window, the PK encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time.
+This approach allows for efficient modeling of multiple effects in $\mathcal{O}(1)$ time.
+Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row.
+Concentration curves are generated for each row of w using the dose at the corresponding time index.
+If no dose is administered at a particular time, $t$, the corresponding row of w indexed by $t$ will be zeroed out.
+The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps:
+w $=\begin{bmatrix}0&1&2&3&\cdots&L-1\\ 0&0&1&2&\cdots&L-2\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ 0&0&0&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{L\times L}$ (4)
+$\widecheck{\textbf{x}}^{(i)}$ $=\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}^{(i)}_{t},\textbf{k}^{(i)})$ (5)
+Here, $f$ is the frequency of the data, $\textbf{x}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of treatment, and $\widecheck{\textbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$ is the new concentration curve feature for patient, $i$.
+Figure 4: Our hybrid global-local architecture combines global model parameters shared across patients with patient-specific pharmacokinetic (PK) parameters.
+Sparse treatment time series features, such as medication doses, are input into the PK encoder to generate patient-specific treatment effect curves, $\check{\textbf{x}}$.
+PK encoder output, $\check{\textbf{x}}$, and any other exogenous features, x, are used as deep learning model inputs.
+During each training step, the model learns relevant hidden states and output parameters, $\hat{\theta}$, across patients, while $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
+3.3 Modeling Dose-Dependent Effects
+A majority of drugs are assumed to adhere to linear PK, where it is assumed that there is a linear relationship between drug dose and concentration in the body (Rosenbaum, 2017).
+A characteristic of this assumption is that PK parameters are constant and do not change with dose.
+In practice, a drug’s PK parameters can vary depending on factors such as the drug’s properties, the type of injection, and the injection site.
+Nonlinear PK occurs when the relationship between drug dose and body concentration is not linear, leading to changes in PK parameters with dose.
+For instance, subcutaneous injections are administered into the subcutaneous tissue, where higher doses can saturate elimination pathways, causing slower absorption rates (Søeborg et al., 2009; Gradel et al., 2018).
+This variability in PK parameters can result in dose-dependent plasma drug concentration profiles.
+CGMs administer the same type of insulin (rapid-acting) for basal and bolus doses.
+However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs.
+9 and 10 in Appendix A.1.
+To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity.
+By treating each dose range as a separate entity, we can capture potential PK variations without resorting to a fully nonlinear PK model.
+This method simplifies the modeling process by maintaining linear assumptions within each dose category.
+It avoids the complexity of a full nonlinear model while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics.
+Specifically, we estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values.
+For a single dose of basal insulin, $\textbf{x}_{\text{basal}}$, or bolus insulin, $\textbf{x}_{\text{bolus}}$, administered at time, $t$, we obtain the plasma insulin concentration at that time for patient, $i$:
+$\widecheck{\textbf{x}}^{(i)}_{\text{basal}}$ $=C(t,\textbf{x}^{(i)}_{\text{basal}},\textbf{k}_{\text{basal}}^{(i)}),$ (6)
+$\widecheck{\textbf{x}}^{(i)}_{\text{bolus}}$ $=C(t,\textbf{x}^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)}).$ (7)
+Treating each dose as a different entity is an approximation and may not capture all nuances of dose-dependent kinetics.
+All physiological models are approximations of complex biological processes.
+Modeling dose-dependent effects in a more granular manner could potentially provide more precise results if such effects are present.
+However, our aim is to develop minimal models validated under established linear PK assumptions from the pharmacology literature, striking a balance between simplicity and accuracy.
+Our framework ensures broader applicability across various medications where a majority of drugs are assumed to adhere to linear PK (Rosenbaum, 2017).
+The equations presented in the paper are shown in a patient-specific manner to highlight that the PK encoder generates individualized information.
+However, these equations can be applied to produce patient-specific outputs for multiple patients simultaneously using vectorized operations.
+3.4 Models
+To assess forecasting performance across various model architectures, we trained models with six different algorithms as baselines.
+We employed Exponential Smoothing (ETS; Hyndman 2008) as our statistical baseline model.
+Our deep learning baselines included the following models: Multi-Layer Perceptron (MLP; Rosenblatt 1958), NBEATS/NHITS Oreshkin et al.
+(2020); Challu et al.
+(2022), Recurrent Neural Network (RNN; Elman (1990); Cho et al.
+(2014)), Long-Short Term Memory (LSTM; Sak et al.
+(2014)), and Temporal Convolution Network with an MLP Decoder (TCN; Bai et al.
+(2018); van den Oord et al.
+(2016)).
+Implementations of these models were obtained from the Neuralforecast (Olivares et al., 2022) and StatsForecast (Garza et al., 2022) libraries.
+More information on these models and related prior work that leverages these are provided in Appendix A.2.
+Details on model training and hyperparameter selection are provided in Appendix A.3
+3.5 Baselines
+We compare our hybrid global-local architecture with the PK encoder to both local and global baselines:
+Global Baseline Models.
+Models were trained on all patient data to create a single global model:
+  • Baseline. Models are trained with the univariate blood glucose signal to demonstrate the forecasting improvement from using exogenous variables, supporting our PK encoder.
+  • Baseline Exogenous (Exog.). Models are trained with the univariate blood glucose signals, static features, and sparse exogenous features (CHO and insulin dose values).
+  • Baseline Exogenous (Exog.) - ST.
+Models are trained with the univariate blood glucose signal, static features, and sparse exogenous features transformed via the sum total (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022).
+Local Baseline Exogenous Models.
+Patient-specific baseline exogenous models were compared against our hybrid global-local PK encoder model.
+Local PK Encoder Models.
+Patient-specific models using the PK encoder were compared against our hybrid global-local PK encoder model.
+3.6 Evaluation
+Trained models were used to generate rolling window forecasts at one-step intervals.
+Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate the blood glucose forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE) computed across 8 trials of individually trained models.
+Formulations for these metrics are provided in Appendix A.4.
+We evaluate model performance across all points within the specified time horizon.
+We also assess the model’s performance exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL).
+For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation.
 4 Cohort
 Simulated data
-An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes.
+An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides blood glucose, CHO, and insulin values for 30 simulated patients with Type-I diabetes.
+We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set.
+Blood glucose time series plots for 6 subjects are shown Fig.
+7 in Appendix A.1.
+Basal and bolus dose distributions for 12 simuluted patients are shown in Fig.
+9 in Appendix A.1.
+OhioT1DM
+12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020).
+Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes.
+The data includes basal rate information, bolus insulin (units), and CHO values.
+The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days.
+Missing values were forward filled to prevent data leakage.
+The sample count and percent missing data for each subject are included in Table 2 in Appendix A.1.
+Blood glucose time series plots for 6 subjects are shown in Fig.
+8 in Appendix A.1.
+Basal and bolus dose distributions for the 12 patients are shown in Fig.
+10 in Appendix A.1.
+More dataset information is in Appendix A.1.
+5 Results
+5.1 Results on Simulated Data
+For the simulated dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous model for forecasts evaluated at all values (t-test p-value: 6.6e-6) as shown in as shown in Table 1.
+For evaluations across all values, the NHITS-PK model provides 3.5% improvement, on average, and up to 9.2% improvement for individual patients over the baseline NHITS sparse exogenous model.
+For evaluations across critical values, the NHITS-PK model provides 5.2% improvement, on average, and up to 16.4% improvement for individual patients over the baseline NHITS sparse exogenous model.
+Additionally, the NHITS-PK model outperforms the NHITS-ST model.
+Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 6 in Appendix C.5.
+5.2 Results on Real Data
+Our hybrid global-local architecture produces models that significantly outperform local models by 15.8%, on average (t-test p-value: 1.7e-5), and up to 31.3% improvement for individual OhioT1DM patients.
+We show that the NHITS-PK hybrid global-local model outperforms NHITS-PK models developed for individual OhioT1DM patients in Fig.
+12.
+Moreover, this improvement in forecast accuracy is consistent across all patients as shown in Fig.
+5 and in Table 4 in Appendix C.1.
+Additionally, the NHITS-PK model with the hybrid global-local architecture also improves MAE by 14.8%, on average, for patients over the patient-specific NHITS sparse exogenous baseline models (t-test p-value: 1.7e-6).
+For the OhioT1DM dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous baseline model for forecasts evaluated at all values (t-test p-value: 5.0e-3) and critical values (t-test p-value: 3.5e-3) as shown in Table 1.
+For evaluations across all values, The NHITS-PK model provides 1.3%  on average, and up to 3.5% improvement for individual patients over the baseline NHITS sparse exogenous model.
+For evaluations across critical values, the NHITS-PK provides 1.6%  on average, and up to 4.9% improvement for individual patients over the baseline NHITS sparse exogenous model.
+Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 7 in Appendix C.5.
+Fig.
+13 in Appendix C.2 shows model forecasts during hyperglycemic events for four randomly selected OhioT1DM patients.
+Figure 5: The NHITS-PK hybrid global-local model (blue, solid) provides more accurate predictions than local NHITS-PK models (purple, dashed).
+Figure 6: In addition to reducing forecasting error during hyperglycemic events, our NHITS-PK encoder enables improved lead time predictions of hypertensive events (blood glucose $\geq$ 180) for 9 of 12 OhioT1DM patients as measured over the 8 trials.
+Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
+Simulated Dataset
+Model All Values Critical Values
+Baseline Exog. Baseline Exog.
+ETS 9.450 13.457
+MLP 10.589 8.433 14.699 10.600
+RNN 13.703 20.190 21.433 35.955
+LSTM 13.557 10.321 20.537 13.550
+TCN 12.096 10.109 17.709 12.660
+NHITS 8.268 7.304 11.114 9.094
+NHITS-ST 7.055 8.648
+NHITS-PK 7.037 8.492
+OhioT1DM Dataset
+ETS 9.022 10.028
+MLP 11.065 11.120 12.630 12.698
+RNN 11.325 11.795 13.627 13.552
+LSTM 9.811 9.467 11.252 11.178
+TCN 11.525 10.760 13.647 12.678
+NHITS 9.060 8.873 10.337 10.137
+NHITS-ST 8.958 10.131
+NHITS-PK 8.758 9.965
+6 Discussion
+Hybrid Global-Local Architecture We demonstrate the efficacy of our novel hybrid global-local model, which produces forecasts with significantly smaller errors than patient-specific models across all patients.
+Our hybrid architecture not only improves accuracy but also enables global models to learn relevant patient-specific information, outperforming global model baselines.
+A hybrid global-local model offers practical benefits beyond improved forecasting accuracy as it requires less computational resources and time to train than individual patient models.
+Training patient-specific NHITS-PK models took seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space.
+The hybrid model can also be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
+PK Encoder As a model-agnostic extension, our novel PK encoder significantly reduces forecasting errors in deep learning models for both simulated and real-world data.
+These improvements are not only statistically significant, but they have potential for positive real-world clinical impact.
+In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts such events earlier, up to 3.6 minutes when compared to the NHITS baseline model, for individual OhioT1DM patients as shown in Fig.
+6.
+This work uses data from T1DM patients but incorporates key variables relevant to both T1DM and T2DM, such as continuous glucose measurements (CGM), carbohydrate intake, and insulin doses, making our method applicable to both populations.
+Improving prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009).
+Our results also highlight the importance of using domain knowledge to incorporate exogenous variables into models.
+Models with sparse features may underutilize valuable information, and feature-engineered representations not informed by domain knowledge, like the ‘sum total’ approach, can underperform compared to baselines.
+Moreover, Appendix C.4 shows that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset.
+As such, dose-dependent models, such as proposed in section 3.3, may be needed for more realistic data simulations and models.
+Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing or drug selection.
+It can also be adapted for other medication administration types, such as oral or intravenous injections.
+Limitations.
+We equate $\sigma$ to our k parameter in the log-normal function and hold $\mu$ constant at 1.
+This simplification reduces the number of potential concentration curve shapes the model can learn.
+Tuning both $\sigma$ and $\mu$ could further improve forecasts by enabling more flexible concentration curves.
+Future Work.
+Future work will explore fine-tuning global forecasting models with patient-specific parameters for new patients with limited historical data as well as modeling dose-dependent effects using more dynamic nonlinear PK parameter frameworks.
+Future work will include applying our approach to additional patient cohorts, such as those with T2DM, to further validate our findings.
+[…]
+This work was partially supported by the NIH award R01HL141916. The authors wish to thank Dr. Frank Guyette, Dr. Leonard Weiss and Dr. Gilles Clermont for their guidance and informative conversations.
+Appendix A Supplemental Methods
+Data and Code Availability
+We use open-source simulated and real-world data that is available to other researchers.
+The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository [Xie, 2018, Visentin et al., 2016].
+The real-world data includes the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020].
+Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu [2020].
+Implementations of models used in our work were obtained from the open-source Neuralforecast [Olivares et al., 2022] and StatsForecast [Garza et al., 2022] libraries.
+We implemented our method using a copy of the Neuralforecast repository to leverage their model training framework.
+Our research code, along with the Neuralforecast models used in our research, is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
+A.1 Data
+Simulated data
+An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) [Xie, 2018, Visentin et al., 2016] provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes.
 Clinical parameters include information on age and weight, and biological parameters include information on insulin-glucose kinetics, such as absorption constants.
 We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set.
-In a similar approach to (Rubin-Falcone et al., 2022), the meal schedule used to generate simulated data was based on the Harrison-Benedict equation (Arthur Harris and Gano Benedict, 21919) with approximately 3 meals per day and no additional snacks.
+In a similar approach to [Rubin-Falcone et al., 2022], the meal schedule used to generate simulated data was based on the Harrison-Benedict equation [Arthur Harris and Gano Benedict, 21919] with approximately 3 meals per day and no additional snacks.
 Height data was not provided in the patient parameters, and so was estimated at 140cm, 170cm, and 175cm for adults, adolescents, and children, respectively.
@@
 The default basal-bolus controller provided in the simulation was used to administer insulin at the time of each meal, where the bolus amount is computed based on the current glucose level, the target glucose level, the patient’s correction factor, and the patient’s carbohydrate ratio.
-However, we adapt the controller to deliver basal doses once per hour, consistent with the OhioT1DM dataset.
-The resulting simulated data consists of CGM (in mg/dL) values, insulin rate (in units per minute), and carbohydrates (in grams), referred to as CHO.
+However, we adapt the controller to deliver basal doses once per hour, consistent with the function of real insulin pumps used in the OhioT1DM dataset.
+The resulting simulated data consists of CGM (in mg/dL) values, insulin (units), and carbohydrates (in grams), referred to as CHO.
 We extract the bolus insulin rate from the overall insulin rate as the difference in the continuous basal insulin rate, which is consistent hourly doses.
@@
 7 in Appendix A.1.
-Figure 3: Instantaneous medication administrations are represented as sparse variables in time series data.
-We propose a pharmacokinetic encoder to effectively capture time-dependent plasma drug concentration.
+Basal and bolus dose distributions for 12 simuluted patients are shown in Fig.
+9 in Appendix A.1.
 OhioT1DM
-12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020).
+12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020].
 Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes.
@@
 Bolus insulin consists of both ‘normal’ insulin, which is recorded as an instantaneous administration, and ‘square dual’, where the insulin dose is stretched over a defined period.
-‘Square dual‘ insulin doses were divided evenly across the specified administration window with the assumption that insulin units are administered consistently across specified timestamps.
+‘Square dual’ insulin doses were divided evenly across the specified administration window with the assumption that insulin units are administered consistently across specified timestamps.
 If multiple bolus administrations occurred within the same minute, the insulin units were summed.
-Basal insulin was included as hourly doses.
+For Medtronic insulin pumps 530G and 630G versions, basal insulin is administered hourly based on the basal rate.
 For timestamps with overlapping basal rates, the last basal rate value was considered.
@@
 8 in Appendix A.1.
-3 Methods
-We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), based on a 10 hour history of blood glucose, carbohydrate, insulin basal, and insulin bolus values.
-3.1 Pharmacokinetic Encoder
-We propose a novel pharamacokinetic encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes.
-The encoder consists of a function, $C$, that takes in sparse exogenous medication variables and generates a concentration-time profile for a specified dose, $d$, using the following equation:
-$C(t,d,\textbf{k}^{(i)})=\frac{d}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
-where $t$ is the time and $\textbf{k}^{(i)}$ is the absorption rate constant specific to patient, $i$.
-Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, $d$.
-Concentration profiles are generated under the assumption that all insulin administrations adhere to 100% bioavailability.
-As bioavailability equates to the area under the concentration curve, the scaled curve thus integrates to the specified dose, $d$.
-Let $N$ be the number of subjects in the data.
-The absorption rate constant, $\textbf{k}_{\text{bolus}}\in[0,1]^{N\times 1}$ and $\textbf{k}_{\text{basal}}\in[0,1]^{N\times 1}$ are represented as network embedding weights that are initialized between values of 0 and 1 prior to model training.
-Initial $\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{bolus}}$ values are tuned as model hyperparameters.
-During model training, sparse basal and bolus features are passed into the pharmacokinetic encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig.
-4.
-The encoder than replaces sparse medication features in the training window with the corresponding concentration curve.
-Non-medication features, such as CHO, are unaffected.
-For each training step, $\textbf{k}^{(i)}_{\text{bolus}}$ and $\textbf{k}^{(i)}_{\text{basal}}$ are updated for each patient, $i$, using gradient descent optimization.
-For a single dose, $d$, of bolus insulin, $d_{\text{bolus}}$, or basal insulin, $d_{\text{basal}}$, administered at time, $t$, we obtain the concentration, $c$, at that time:
-$c^{(i)}_{\text{bolus}}$ $=C(t,d^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)})$ (2)
-$c^{(i)}_{\text{basal}}$ $=C(t,d^{(i)}_{\text{basal}},\textbf{k}_{\text{basal}}^{(i)})$ (3)
-Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig.
-3.
-The encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time, to perform vectorized operations that enable generating concentration curves for multiple doses in an input window in $\mathcal{O}(1)$ time.
-Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row.
-Concentration curves are generated for each row of w using the dose at the corresponding time index.
-The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps.
-w $=\begin{bmatrix}0&1&2&3&\cdots&L-1\\ 0&0&1&2&\cdots&L-2\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ 0&0&0&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{L\times L}$ (8)
-$\textbf{c}^{(i)}$ $=\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}^{(i)}_{t},\textbf{k}^{(i)})$ (9)
-Here, $f$ is the frequency of the data, $\textbf{d}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of doses, and $\textbf{c}^{(i)}\in\mathbb{R}^{1\times L}$ is the new plasma insulin concentration feature for patient, $i$.
-For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of concentration curves, c, other non-medication time series, x, and static features, s,
-$\hat{\textbf{y}}^{(i)}_{t:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\textbf{c}^{(i)}_{t-L:t},\textbf{x}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (10)
-In the proposed context, $\textbf{c}_{t-L:t}$ includes concentration curves for bolus insulin, $\textbf{c}_{\text{bolus}}$, and basal insulin $\textbf{c}_{\text{basal}}$, x includes CHO intake, and s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type, if applicable.
-Figure 4: Hybrid Global-Local architecture.
-Global architectures with parameters shared across patients’ time series with local parameters from a patient-specific pharmacokinetic (Ph-K.) encoder.
-Sparse medication dose time series are passed into the pharmacokinetic encoder along with their respective $k_{a}$ parameters to generate concentration curves.
-Concentration curves are then passed into a deep learning model which learns relevant hidden states (hidden) and output parameters, $\hat{\theta}$.
-3.2 Models
-To assess forecasting performance across various model architectures, we trained models with six different algorithms as baselines.
-The implementations of these models were obtained from the Neuralforecast (Olivares et al., 2022b) and StatsForecast (Garza et al., 2022) libraries.
-More information on these models is provided in Appendix A.2.
-Statistical Model
-We employed Exponential Smoothing (ETS; Hyndman 2008) as our statistical baseline model.
-Deep Learning Models
-Our deep learning baselines included the following models: Multi-Layer Perceptron (MLP; Rosenblatt 1958), NBEATS/NHITS Oreshkin et al.
-(2020); Challu et al.
-(2022), Recurrent Neural Network (RNN; Elman (1990); Cho et al.
-(2014)), Long-Short Term Memory (LSTM; Sak et al.
-(2014)), Temporal Convolution Network with an MLP Decoder (TCN; Bai et al.
-(2018); van den Oord et al.
-(2016)), and Temporal Fusion Transformer (TFT; Lim et al.
-(2021))
-In contrast to previous efforts that focused on training models for individual subjects, or local models, we employ global architectures with parameters shared across subjects Rubin-Falcone et al.
-(2022); Xie and Wang (2020).
-For each dataset, models were trained using all patient data to generate one global model that was applied across patients.
-We showcase the efficacy of global architectures in producing improved forecasts by training local models for comparison against our global model.
-Moreover, apart from reducing computational and storage costs, global models could prove more applicable in real-world scenarios since they can be utilized for new patients with limited or no available training data.
-Models were trained using Adam optimizer (Kingma and Ba, 2017).
-Optimal hyperparameters were selected based on a validation set with cross-validation.
-Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected.
-Huber estimator was selected as a robust loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
-$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)$ =$=\begin{cases}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta\cdot(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{cases}$ (11)
-Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set.
-The deep learning models were trained using an NVIDIA A100 Tensor Core GPU.
-Figure 5: The boxplots show average MAE computed across 8 trials for global (orange) and local (blue) models for OhioT1DM subjects.
-Models trained on all subject data, or global models, have lower MAE, on average, than models trained on individual subjects, or local models.
-3.3 Baselines
-We implement three baselines against which we compare our proposed hybrid global-local architexture and pharmacokinetic encoder approach:
-Univariate signal
-We train models with the univariate blood glucose signal.
-Univariate signal with exogenous features
-We train models with the univariate blood glucose signal, static features, and exogenous features, consisting of sparse basal and bolus insulin dose values.
-Univariate signal with sum-total feature engineering
-We train models with the univariate blood glucose signal, static features, and exogenous features, consisting of insulin dose values.
-The sum total feature preprocessing approach was used to transform sparse features into non-sparse features of cumulative dose across an input window, as shown in Fig.
-3.
-3.4 Evaluation
-Trained models were used to generate rolling window forecasts at one-step intervals.
-In accordance with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate the blood glucose forecasts using two metrics: mean absolute error (MAE) and root mean squared error (RMSE) computed across 8 trials of individually trained models.
-Mathematical formulations for these metrics are provided in Appendix A.4.
-We evaluate model performance across all points within the specified time horizon.
-Furthermore, given the critical importance of healthcare outcomes related to blood glucose levels exceeding or falling below safe thresholds, we also assess the model’s performance exclusively at time points where blood glucose levels reach or fall below certain thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL).
-For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation.
-Our code implementation is open-source and can be found on Github. [^1] https://github.com/Nixtla/neuralforecast/tree/treat_module
-4 Results
-The pharmacokinetic encoder significantly improves forecasting accuracy for the top performing deep learning model by approximately 4.4% and 2.1%, for simulated and OhioT1DM datasets, respectively, as shown in Table 1.
-For the OhioT1DM dataset, the NHITS-Pharamacokinetic model has significantly smaller MAE than the NHITS-Sparse Exogenous model for forecasts evaluated at all values (t-test p-value: 4.44e-05) and critical values (t-test p-value: 1.6e-2).
+Basal and bolus dose distributions for the 12 patients are shown in Fig.
+10 in Appendix A.1.
 Fig.
-6 demonstrates improved forecasting performance using the the NHITS-Pharamacokinetic model during hyperglycemic and hypoglycemic events for three randomly selected OhioT1DM subjects.
-Mean MAE and RMSE results with standard deviations computed across 8 trials are shown in Tables 7 and 8 in Appendix B.3.
-We showcase the efficacy of our novel hybrid global-local model in producing more accuracy forecasts than models trained per patient.
-Global models consistently outperform local models, on average, for each OhioT1DM subject as shown in Table 4 and Fig.
-9 in Appendix B.1.
-For the OhioT1DM dataset, the hybrid global-local pharmacokinetic model achieves a significantly lower (t-test p-value: 1.29e-2) MAE by 14.6% over patient-specific pharmacokinetic models.
+7 shows the first 10 days of data for 6 subjects from the simulated dataset.
 Fig.
-5 shows that the global-local pharmacokinetic model has lower MAE, on average, across all OhioT1DM patients compared to the local models.
-Furthermore, the hybrid global-local pharmacokinetic model also achieves a significantly lower (t-test p-value: 3.83e-2) MAE by 9.2% over the patient-specific sparse exogenous models.
-Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
-Simulated Dataset
-Model All Values Critical Values
-Baseline Exogenous Baseline Exogenous
-ETS 10.040 13.531
-MLP 9.734 8.653 12.189 9.887
-LSTM 8.731 8.372 10.863 9.996
-TCN 16.772 12.234 28.913 18.122
-TFT 7.788 6.762 9.664 7.828
-NHITS 8.333 7.282 10.184 8.353
-TFT- Sum-Total 6.698 7.634
-NHITS- Pharmacokinetic 7.027 7.917
-TFT- Pharmacokinetic 6.466 7.514
-OhioT1DM Dataset
-ETS 10.126 11.781
-MLP 10.568 10.275 11.805 11.434
-LSTM 10.604 9.606 12.999 11.144
-TCN 11.011 11.090 12.693 13.212
-TFT 9.605 9.515 10.686 10.555
-NHITS 9.234 8.954 10.436 10.117
-NHITS- Sum-Total 8.988 10.092
-NHITS- Pharmacokinetic 8.769 10.007
-TFT- Pharmacokinetic 9.492 10.484
-Figure 6: Blood glucose forecasts of the final horizon point for three OhioT1DM subjects.
-The NHITS -Pharmacokinetic model (orange) produces noticeably improved forecasts of the ground truth (blue) compared to the NHITS model with sparse exogenous features (yellow).
-5 Discussion
-Our pharmacokinetic encoder enables deep learning models to achieve significantly lower forecasting errors.
-As such, models that only include sparse features may be under-leveraging the potential utility of exogenous information.
-Our pharmacokinetic encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings about unexpected treatment response, or helping to characterize patient-specific treatment effects in terms of drug absorption and elimination characteristics.
-The characterization of patient-specific pharmacokinetic parameters holds substantial potential for advancing the development of more effective treatments by tailoring medication to individual patients.
-Assessing a forecasting model’s ability to effectively capture the impact of variables on the target signal is essential for ensuring the reliability of forecasts.
-We examine whether the pharmacokinetic model can accurately capture the treatment effects of insulin on blood glucose by generating forecasts during time periods for three scenarios: with the original bolus insulin doses, with bolus insulin doses removed, and with bolus insulin doses increased by a factor of 10.
-When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig.
-10 in Appendix B.2.
-Our model thus effectively learns the inhibitory treatment effects of insulin on blood glucose.
-Bolus insulin is quick-acting and often taken before meals compared to basal insulin, which is longer-acting.
-Our pharamcokinetic encoder learns $k_{a}$ values that align with this established knowledge.
-For OhioT1DM data, the average $k_{a_{bolus}}$ and $k_{a_{basal}}$ were 1.622 and 1.486, respectively.
-For our approach, larger $k_{a}$ values correspond to concentration curves with shorter duration, or steeper curves.
-The average $k_{a}$ values for bolus insulin were higher than those for basal insulin, on average, resulting in shorter durations for bolus insulin concentration profiles.
-To evaluate whether the difference in $k_{a_{bolus}}$ and $k_{a_{basal}}$ values was statistically significant, we compute a one-sided paired t-test of model-inferred $k_{a}$ values for patients across the 8 trials.
-The alternative hypothesis claims that $k_{a}{{}_{bolus}}$ values are larger than $k_{a_{basal}}$.
-The test yielded a p-value of 0.043 indicating that the $k_{a}$ parameter for bolus insulin is significantly higher than that for basal insulin when considering a significance level of 0.05.
-Thus, our approach can learn the relative differences in pharamcokinetic characteristics between basal and bolus insulin.
-Limitations.
-We equate $\sigma$ to our $k_{a}$ parameter in the log-normal function and hold $\mu$ constant at 1.
-This simplification reduces the number of potential concentration curve shapes the model can learn.
-Tuning both $\sigma$ and $\mu$ would enable more flexible concentration curves with potentially more accurate duration and peak characteristics, which could further improve forecasts.
-Future Work.
-Future work will consider adaptations to the concentrator log-normal function to enable more flexible concentration curves.
-Future work will also consider expanding our pharmacokinetic encoder with functions relevant to other medication administration methods, such as intravenous injections.
-Finally, another interesting line of research is determining how to leverage the advantages of the global forecasting model with patient-specific $k_{a}$ values for new patients.
-[…]
-This work was partially supported by the NIH award R01HL141916. The authors wish to thank Dr. Frank Guyette, Dr. Leonard Weiss and Dr. Gilles Clermont for their guidance and informative conversations.
-Appendix A First Appendix
-A.1 Data
-Fig. 7 shows the first 10 days of data for 6 subjects from the simulated dataset.
-Fig. 8 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset.
+8 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset.
+Figs.
+9 and 10 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively.
 Table 2 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
-Table 2: Sample count (miss- ing samples) for the original OhioT1DM datasets and prepro- cessed test set
+Table 2: Sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set
 ID Original Preprocessed
@@
 (20.20%) (8.66%) (9.66%)
-Figure 7: Time series examples for the first 10 days of 6 subjects from the simulated dataset.
-Figure 8: Time series for the first 10 days of 6 subjects from the OhioT1DM dataset.
 A.2 Models
-Exponential Smoothing (ETS) The ETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. (Hyndman, 2008).
-Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
-Neural Hierarchical Interpolation for Time Series (NHITS) - A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands.
-The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation (Challu et al., 2022).
+The baseline models used in this work have been extensively used in previous blood glucose forecasting work.
+For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies [Qian et al., 2021, Li et al., 2019a, Fox et al., 2018, Rabby et al., 2021, Rubin-Falcone et al., 2020, 2022, Miller et al., 2020].
+Similarly, CNNs [Xie and Wang, 2020, Li et al., 2019a, b] and MLPs [McShinsky and Marshall, 2020, Miller et al., 2020] have been utilized in proposed models and as baselines.
+NHITS achieves state-of-the-art performance, surpassing RNN-based, transformer-based, and other deep learning time series decomposition models (such as N-BEATS [Oreshkin et al., 2020]) in long-horizon forecasting tasks [Challu et al., 2022].
+Exponential Smoothing (ETS) The ETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. [Hyndman, 2008].
+Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation [Nair and Hinton, 2010, Fukushima, 1975, Rosenblatt, 1958].
+Neural Hierarchical Interpolation for Time Series (NHITS) A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands.
+The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation [Challu et al., 2022].
 Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states.
-The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts.(Elman, 1990; Cho et al., 2014)
-Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.(Sak et al., 2014)
+The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts.[Elman, 1990, Cho et al., 2014]
+Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.[Sak et al., 2014]
 Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts.
-To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions (Bai et al., 2018; van den Oord et al., 2016).
-Temporal Fusion Transformer (TFT) An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies.
-TFT also leverages variable selection networks as well as a series of gating layers to suppress unnecessary parts of the architecture (Lim et al., 2021).
-A.3 Hyperparameters
+To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions [Bai et al., 2018, van den Oord et al., 2016].
+A.3 Model Training and Hyperparameters
+Models were trained using Adam optimizer [Kingma and Ba, 2017].
+Optimal hyperparameters were selected based on a validation set with cross-validation.
+Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected.
+Huber estimator was selected as a robust loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
+$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)$ =$=\begin{cases}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta\cdot(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{cases}$ (8)
+Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set.
+The deep learning models were trained using an NVIDIA A100 Tensor Core GPU.
 For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 3.
 Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 3.
+We tuned the hyperparameters of our model using ‘HyperOptSearch’ with the Tree-structured Parzen Estimator (TPE) algorithm and 20 samples.
 Hyperparameter Considered Values
-Learning rate Loguniform(1e-4, 1e-2)
+Learning rate Loguniform(1e-4, 1e-1)
 Batch size 4
-Dropout {0.0, 0.1, 0.2, 0.3, 0.4, 0.5}
-Training steps {1000, 2000}
+Windows batch size’ 256
+Dropout 0.0
+Training steps 2000
+Validation check steps 100
+Early stop 5
+patience steps
 Random seed DiscreteRange(1, 10)
-Basal dose initial $k_{a}$ Uniform(1, 2, 11)
-Bolus dose initial $k_{a}$ Uniform(1, 2, 11)
+Initial $\textbf{k}_{\text{basal}}$1.0, 1.1, 1.2}
+Initial $\textbf{k}_{\text{bolus}}$1.7, 1.8, 1.9}
+Initial $\textbf{k}_{\text{CHO}}$1.7, 1.8, 1.9}
 Table 3: Common hyperparameter search space
 We tune additional specific hyperparameters for each architecture.
-For the MLP  we tune the number of hidden units per layer ([128,256]); for the LSTM  the hidden size ([128,256]); for the TCN the hidden size ([128,256]) and number of dilations ([2,4,8]); for the TFT the hidden size ([128,256]); for the NHITS we did not tune additional hyperparameters.
+For the MLP  we tune the number of hidden units per layer ([128,256]); for the LSTM  the hidden size ([128,256]); for the TCN the hidden size ([128,256]) and number of dilations ([2,4,8]); for the NHITS we did not tune additional hyperparameters.
 The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
 A.4 Metrics
+We use the following metrics to evaluate model performance. Here, $H$ refers to the forecast horizon.
 Mean Absolute Error (MAE)
@@
 $RMSE(y_{t+1,t+H},\hat{y}_{t+1,t+H})=\sqrt{\frac{1}{H}\sum_{i=t+1}^{t+H}(y_{i}-\hat{y}_{i})^{2}}$
-Appendix B Second Appendix
-B.1 Patient-specific (local) models
+A.5 Propositions
+Proposition A.1.
+Under linear pharmacokinetic (PK) assumptions, the area under the concentration curve is directly proportional to the dose of the drug.
+As such, PK parameters, such as bioavailability, are assumed to be constant for a given drug.
+The bioavailability, $F$, can be assessed from the dose-normalized area under the concentration curve (AUC).
+When bioavailability is constant, we show that the AUC is bounded by the dose, $x$, given that $0\leq F\leq 1$,
+$F$ $\propto\frac{\text{AUC}}{x}$ (9)
+$n\cdot F$ $=n\cdot\frac{\text{AUC}}{x}$ (10)
+AUC $=F\cdot x$ (11)
+$\leq x$ (12)
+where $n$ where is a constant of proportionality.
+Proposition A.2.
+Given some mathematical model, $C(t,\cdot)_{x}$, of concentration for a drug dose, $x$, that is a function of time, $t$, and pharmacokinetic parameters, $\cdot$, we show that the overall effect (concentration) of $n$ drug doses of the same drug is equal to the sum of the individual effects under linear pharmacokinetic assumptions:
+[…]
+[…]
+[…]
+[…]
+$=\int_{0}^{T}\sum_{i=1}^{n}C(t,\cdot)_{x}dt$ (17)
+Here, $T$ represents the total duration of drug exposure or the time interval over which drug concentrations are being measured and $t$ represents a specific point in time within that interval.
+A.6 Algorithm
+We outline the algorithm for training our proposed PK encoder that transforms sparse exogenous variables into time-dependent concentration profiles.
+This framework can be implemented in any deep learning architecture.
+It can also be adapted to utilize other concentration-time profile functions pertinent to a specific healthcare context.
+The algorithm initializes k as a network embeding weight using a value $\textbf{k}^{*}$ from the specified range $[0,1]$ at the beginning of each epoch.
+This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library [Liaw et al., 2018].
+For each training step, the algorithm iterates over batches of data.
+It computes a context vector c based on the input data x, k, and other parameters.
+It then updates the hyperparameters k for each subject using stochastic gradient descent (SGD) based on the loss function $\mathcal{L}$ between the predicted $\hat{\textbf{y}}$ and actual y values.
+After a defined number of training steps (determined by a validation step parameter), the algorithm performs a validation step.
+It calculates the total loss $\ell(\textbf{k})_{\text{total}}$ over the validation set by summing the loss $\ell(\textbf{k})$ and the current batch loss $\mathcal{L}(\textbf{y},\hat{\textbf{y}})$.
+It then updates the hyperparameter vector $\textbf{k}^{*}$ to minimize the total loss.
+Finally, the algorithm returns the final k values.
+Algorithm 1 PK Encoder Training.
+$E$ is the number of trainings steps, $\eta$ is the learning rate, $N$ is the number of subjects, $H$ is the forecast horizon, and $L$ is the lag time.
+Data: $N>0,f\in\mathbb{R}$, $L\in\mathbb{R}$, $H\in\mathbb{R}$
+Result: k
+$\mathcal{B}_{train}\leftarrow$ (split training data into batches)
+$\mathcal{B}_{val}\leftarrow$ (split validation data into batches)
+$\textbf{k}^{*}\leftarrow$ (Randomly select $\textbf{k}^{*}$ from defined hyperparmeter options, $\textbf{k}^{*}\in\{1,\ldots,2\}$)
+for each step i from 1 to E do
+       $\textbf{k}\leftarrow\textbf{k}^{*}$triangleright Initialize k
+for batch $b\in\mathcal{B}_{train}$ do
+               $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}_{t},\textbf{k})$
+$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
+$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
+$\textbf{k}\leftarrow\textbf{k}-\eta\cdot\nabla_{\textbf{k}}\ell(\textbf{k})$triangleright Update k using SGD
+       end for
+      if i % validation step == 0 then
+               $\ell(\textbf{k})\leftarrow 0$
+for batch $b\in\mathcal{B}_{val}$ do
+                     $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}_{t},\textbf{k})$
+$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
+$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
+             end for
+            $\textbf{k}^{*}\leftarrow\arg\min_{\textbf{k}^{*}\in\{1,\ldots,2\}}\ell(\textbf{k})_{total}$triangleright Select the hyperparameter that minimizes loss
+       end if
+end for
+return k
+Figure 7: Time series for the first 10 days of 6 subjects from the simulated dataset.
+Figure 8: Time series for the first 10 days of 6 subjects from the OhioT1DM dataset.
+Figure 9: Density plots of bolus and basal insulin for 12 simulated data subjects. Bolus doses are general larger than basal doses.
+Figure 10: Density plots of bolus and basal insulin for OhioT1DM subjects. Bolus doses are general larger than basal doses.
+Appendix B Pharmacokinetic Background
+Concentration-time Profiles
+From PK literature, exponential decay functions are commonly used to model drug concentration over time for intravenous injections, where the drug is directly injected into the bloodstream, leading to instantaneous effects [Rosenbaum, 2017, Angelika et al., 2005].
+Alternatively, subcutaneous injections involve absorption and elimination processes where the peak onset time may not be instantaneous depending on the drug profile.
+Regarding knowledge on concentration-time profiles established through empirical and invasive studies, we know that concentration-time profiles for subcutaneous injections adhere to right-skewed curves.
+The degree of right-skew in the concentration curve will differ depending on the type of insulin, as illustrated in Fig.
+2, along with various patient-specific factors.
+Insulin Analogs
+Insulin analogs are modified forms of insulin designed to exhibit certain PK properties, such as changes in drug onset and duration of action.
+Fig.
+2 is adapted from PK literature and depicts PK modeling of plasma insulin concentration over time for different insulin analogs [Angelika et al., 2005].
+Generally, bolus doses are quick-acting and, thus, often taken before meals.
+Contrastingly, basal insulin is longer-acting and serves to regulate blood glucose levels throughout the day.
+However, continuous glucose monitors (CGMs) and insulin pumps generally only hold and administer one type of insulin, such as rapid-acting insulin.
+For these devices, ‘bolus’ and ‘basal’ are terms used to refer to different timings of injections for different purposes.
+Bolus doses are administered discretely, generally around meals, while basal doses are administered continuously at a defined frequency, such as hourly, to regulate blood glucose levels throughout the day.
+Bioavailability
+In Fig.
+2, the area under the concentration curve measures drug bioavailability, the fraction of the absorbed dose that reaches the intended site of systemic circulation intact.
+Bioavailability varies based on many factors, including type of medication, dosage, and patient-specific characteristics.
+For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent [Søeborg et al., 2012].
+However, for subcutaneous injections, bioavailability is typically lower due to unintended drug absorption or elimination.
+First-order Kinetics
+First-order kinetics, a common assumption in pharmacokinetics, describes the rate at which a drug is eliminated from the body as being proportional to its concentration.
+This means that a constant fraction of the drug is eliminated per unit of time, regardless of its concentration levels.
+Exponential functions are often used to represent concentration-time profiles under the assumption of first-order kinetics, simplifying the modeling of plasma drug concentration in the body.
+Linear Pharmacokinetics
+Linear pharmacokinetics refers to a situation where the relationship between the dose of a drug and its concentration in the body is linear [Rosenbaum, 2017].
+In linear pharmacokinetics, the area under the concentration curve is directly proportional to the dose of the drug.
+If you double the dose of a drug, the area under the curve will also double.
+Linearity results from the assumption that the drug processes exhibit first-order kinetics and that the PK parameters remain constant with dose, indicating no significant changes in the body’s ability to eliminate the drug with higher doses.
+The majority of drugs used in clinical practice are assumed to follow linear pharmacokinetics [Rosenbaum, 2017].
+Figure 11: When doses of insulin are administered at close intervals, the subsequent dose may start to act before the previous dose has fully finished its effect.
+This can lead to a stacking of insulin effects.
+The figure depicts additive effects under linear pharmacokinetic assumptions.
+Appendix C Additional Results
+C.1 Patient-specific (Local) Models
 We showcase the effectiveness of global models, with parameters shared across subjects, in generating more accuracy forecasts compared to local models trained for individual subjects.
 Global models consistently outperform local models, on average across 8 trials, for each OhioT1DM subject as shown in Table.
-4 and Fig 9b.
-Furthermore, the global NHITS-Pharamcokinetic model consistently outperforms the global NHITS-Exogenous model, on average, for each OhioT1DM subject as shown in Fig 9a.
+4 and Fig.
+5.
+Furthermore, the global NHITS-PK model consistently outperforms the global NHITS-Exogenous model, on average, for OhioT1DM subjects as shown in Fig.
+12.
 Table 4: MAE computed across forecast horizon values using local and global models.
-ID Exogenous Pharmacokinetic
+ID Exogenous PK
 Local Global Local Global
-540 11.358 10.325 12.059 10.050
+540 12.106 10.119 11.880 10.015
 (0.779) (0.139) (0.596) (0.147)
-544 7.513 7.312 7.839 7.028
+544 7.952 7.188 8.067 6.967
 (0.679) (0.062) (0.490) (0.083)
-552 8.400 7.547 8.854 7.308
+552 8.559 7.400 9.132 7.374
 (0.243) (0.118) (0.403) (0.070)
-559 9.888 9.093 10.486 9.000
+559 10.845 9.013 13.051 8.967
 (1.266) (0.176) (7.027) (0.102)
-563 9.247 8.610 9.172 8.580
-(0.102) (0.071 (0.302) (0.107)
-567 11.254 9.835 11.455 9.662
+563 9.527 8.731 9.427 8.506
 (0.379) (0.065) (0.402) (0.084)
-570 8.160 7.470 8.259 7.280
+567 12.599 9.640 12.132 9.562
 (1.649) (0.138) (1.667) (0.189)
-575 10.500 9.734 11.149 9.539
+570 8.361 7.523 8.466 7.258
 (0.496) (0.141) (0.223) (0.121)
-584 10.605 9.958 13.706 9.787
+575 10.891 9.712 11.095 9.496
 (0.419) (0.099) (0.470) (0.151)
-588 8.966 8.295 9.454 8.158
+584 11.111 9.846 11.518 9.905
 (0.547) (0.091) (0.538) (0.156)
-591 11.237 10.344 11.275 10.103
+588 10.172 8.172 9.376 8.137
 (1.737) (0.081) (0.321) (0.087)
-596 8.797 8.230 9.471 8.008
+591 11.987 10.246 11.428 10.151
 (1.274) (0.083) (0.172) (0.134)
-All 9.661 8.954 10.265 8.769
+596 8.990 8.133 9.147 8.059
 (0.246) (0.103) (0.502) (0.141)
-Figure 9: The pharmacokinetic NHITS  model outperforms the NHITS baseline exogenous model across all patients, on average (a).
-Additionally, the pharmacokinetic model trained on data from subjects, or the global model, achieves lower MAE than models trained on individual subjects, or local models (b).
-B.2 Forecasting treatment effects
-In this appendix, we explore the effect of changing the insulin dosage on the forecasts produced by our approach.
-Fig.
-10, demonstrates that the pharmacokinetic model appropriately learns the treatment effect of insulin on blood glucose for OhioT1DM subjects, decreasing the forecast when the dosage increases.
-Figure 10: Treatment effect of insulin on the NHITS pharmacokinetic model.
+All 10.324 8.873 10.425 8.758
+(0.401) (0.064) (0.659) (0.080)
+Figure 12: The NHITS-PK hybrid global-local model (Hybrid) trained on all OhioT1DM subject data (blue, solid) has a lower MAE, on average, than NHITS-PK models trained on individual subjects, or local models (purple, dashed) .
+C.2 Forecast Comparison
+Figure 13: Blood glucose forecasts of the final horizon point during 5 randomly selected hypoglycemic threshold periods (blood glucose $\geq$ 180) for 4 randomly selected OhioT1dm patients.
+The NHITS-PK hybrid global-local model (orange, dashed) produces improved forecasts of the ground truth (blue, solid) compared to the NHITS baseline model with sparse exogenous features (yellow, dash-dotted).
+C.3 Forecasting Response to Treatment
+Predicting treatment effects using accurate forecasting algorithms can assist individuals with diabetes and their care providers in making informed decisions about the ideal medication type and dosage, leading to improved glycemic control and reduced risk of both short- and long-term complications.
+Assessing a forecasting model’s ability to effectively capture the impact of variables on the target signal is essential for ensuring the reliability of forecasts.
+We examine whether the NHITS-PK model can accurately capture the treatment effects of insulin on blood glucose by generating forecasts during time periods for three scenarios: with the original bolus insulin doses, with bolus insulin doses removed, and with bolus insulin doses increased by a factor of 10.
+When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig.
+14.
+The NHITS-PK model thus effectively learns the inhibitory treatment effects of insulin on blood glucose.
+Figure 14: Treatment effect of insulin on the NHITS-PK hybrid global-local model.
 Each point in the plot corresponds to the final horizon point.
-Forecasts with original bolus doses (orange), removing bolus insulin doses results in higher forecasts (blue), and augmenting doses results in lower blood glucose forecasts (yellow).
-B.3 Model MAE and RMSE results
-This appendix presents the complete results for all models, metrics of evaluation, and both simulated data and OhioT1DM. Standard deviation across 8 trials are provided in parenthesis.
+Forecasts with original bolus doses (orange, solid), removing bolus insulin doses results in higher forecasts (blue, dashed), and augmenting doses results in lower blood glucose forecasts (yellow, dotted).
+C.4 Forecasting with Dose-Independent Parameters
+[…]
+We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values.
+However, this approach led to worse performance on the OhioT1DM dataset.
+The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters [Xie, 2018, Visentin et al., 2016].
+In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3.
+Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data.
+These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
+Table 5: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
+Simulated Dataset
+Model All Values Critical Values
+Baseline Exog. Baseline Exog.
+NHITS 8.268 7.304 11.114 9.094
+(0.039) (0.055) (0.237) (0.236)
+NHITS-ST 7.055 8.648
+(0.038) (0.068)
+NHITS-PK 7.103 8.449
+linear (0.069) (0.139)
+NHITS-PK 7.037 8.492
+nonlinear (0.093) (0.220)
+OhioT1DM Dataset
+NHITS 9.060 8.873 10.337 10.137
+(0.142) (0.064) (0.137) (0.108)
+NHITS-ST 8.958 10.131
+(0.087 ) (0.100)
+NHITS-PK 8.830 10.026
+linear (0.060) (0.064)
+NHITS-PK 8.758 9.965
+nonlinear (0.080) (0.089)
+C.5 Model MAE and RMSE Results
+This appendix presents the complete results for all models for both simulated data and OhioT1DM.
+Standard deviation across 8 trials are provided in parenthesis.
+Table 6 provides complete results for the simulated data and Table 7 provides complete results for the OhioT1DM dataset.
 Table 6: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated data.
@@
 Model All Values Critical Values
-Baseline Exogenous Baseline Exogenous
-ETS 10.040 13.531
+Baseline Exog. Baseline Exog.
+MAE
+ETS 9.450 13.457
 (0.00) (0.00)
-MLP 9.734 8.653 12.189 9.887
+MLP 10.589 8.433 14.699 10.600
 (2.085) (0.490) (3.180) (0.464)
-LSTM 8.731 8.372 10.863 9.996
+RNN 13.703 20.190 21.433 35.955
 (2.645) (9.966) (6.123) (23.235)
-TCN 16.772 12.234 28.913 18.122
+LSTM 13.557 10.321 20.537 13.550
 (9.112) (2.074) (18.191) (2.817)
-TFT 7.788 6.762 9.664 7.828
+TCN 12.096 10.109 17.709 12.660
 (1.681) (0.538) (2.995) (0.826)
-NHITS 8.333 7.282 10.184 8.353
+NHITS 8.268 7.304 11.114 9.094
 (0.039) (0.055) (0.237) (0.236)
-TFT-Sum Total 6.698 7.634
+NHITS-ST 7.055 8.648
 (0.038) (0.068)
-NHITS-Pharmacokinetic 7.027 7.917
+NHITS-PK 7.037 8.492
 (0.093) (0.220)
-TFT-Pharmacokinetic 6.466 7.514
+RMSE
+ETS 14.751 21.454
 (0.00) (0.00)
-Table 6: Root Mean Square error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated data.
-Model All Values Critical Values
-Baseline Exogenous Baseline Exogenous
-ETS 15.754 23.151
+MLP 15.561 11.461 22.380 14.449
+(2.318) (0.553) (3.568) (0.502)
+RNN 19.033 27.116 28.834 43.256
+(3.479) (12.484) (6.974) (23.190)
+LSTM 18.567 13.723 26.863 18.028
+(11.191) (2.373) (18.642) (3.661)
+TCN 17.094 13.551 25.163 16.908
+(1.719) (0.628) (3.092) (1.074)
+NHITS 12.918 10.546 18.436 13.371
+(0.064) (0.106) (0.369) (0.387)
+NHITS-ST 10.082 12.679
 (0.054) (0.135)
-MLP 14.285 11.607 19.443 13.731
-(0.981) (1.756) (0.737) (2.321)
-LSTM 12.802 11.273 16.903 13.595
-(0.759) (1.918) (2.100) (3.260)
-TCN 22.874 16.194 36.468 22.533
-(11.937) (10.309) (22.454) (20.991)
-TFT 11.930 9.801 15.736 11.708
-(1.424) (0.690) (1.676) (0.948)
-NHITS 13.058 10.479 18.036 12.822
-(0.255) (0.372) (0.644) (0.855)
-TFT-Sum Total 9.711 11.283
+NHITS-PK 10.039 12.339
 (0.188) (0.435)
-NHITS-Pharmacokinetic 10.024 11.931
-(0.204) (0.549)
-TFT-Pharmacokinetic 9.407 11.224
-(0.654) (1.152)
-Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data.
+Table 7: Mean absolute error (MAE) (top) and Root Mean Square Error (RMSE) (bottom) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data.
 Average results are computed across 8 trials.
 Model All Values Critical Values
-Baseline Exogenous Baseline Exogenous
-ETS 10.126 11.781
+Baseline Exog. Baseline Exog.
+MAE
+ETS 9.022 10.028
 (0.00) (0.00)
-MLP 10.568 10.275 11.805 11.434
+MLP 11.065 11.120 12.630 12.698
 (1.461) (0.381) (1.727) (0.821)
-LSTM 10.604 9.606 12.999 11.144
+RNN 11.325 11.795 13.627 13.552
 (1.286) (1.985) (1.573) (2.480)
-TCN 11.011 11.090 12.693 13.212
+LSTM 9.811 9.467 11.252 11.178
 (0.570) (0.230) (0.512) (0.396)
-TFT 9.605 9.515 10.686 10.555
+TCN 11.525 10.760 13.647 12.678
 (2.008) (0.925) (3.349) (1.459)
-NHITS 9.234 8.954 10.436 10.117
+NHITS 9.060 8.873 10.337 10.137
 (0.142) (0.064) (0.137) (0.108)
-NHITS-Sum Total 8.988 10.092
+NHITS-ST 8.958 10.131
+(0.087 ) (0.100)
+NHITS-PK 8.758 9.965
 (0.080) (0.095)
-NHITS-Pharmacokinetic 8.769 10.007
+RMSE
+ETS 14.569 16.225
 (0.00) (0.00)
-TFT-Pharmacokinetic 9.492 10.484
+MLP 16.534 16.948 18.651 19.289
+(1.280) (0.481) (1.403) (0.943)
+RNN 16.498 17.066 19.446 19.423
+(1.252) (2.037) (1.709) (2.532)
+LSTM 14.915 14.597 16.894 16.851
+(0.494) (0.279) (0.490) (0.394)
+TCN 16.690 15.971 19.213 18.418
+(1.891) (0.710) (2.706) (1.295)
+NHITS 14.844 14.505 16.789 16.453
+(0.192) (0.096) (0.189) (0.132)
+NHITS-ST 14.640 16.492
 (0.205) (0.203)
-Table 8: Root mean square error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data.
-Model All Values Critical Values
-Baseline Exogenous Baseline Exogenous
-ETS 15.351 17.478
+NHITS-PK 14.385 16.251
 (0.092) (0.089)
-MLP 16.128 15.625 17.942 17.482
-(0.580) (0.131) (0.716) (0.232)
-LSTM 15.766 14.704 18.565 16.784
-(0.930) (0.297) (1.818) (0.688)
-TCN 16.173 16.192 18.432 18.978
-(0.932) (0.545) (1.097) (0.806)
-TFT 15.031 14.820 16.720 16.455
-(0.576) (0.186) (0.954) (0.413)
-NHITS 15.106 14.596 16.901 16.421
-(0.075) (0.080) (0.168) (0.133)
-NHITS-Sum Total 14.714 16.438
-(0.043) (0.135)
-NHITS-Pharmacokinetic 14.375 16.287
-(0.126) (0.154)
-TFT-Pharmacokinetic 14.846 16.351
-(0.366) (0.284)

### v8

@@
-Forecasting Treatment Response with Deep Pharmacokinetic Encoders
+Global Deep Forecasting with Patient-Specific Pharmacokinetics
 Abstract
 Forecasting healthcare time series data is vital for early detection of adverse outcomes and patient monitoring.
-However, forecasting is challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties for each patient.
+However, it can be challenging in practice due to variable medication administration and unique pharmacokinetic (PK) properties of each patient.
 To address these challenges, we propose a novel hybrid global-local architecture and a PK encoder that informs deep learning models of patient-specific treatment effects.
-We showcase the efficacy of our approach in achieving significant accuracy gains for a blood glucose forecasting task using both realistically simulated and real-world data.
-Our hybrid global-local architecture improves over patient-specific models by 15.8%  on average.
-Additionally, our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels.
-Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
-keywords:
-Multi-Input Forecasting, Healthcare, Pharmacokinetics, Deep Learning
+We showcase the efficacy of our approach in achieving significant accuracy gains in a blood glucose forecasting task using both realistically simulated and real-world data.
+Our PK encoder surpasses baselines by up to 16.4% on simulated data and 4.9% on real-world data for individual patients during critical events of severely high and low glucose levels.
+Furthermore, our proposed hybrid global-local architecture outperforms patient-specific PK models by 15.8%, on average.
+Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code
 1 Introduction
 Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution.
-Furthermore, while the medication dose is usually known, one cannot directly observe the plasma concentration of the drug, as illustrated in Fig.
+Furthermore, while the drug dose can be observed and recorded, the body’s interaction with the drug cannot be directly observed and is typically inferred from its plasma concentration, as illustrated in Fig.
 1.
-Consequently, the impact of dynamic changes in treatment resulting from drug absorption and elimination remains hidden.
-Figure 1: Insulin dose is usually known and recorded in data, whereas plasma insulin concentration is unobservable and varies over time.
-We consider the blood glucose forecasting task while taking into account latent, time-dependent treatment effects.
-Patients with diabetes must regularly monitor and regulate their blood glucose values to prevent hypoglycemic and hyperglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal.
+Consequently, the recorded drug dose alone cannot capture time-dependent drug kinetics, including processes such as absorption and elimination.
+Figure 1: Insulin dose is recorded, but plasma insulin concentration is not directly measurable and typically requires invasive procedures and laboratory tests.
+Consider the task of blood glucose monitoring for patients with diabetes who must regularly monitor and regulate their blood glucose level.
+Careful and prompt administration of insulin and intake of carbohydrates is required to prevent hyper- and hypoglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal.
 Hyperglycemic and hypoglycemic episodes are responsible for numerous hospital admissions each year.
-A 2009 report on hospital admissions for hyperglycemia indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009).
-Another report on hypoglycemic hospital admissions suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021).
+A 2009 report on hospital admissions indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009).
+Another report on hypoglycemic hospitalizations suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021).
 With over 38 million patients with Diabetes Mellitus in the United States alone, this equates to a conservative estimate of 212,800 patients hospitalized for hypoglycemic episodes each year.
-Capturing the dynamic effects of insulin is crucial for accurately predicting its impact on blood glucose levels to anticipate and proactively address abnormal levels to prevent adverse patient events.
+For healthcare challenges like this, advanced forecasting technologies can help patients and clinicians monitor, anticipate, and proactively address emerging abnormalities to prevent adverse events.
+However, accurate forecasting requires models that can leverage exogenous information, such as treatments, while capturing latent treatment kinetics and their effects to generate reliable forecast trajectories.
+To address this, we propose a hybrid model that integrates cohort-level information with patient-specific pharmacokinetic characteristics to improve the predictability of blood glucose responses to treatment.
 The main contributions of this paper are:
+  1. (i)
     Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific encoder.
+  2. (ii)
     Pharmacokinetic (PK) Encoder. We introduce a novel deep learning model-agnostic module that generates plasma concentration profiles to capture time-dependent treatment effects.
+  3. (iii)
     Multiple Dose Effects Framework.
 We provide a computationally efficient framework for encoding the effects of multiple doses.
-We verify this approach mathematically under the assumption of linear PK.
-  • State-of-the-Art Results. We obtain significant accuracy improvements on large scale simulated and real-world patient blood glucose forecasting datasets.
+We verify this approach mathematically under the assumption of linear pharmacokinetics.
+  4. (iv)
+    State-of-the-Art Results. We obtain significant accuracy improvements on large-scale simulated and real-world patient blood glucose forecasting datasets.
 1.1 Pharmacokinetics Background
-PK modeling is a natural solution to the aforementioned problem, as it can unveil medication effects through knowledge of drug movement within the body.
-For example, PK provides a well-established understanding of plasma insulin concentration changes over time based on drug absorption and elimination characteristics  (Binder et al., 1984).
+PK modeling is a natural solution to the aforementioned problem, as it can reveal the effects of the medication over time through knowledge of drug movement within the body.
+For example, PK provides a well-established understanding of the changes in plasma insulin concentration over time based on drug absorption and elimination characteristics  (Binder et al., 1984).
 Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves.
-The degree of skew varies with insulin type and patient-specific factors, as shown in Fig.
+The extent of skew varies with insulin type and patient-specific factors, as shown in Fig.
 2, where the area under the concentration curve measures drug bioavailability, or the fraction of the absorbed dose that reaches the intended site of systemic circulation intact.
@@
 Blood Glucose Forecasting.
-Neural Forecasting has previously been applied to the field of blood glucose forecasting, outperforming the accuracy of classic statistical methods.
-RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020).
-Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020; Jahangir et al., 2017) have been utilized in proposed models and as baselines.
-Blood Glucose Treatment Effect Modeling.
-Previous studies have successfully modeled treatment effects on patient blood glucose levels.
+Neural Forecasting has previously been used to predict blood glucose levels, outperforming traditional statistical methods.
+RNNs and LSTMs have been used both as models and as baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020).
+Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020; Jahangir et al., 2017) have been used in proposed models and as baselines.
+Modeling Blood Glucose Treatment Effects.
+Prior works have looked at modeling treatment effects on blood glucose levels.
 Miller et al.
-present one of the first hybrid statistical-and-physiological model that combines glucose-insulin dynamics and deep learning (Miller et al., 2020).
-They present the Deep T1D Simulator (DTD-Sim) model, which uses a deep state-space model to describe parameters representing insulin and carbohydrate absorption rates and sensitivities of the UVA/Padova simulator that simulates blood glucose levels.
+introduced one of the first hybrid statistical-physiological models, the Deep T1D Simulator (DTD-Sim), which integrates glucose-insulin dynamics with deep learning through a deep state-space model (Miller et al., 2020).
 Odnoblyudova et al.
-demonstrated enhanced accuracy in predicting blood glucose levels using Multi-output Gaussian Processes (MOGPs) to model composite effects of carbohydrate and fat components from meals.
-(Odnoblyudova et al., 2023).
-Despite recent advancements in forecasting treatment effects, significant challenges remain, including:
-Sparse data.
-Limited treatment data can hinder deep learning models from accurately learning treatment responses (Rubin-Falcone et al., 2022; Hussain et al., 2021).
-To address this challenge, we integrate domain knowledge into the deep learning models, allowing them to more effectively encode the time-dependent effects of these variables and improve forecasting performance.
-Composite treatment effects.
-Unlike meals that consist of various components such as carbohydrates and fats, insulin is a homogenous protein hormone composed of a single type of molecule.
-Therefore, modeling its component effects is not applicable (Odnoblyudova et al., 2023).
-Instead, we aim to model patient-specific holistic effects of the drug as well as composite effects of multiple doses.
-Inflexible models.
-Previous glucose forecasting models, like the UVA/Padova simulator, are specific to blood glucose dynamics and inflexible to other forecasting targets and variables (Visentin et al., 2016; Miller et al., 2020).
-Our approach leverages a deep learning model to generate predictions.
-This distinction provides greater flexibility, as our forecasting model is not constrained by a specific physiological model and can accommodate various deep learning models and forecasting targets and inputs.
-Model application scope.
-Prior work that integrates PK domain knowledge in deep learning models has been used in the context of predicting disease progression for cancer and COVID-19 with long-term horizons of interest (Hussain et al., 2021; Qian et al., 2021).
-In contrast, we leverage linear PK assumptions to create a framework that can be directly applied to various medications used in clinical practice and is easily adapted to other treatments.
-Individualized networks.
-Studies often use individualized networks, limiting cross-learning. Our novel global-local architecture allows efficient training of a more accurate global model applicable to multiple patients.
+improved the prediction of blood glucose using multi-output Gaussian processes (MOGP) to model the composite effects of meal components, such as carbohydrates and fats (Odnoblyudova et al., 2023).
+Unlike meals, insulin is a homogeneous protein hormone, making component-based modeling inapplicable to our task (Odnoblyudova et al., 2023).
+Instead, our objective is to capture the holistic patient-specific effects of the drug and the composite effects of multiple doses.
+Despite advances in treatment effect modeling, current approaches face limitations, including (1) reliance on individualized, subject-specific models and (2) inflexible architectures that hinder adaptability to various base models, forecasting tasks, and model inputs.
+The next paragraph outlines these limitations and how our proposed contributions address them.
+Previous work has developed individualized networks that overlook the benefits of developing global models that learn from various, diverse data series.
+Such cross-learning on diverse datasets can enhance generalization abilities, as demonstrated with Time Series Foundation Models (TSFMs) (Goswami et al., 2024).
+Our hybrid global-local architecture enables cohort-level training to develop a global model applicable to multiple patients while also preserving patient-specific information for personalized forecasts.
+Regarding architectures, models such as the UVA / Padova simulator depend on predefined ODEs for glucose dynamics, restricting adaptability to other targets and exogenous inputs (Visentin et al., 2016; Miller et al., 2020).
+Similarly, deep learning approaches for sparse exogenous variables often rely on fixed base models like LSTMs (Rubin-Falcone et al., 2022).
+Our proposed contributions support diverse deep learning models, and are flexible to forecasting targets and input variables, ensuring future adaptability as more advanced tools emerge.
+Other prior work that proposes the integration of PK knowledge with deep learning focuses on domains such as long-term disease progression in cancer and COVID-19, where such PK models are not directly applicable to blood glucose and insulin kinetics.
+(Hussain et al., 2021; Qian et al., 2021).
+Our study provides a formal framework for integrating PK in deep learning under linear PK assumptions that are applicable to many drugs used in clinical practice (Rosenbaum, 2017).
 3 Methods
 We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10 hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values.
-We choose a 10-hour duration to balance sufficient historical information with GPU memory constraints.
-the duration of the rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (ins, 2018, ).
-Therefore, the selection of a 10-hour lag time can fully cover multiple dose durations, allowing the model to learn treatment effects across various dosages and durations.
+We choose a 10-hour duration to balance sufficient historical information with our GPU memory constraints.
+Moreover, the duration of rapid-acting insulin (Humalog and Novolog) used in insulin pumps by patients in our cohort typically ranges from 2 to 5 hours (Clinic, 2018; Association, n.d.).
+Thus, a 10-hour timespan fully captures multiple dose durations, enabling the model to learn treatment effects across various dosages and durations.
 3.1 PK Encoder
-Medication administrations are represented as instantaneous, sparse variables in time series data.
+Medication administrations are represented as instantaneous sparse variables in time series data.
 We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes as shown in Fig.
@@
 $C(t,\textbf{x}^{(i)},\textbf{k}^{(i)})=\frac{\textbf{x}^{(i)}}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
-where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate that specific to patient, $i$.
+where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate specific to the patient, $i$.
 Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, x.
-The intuition for the log-normal shape of the concentration curves stems from pharmacokinetics literature as outlined in Section 1.1 and Appendix B.
+The intuition for the log-normal shape of the concentration curves stems from PK literature, as outlined in Section 1.1 and Appendix B.
 Figure 3: Medication administrations are represented as sparse variables in time series data.
 We propose a pharmacokinetic (PK) encoder to effectively capture time-dependent plasma drug concentration (Conc.).
-Under linear PK assumptions, PK parameters, including bioavailability, are assumed to be constant with respect to the drug.
+Under linear PK assumptions, the PK parameters, including bioavailability, are assumed to be constant with respect to the drug.
 Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.4.
-To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose, x, as a scaling factor in $C$.
+To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose x, as a scaling factor in $C$.
 This ensures that the area under the curve integrates to the dose value, under the simplifying assumption that all insulin administrations adhere to 100% bioavailability, which we show is upper bounded by the dose.
-Scaling the concentration curve changes its magnitude but does not alter the PK parameters or shape of the curve.
+Scaling the concentration curve changes its magnitude, but does not alter the PK parameters or shape of the curve.
 Let $N$ be the number of subjects in the data.
@@
 The encoder is trained end-to-end with the model.
-During model training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig.
+During training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig.
 4.
-The encoder than replaces sparse medication features in the training window with the corresponding concentration curve.
+The encoder then replaces sparse medication features in the training window with the corresponding concentration curves.
 For a single dose, x, of insulin administered at time, $t$, we obtain the concentration, $\widecheck{\textbf{x}}^{(i)}$, at that time for patient, $i$:
@@
 For each training step, $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
-For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded in the time prior to $t$ based on the lag time, $L$,
-$\hat{\textbf{y}}^{(i)}_{t:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
+For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded up to and including time $t$ based on the lag time, $L$,
+$\hat{\textbf{y}}^{(i)}_{t+1:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
 In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type.
 $\widecheck{\textbf{x}}_{t-L:t}$ includes concentration curves for basal insulin, $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin, $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO, $\widecheck{\textbf{x}}_{\text{CHO}}$.
-We include plasma treatment effects of CHO because, similar to drugs, it undergoes absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
-The PK encoder training framework is outlined in Appendix A.6, which can be integrated into any deep learning architecture capable of handling exogenous variables alongside modeling the target variable.
+We include treatment effects of carbohydrates because, similar to drugs, they undergo absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
 3.2 Modeling Multiple Dose Effects
 Multiple insulin doses may occur at close intervals in an event often referred to as “insulin stacking” as shown in Fig.
-11.
+3 and Fig.
+14.
 To model PK effects for multiple doses in an input window, we generate a concentration curve for each dose event and sum concentration curves across the windows.
@@
 Here, $f$ is the frequency of the data, $\textbf{x}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of treatment, and $\widecheck{\textbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$ is the new concentration curve feature for patient, $i$.
+The PK encoder training algorithm is described in Appendix A.5, which can be integrated into any deep learning architecture capable of handling exogenous variables in addition to modeling the target variable.
 Figure 4: Our hybrid global-local architecture combines global model parameters shared across patients with patient-specific pharmacokinetic (PK) parameters.
@@
 This variability in PK parameters can result in dose-dependent plasma drug concentration profiles.
-CGMs administer the same type of insulin (rapid-acting) for basal and bolus doses.
+Insulin pumps administer the same type of insulin (rapid-acting) for the basal and bolus doses.
 However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs.
-9 and 10 in Appendix A.1.
-To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity.
-By treating each dose range as a separate entity, we can capture potential PK variations without resorting to a fully nonlinear PK model.
+12 and 13 in Appendix A.1.
+To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity, so that we can capture potential PK variations without resorting to a fully nonlinear PK model.
 This method simplifies the modeling process by maintaining linear assumptions within each dose category.
-It avoids the complexity of a full nonlinear model while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics.
-Specifically, we estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values.
+It avoids the complexity while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics.
+Specifically, we separately estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values.
 For a single dose of basal insulin, $\textbf{x}_{\text{basal}}$, or bolus insulin, $\textbf{x}_{\text{bolus}}$, administered at time, $t$, we obtain the plasma insulin concentration at that time for patient, $i$:
@@
 $\widecheck{\textbf{x}}^{(i)}_{\text{bolus}}$ $=C(t,\textbf{x}^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)}).$ (7)
-Treating each dose as a different entity is an approximation and may not capture all nuances of dose-dependent kinetics.
-All physiological models are approximations of complex biological processes.
-Modeling dose-dependent effects in a more granular manner could potentially provide more precise results if such effects are present.
-However, our aim is to develop minimal models validated under established linear PK assumptions from the pharmacology literature, striking a balance between simplicity and accuracy.
-Our framework ensures broader applicability across various medications where a majority of drugs are assumed to adhere to linear PK (Rosenbaum, 2017).
-The equations presented in the paper are shown in a patient-specific manner to highlight that the PK encoder generates individualized information.
+Treating each dose category as a distinct entity is an approximation and may not capture all nuances of dose-dependent kinetics.
+However, all physiological models approximate complex biological processes.
+Our goal is to develop models validated under established linear PK assumptions from the pharmacology literature while still accounting for dose-dependent effects, balancing simplicity and precision.
+This framework also ensures broader applicability across various medications, as most drugs are assumed to follow linear PK (Rosenbaum, 2017).
+The equations presented are shown in a patient-specific manner to highlight that the PK encoder generates individualized information.
 However, these equations can be applied to produce patient-specific outputs for multiple patients simultaneously using vectorized operations.
 3.4 Models
-To assess forecasting performance across various model architectures, we trained models with six different algorithms as baselines.
-We employed Exponential Smoothing (ETS; Hyndman 2008) as our statistical baseline model.
-Our deep learning baselines included the following models: Multi-Layer Perceptron (MLP; Rosenblatt 1958), NBEATS/NHITS Oreshkin et al.
-(2020); Challu et al.
-(2022), Recurrent Neural Network (RNN; Elman (1990); Cho et al.
-(2014)), Long-Short Term Memory (LSTM; Sak et al.
-(2014)), and Temporal Convolution Network with an MLP Decoder (TCN; Bai et al.
+To assess forecasting performance across various model architectures, we trained global models with 11 different algorithms as baselines.
+Additionally, to assess our proposed hybrid global-local architecture, we compare the best performing model with the PK encoder to both purely local and global models.
+Statistical and Deep Learning Baselines.
+We employed Exponential Smoothing (AutoETS; Hyndman 2008) as our statistical baseline model.
+Our deep learning baselines include models from various architectural categories, such as Linear models, Multi-Layer Perceptron (MLP)-based, Recurrent Neural Network (RNN)-based, Convolutional Neural Network (CNN)-based, and Transformer-based models.
+We employ the following deep learning models: DLinear (Zeng et al., 2023), MLP (Rosenblatt, 1958), NHITS (Challu et al., 2022), NBEATSx (Oreshkin et al., 2020; Olivares et al., 2022a), RNN (Elman, 1990; Cho et al., 2014), Long-Short-Term Memory (LSTM; Sak et al.
+(2014)), Temporal Convolutional Network with an MLP decoder (TCN; Bai et al.
 (2018); van den Oord et al.
+(2016)), Informer (Zhou et al., 2021), patch time series Transformer (PatchTST; Nie et al.
+(2023)), and Temporal Fusion Transformer (TFT; Lim et al.
 (2021)).
-Implementations of these models were obtained from the Neuralforecast (Olivares et al., 2022) and StatsForecast (Garza et al., 2022) libraries.
-More information on these models and related prior work that leverages these are provided in Appendix A.2.
-Details on model training and hyperparameter selection are provided in Appendix A.3
-3.5 Baselines
-We compare our hybrid global-local architecture with the PK encoder to both local and global baselines:
-Global Baseline Models.
-Models were trained on all patient data to create a single global model:
-  • Baseline. Models are trained with the univariate blood glucose signal to demonstrate the forecasting improvement from using exogenous variables, supporting our PK encoder.
-  • Baseline Exogenous (Exog.). Models are trained with the univariate blood glucose signals, static features, and sparse exogenous features (CHO and insulin dose values).
-  • Baseline Exogenous (Exog.) - ST.
-Models are trained with the univariate blood glucose signal, static features, and sparse exogenous features transformed via the sum total (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022).
-Local Baseline Exogenous Models.
-Patient-specific baseline exogenous models were compared against our hybrid global-local PK encoder model.
-Local PK Encoder Models.
-Patient-specific models using the PK encoder were compared against our hybrid global-local PK encoder model.
-3.6 Evaluation
+More information on these models and related prior work is provided in Appendix A.2.
+Details on model training and hyperparameter selection are provided in Appendix A.3.
+Global Model Baselines.
+Global models are trained on all patient data to create a single model.
+Global model baselines include: (1) univariate models trained with only the blood glucose signal and static features, (2) models trained with univariate blood glucose signals, static features, and sparse historical exogenous features (CHO and insulin dose values), and (3) models trained with the univariate blood glucose signal, static features, and sparse historical exogenous features transformed using the “Sum Total” (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022).
+Univariate models are included as baselines to highlight forecast improvements from training on historical exogenous variables, motivating advanced transformations for sparse, continuous exogenous features using methods like our PK encoder.
+With the exception of TFT, most Transformer-based models, including Informer and PatchTST, are not designed to model covariates.
+For these two models, we can only report global model results for those trained on univariate blood glucose data.
+Since our proposed PK encoder is intended to supplement model architectures capable of incorporating exogenous variables alongside the endogenous signal, we focus on evaluating the PK encoder in top-performing models that are capable of modeling covariates, such as TFT, NBEATSx, and NHITS.
+Beyond modeling covariates, residualized MLP-based models like NBEATSx and NHITS are valuable to demonstrate the utility of our PK encoder as NHITS has achieved SOTA performance, outperforming RNNs and Transformers in long-range forecasting tasks while avoiding the quadratic complexity of self-attention in Transformer-based architectures (Challu et al., 2022; Ansari et al., 2024).
+Local Model Baselines.
+Patient-specific sparse exogenous models as well as patient-specific exogenous models using the PK encoder were compared against our hybrid global-local PK encoder model.
+3.5 Model Training
+Models were trained using Adam optimizer (Kingma and Ba, 2017).
+Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected.
+Huber estimator was selected as the loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
+$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)=\left\{\begin{array}[]{ll}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{array}\right.$ (8)
+Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set.
+Optimal hyperparameters were selected based on a validation set with cross-validation.
+More information on model training and hyperparameters in provided in Appendix A.3.
+3.6 Model Evaluation
 Trained models were used to generate rolling window forecasts at one-step intervals.
-Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate the blood glucose forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE) computed across 8 trials of individually trained models.
-Formulations for these metrics are provided in Appendix A.4.
-We evaluate model performance across all points within the specified time horizon.
-We also assess the model’s performance exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL).
+Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE):
+$\mathrm{MAE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\frac{1}{H}\sum_{{\tau}=t+1}^{t+H}|\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau}|\quad\text{and}$ (9)
+$\mathrm{RMSE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\sqrt{\frac{1}{H}\sum_{\tau=t+1}^{t+H}(\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau})^{2}}.$ (10)
+We compute results across all points within the forecast horizon, and present the average results over 8 trials of individually trained models.
+We also evaluate the performance of the models exclusively at time points where blood glucose levels reach or fall below critical thresholds relevant to hypoglycemic and hyperglycemic events (i.e., $\leq$70 mg/dL and $\geq$180 mg/dL).
 For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation.
+To evaluate whether hybrid global-local PK models significantly outperform baselines, we use paired t-test statistics computed on the patients’ mean MAE results.
 4 Cohort
-Simulated data
+We use simulated and real-world datasets containing patient insulin, CHO time series, and blood glucose measurements.
+These open-source datasets ensure reproducibility and are widely used in blood glucose forecasting research, featuring some of the largest patient cohorts (Xie and Wang, 2020; Rubin-Falcone et al., 2020, 2022; Rabby et al., 2021; Li et al., 2019b).
+Simulated data.
 An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides blood glucose, CHO, and insulin values for 30 simulated patients with Type-I diabetes.
 We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set.
-Blood glucose time series plots for 6 subjects are shown Fig.
-7 in Appendix A.1.
-Basal and bolus dose distributions for 12 simuluted patients are shown in Fig.
-9 in Appendix A.1.
-OhioT1DM
-12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020).
-Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes.
-The data includes basal rate information, bolus insulin (units), and CHO values.
+Appendix A.1 contains examples of blood glucose time series for 6 individuals, shown in Fig.
+10, as well as basal and bolus dose distributions for 12 individuals, shown in Fig.
+12.
+Figure 5: The hybrid global-local NHITS-PK model trained on all OhioT1DM individuals (orange, solid median) has a lower average MAE than NHITS-PK local models (blue, dashed median).
+Figure 6: The NHITS-PK hybrid global-local model (orange, solid median) consistently outperforms the local NHITS-PK model (blue, dashed median).
+[…]
+Simulated Dataset OhioT1DM Dataset
+Model All Values Critical Values All Values Critical Values
+w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
+AutoETS 9.450 13.457 9.022 10.028
+DLinear 7.802 7.989 10.575 10.790 11.514 13.011 12.512 14.580
+RNN 13.703 20.190 21.433 35.955 11.325 11.795 13.627 13.552
+LSTM 13.557 10.321 20.537 13.550 9.811 9.467 11.252 11.178
+TCN 12.096 10.109 17.709 12.660 11.525 10.760 13.647 12.678
+Informer 8.599 12.074 13.714 14.867
+PatchTST 7.284 9.842 9.922 10.999
+TFT 6.587 5.826 8.518 7.132 9.402 9.197 10.577 10.643
+MLP 10.589 8.433 14.699 10.600 11.065 11.120 12.630 12.698
+NBEATSx 9.187 6.886 12.757 8.145 9.868 10.280 11.037 11.359
+NHITS 8.268 7.304 11.114 9.094 9.060 8.873 10.337 10.137
+TFT-PK 5.743 7.042 9.529 10.461
+NBEATSx-PK 6.806 7.998 9.983 10.927
+NHITS-PK 7.037 8.492 8.758 9.965
+Real-world data.
+Twelve de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020).
+Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a continuous glucose monitor (CGM) at a frequency of 5-minutes, basal rate information, bolus insulin (units), and CHO values.
 The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days.
 Missing values were forward filled to prevent data leakage.
-The sample count and percent missing data for each subject are included in Table 2 in Appendix A.1.
-Blood glucose time series plots for 6 subjects are shown in Fig.
-8 in Appendix A.1.
-Basal and bolus dose distributions for the 12 patients are shown in Fig.
-10 in Appendix A.1.
-More dataset information is in Appendix A.1.
+The sample count and percentages of missing data for each subject are included in Table 4 in Appendix A.1.
+Appendix A.1 also contains examples of blood glucose time series plots for 6 subjects, shown in Fig.
+11, as well as basal and bolus dose distributions for the 12 patients, shown in Fig.
+13.
 5 Results
-5.1 Results on Simulated Data
-For the simulated dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous model for forecasts evaluated at all values (t-test p-value: 6.6e-6) as shown in as shown in Table 1.
-For evaluations across all values, the NHITS-PK model provides 3.5% improvement, on average, and up to 9.2% improvement for individual patients over the baseline NHITS sparse exogenous model.
-For evaluations across critical values, the NHITS-PK model provides 5.2% improvement, on average, and up to 16.4% improvement for individual patients over the baseline NHITS sparse exogenous model.
-Additionally, the NHITS-PK model outperforms the NHITS-ST model.
-Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 6 in Appendix C.5.
-5.2 Results on Real Data
-Our hybrid global-local architecture produces models that significantly outperform local models by 15.8%, on average (t-test p-value: 1.7e-5), and up to 31.3% improvement for individual OhioT1DM patients.
-We show that the NHITS-PK hybrid global-local model outperforms NHITS-PK models developed for individual OhioT1DM patients in Fig.
-12.
-Moreover, this improvement in forecast accuracy is consistent across all patients as shown in Fig.
-5 and in Table 4 in Appendix C.1.
-Additionally, the NHITS-PK model with the hybrid global-local architecture also improves MAE by 14.8%, on average, for patients over the patient-specific NHITS sparse exogenous baseline models (t-test p-value: 1.7e-6).
-For the OhioT1DM dataset, the NHITS-PK model has significantly smaller MAE than the NHITS sparse exogenous baseline model for forecasts evaluated at all values (t-test p-value: 5.0e-3) and critical values (t-test p-value: 3.5e-3) as shown in Table 1.
-For evaluations across all values, The NHITS-PK model provides 1.3%  on average, and up to 3.5% improvement for individual patients over the baseline NHITS sparse exogenous model.
-For evaluations across critical values, the NHITS-PK provides 1.6%  on average, and up to 4.9% improvement for individual patients over the baseline NHITS sparse exogenous model.
-Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Table 7 in Appendix C.5.
-Fig.
-13 in Appendix C.2 shows model forecasts during hyperglycemic events for four randomly selected OhioT1DM patients.
-Figure 5: The NHITS-PK hybrid global-local model (blue, solid) provides more accurate predictions than local NHITS-PK models (purple, dashed).
-Figure 6: In addition to reducing forecasting error during hyperglycemic events, our NHITS-PK encoder enables improved lead time predictions of hypertensive events (blood glucose $\geq$ 180) for 9 of 12 OhioT1DM patients as measured over the 8 trials.
-Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
-Simulated Dataset
-Model All Values Critical Values
-Baseline Exog. Baseline Exog.
-ETS 9.450 13.457
-MLP 10.589 8.433 14.699 10.600
-RNN 13.703 20.190 21.433 35.955
-LSTM 13.557 10.321 20.537 13.550
-TCN 12.096 10.109 17.709 12.660
-NHITS 8.268 7.304 11.114 9.094
-NHITS-ST 7.055 8.648
-NHITS-PK 7.037 8.492
-OhioT1DM Dataset
-ETS 9.022 10.028
-MLP 11.065 11.120 12.630 12.698
-RNN 11.325 11.795 13.627 13.552
-LSTM 9.811 9.467 11.252 11.178
-TCN 11.525 10.760 13.647 12.678
-NHITS 9.060 8.873 10.337 10.137
-NHITS-ST 8.958 10.131
-NHITS-PK 8.758 9.965
+PK-encoder models outperform their sparse exogenous model counterparts.
+For simulated data, the TFT-PK model achieves the lowest MAE, on average, among all models for both evaluations across all models and critical values.
+The PK models, including TFT-PK, NBEATSx-PK, and NHITS-PK outperform their sparse exogenous model counterparts by up to 5.4%, 3.7%, 9.2%, respectively, on evaluations across all values and 7.7%, 6.6%, 16.4% on evaluations across critical values for individual subjects.
+For the OhioT1DM dataset, the NHITS-PK model achieves the lowest MAE, on average, among all models for both evaluations across all values and critical values.
+The PK models, including, TFT-PK, NBEATSx-PK, and NHITS-PK, outperform their sparse exogenous model counterparts by up to 5.2%, 8.5%, 4.9% on evaluations across critical values for individual patients.
+Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Appendix C.3 Tables 8 and 9, respectively.
+Hybrid global-local models significantly improve forecasts over global models.
+For the simulated dataset, the TFT-PK, NBEATSx-PK, and NHITS-PK models achieve significantly lower forecast errors for subjects for evaluations on all values (p-values: 4.6e-4 [TFT-PK]; 7.9e-5 [NBEATSx-PK]; 2.0e-7 [NHITS-PK]) and critical values (p-values: 5.7e-4 [NBEATSx-PK]; 3.0e-5 [NHITS-PK]) compared to their sparse exogenous global model counterparts.
+For the OhioT1DM dataset, the NBEATSx-PK and NHITS-PK models also achieve significantly lower forecast errors for subjects for both evaluations on all values (p-values: 3.4e-3 [NBEATSx-PK]; 1.3e-3 [NHITS-PK]) and critical values (p-values: 4.4e-4 [NBEATSx-PK]; 7.0e-3 [NHITS-PK]) compared to their sparse exogenous global model counterparts.
+Additionally, Table 2 shows that TFT-PK and NHITS-PK models achieve significantly lower errors, on average—2.4% (p-value: 3.6e-12) and 2.3% (p-value: 4.1e-5), respectively—for subjects in evaluations across all values, compared to their “Sum Total” model counterparts.
+Hybrid global-local models significantly improve forecasts over local models.
+For the OhioT1DM dataset, the NHITS-PK hybrid global-local model outperforms local NHITS-PK models developed for individual OhioT1DM patients by 15.8%, on average (p-value: 1.7e-5) as shown in Fig.
+5.
+Furthermore, this improvement in forecast accuracy is consistent across all patients as shown in Fig.
+6 and Table 3 (standard deviations are provided in Table 6 in Appendix C.1).
+Our hybrid global-local NHITS-PK model also significantly outperforms local sparse exogenous baseline models by 14.8%, on average, for individual subjects (p-value: 1.7e-6).
+Table 2: PK models outperform Sum Total (ST) models for both Simulated and OhioT1DM datasets. Best result is in bold.
+Simulated Dataset OhioT1DM Dataset
+TFT-PK TFT-ST NHITS-PK NHITS-ST
+All Values 5.743 5.885 8.758 8.958
+(0.159) (0.187) (0.080) ((0.087)
+Critical Values 7.042 7.233 9.965 10.131
+(0.270) (0.282) (0.095) (0.100)
+Table 3: MAE computed across forecast horizon values for local NHITS, global NHITS, local NHITS-PK, and hybrid global-local NHITS-PK on the 12 subjects from the OhioT1DM dataset. Best results in bold.
+ID w/ Exog. (Sparse) w/ Exog. (PK)
+Local Global Local Hybrid
+540 12.106 10.119 11.880 10.015
+544 7.952 7.188 8.067 6.967
+552 8.559 7.400 9.132 7.374
+559 10.845 9.013 13.051 8.967
+563 9.527 8.731 9.427 8.506
+567 12.599 9.640 12.132 9.562
+570 8.361 7.523 8.466 7.258
+575 10.891 9.712 11.095 9.496
+584 11.111 9.846 11.518 9.905
+588 10.172 8.172 9.376 8.137
+591 11.987 10.246 11.428 10.151
+596 8.990 8.133 9.147 8.059
+All 10.324 8.873 10.425 8.758
+The NHITS-PK model effectively learns the inhibitory treatment effects of insulin on blood glucose.
+We examine whether the NHITS-PK model can accurately capture treatment effects of insulin on blood glucose by generating forecasts for three scenarios: the original bolus insulin doses, bolus insulin doses removed, and bolus insulin doses increased by a factor of 10.
+When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig.
+8.
+This result underscores the importance of models that can effectively capture treatment effects, enabling them to adapt to new treatment doses for more reliable and accurate forecasts.
+Figure 7: Blood glucose forecasts of the final horizon point during 5 randomly selected hyperglycemic periods (blood glucose $\geq$ 180) for 4 randomly selected OhioT1dm patients.
+Forecasts for the NHITS-PK model (orange, dashed) are more aligned with ground truth (black, solid) compared to the NHITS sparse exogenous model blue, dash-dotted).
+Figure 8: The NHITS-PK model captures insulin treatment effects: forecasts with removed bolus doses (yellow, dash-dotted) and augmented doses (red, dotted), result in higher and lower glucose predictions, respectively, compared to the original doses (orange, dashed).
+Each point represents the final forecasted timestamp from a rolling window forecast.
 6 Discussion
-Hybrid Global-Local Architecture We demonstrate the efficacy of our novel hybrid global-local model, which produces forecasts with significantly smaller errors than patient-specific models across all patients.
-Our hybrid architecture not only improves accuracy but also enables global models to learn relevant patient-specific information, outperforming global model baselines.
-A hybrid global-local model offers practical benefits beyond improved forecasting accuracy as it requires less computational resources and time to train than individual patient models.
-Training patient-specific NHITS-PK models took seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space.
-The hybrid model can also be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
-PK Encoder As a model-agnostic extension, our novel PK encoder significantly reduces forecasting errors in deep learning models for both simulated and real-world data.
+The proposed PK encoder can substantially reduce forecasting errors in deep learning models for both simulated and real-world data.
 These improvements are not only statistically significant, but they have potential for positive real-world clinical impact.
-In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts such events earlier, up to 3.6 minutes when compared to the NHITS baseline model, for individual OhioT1DM patients as shown in Fig.
-6.
+In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts these events earlier, up to 3.6 minutes compared to the NHITS sparse exogenous baseline model, for individual OhioT1DM patients, as shown in Fig.
+9.
 This work uses data from T1DM patients but incorporates key variables relevant to both T1DM and T2DM, such as continuous glucose measurements (CGM), carbohydrate intake, and insulin doses, making our method applicable to both populations.
-Improving prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009).
-Our results also highlight the importance of using domain knowledge to incorporate exogenous variables into models.
-Models with sparse features may underutilize valuable information, and feature-engineered representations not informed by domain knowledge, like the ‘sum total’ approach, can underperform compared to baselines.
-Moreover, Appendix C.4 shows that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset.
-As such, dose-dependent models, such as proposed in section 3.3, may be needed for more realistic data simulations and models.
-Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing or drug selection.
-It can also be adapted for other medication administration types, such as oral or intravenous injections.
-Limitations.
-We equate $\sigma$ to our k parameter in the log-normal function and hold $\mu$ constant at 1.
-This simplification reduces the number of potential concentration curve shapes the model can learn.
-Tuning both $\sigma$ and $\mu$ could further improve forecasts by enabling more flexible concentration curves.
-Future Work.
-Future work will explore fine-tuning global forecasting models with patient-specific parameters for new patients with limited historical data as well as modeling dose-dependent effects using more dynamic nonlinear PK parameter frameworks.
-Future work will include applying our approach to additional patient cohorts, such as those with T2DM, to further validate our findings.
+Improving the prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009).
+Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing.
+While we demonstrate performance improvements across the top three models, our approach can be easily integrated into other deep learning architectures as new tools emerge.
+It can also be easily adapted for other types of medication administration, such as oral or intravenous injections.
+In addition to improving performance over global model counterparts, our hybrid global-local model architecture produces forecasts with significantly smaller errors than patient-specific models.
+This result is of substantial importance in demonstrating that global models can integrate cohort-level information with patient-specific parameters to produce more accurate personalized forecasts.
+Moreover, our proposed hybrid global-local model offers practical benefits beyond improved forecasting performance as it requires less computational resources and time to train than individual patient models.
+Training patient-specific NHITS-PK models took approximately seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space.
+The hybrid model also has the potential to be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
+Our study demonstrates that leveraging domain knowledge to integrate exogenous variables in models offers performance benefits over using sparse features or the “Sum Total” feature engineering approach, both of which are outperformed by our proposed deep PK encoding method.
+Furthermore, dose-dependent PK encoding strategies, such as those proposed in Section 3.3, can lead to more realistic models, as highlighted by our ablation study in Appendix C.2 where we demonstrate that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset.
+Limitations and Future Work.
+This study has two limitations related to the PK encoder: we assume that the concentration curves exhibit 100% bioavailability and maintain a consistent peak concentration across patients.
+These simplifying assumptions are intended to focus model learning on the parameter $k$, rather than on the area or peak of the concentration curve, to emphasize the spread or duration of the concentration curve and to reduce potential issues with parameter identification.
+Future work will explore expanding concentration curve parameters to estimate area and peak parameters to potentially enable more flexible concentration curves.
+We will also aim to apply our approach to additional cohorts of patients, such as those with T2DM, as more data become available, to further validate our findings.
+Future work will also explore modeling dose-dependent effects using more dynamic nonlinear PK frameworks.
+Figure 9: The PK encoder enables improved NHITS model prediction lead times for hyperglycemic events (CGM $\geq$ 180) for 9 of 12 OhioT1DM patients.
 […]
 This work was partially supported by the NIH award R01HL141916. The authors wish to thank Dr. Frank Guyette, Dr. Leonard Weiss and Dr. Gilles Clermont for their guidance and informative conversations.
-Appendix A Supplemental Methods
 Data and Code Availability
 We use open-source simulated and real-world data that is available to other researchers.
-The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository [Xie, 2018, Visentin et al., 2016].
-The real-world data includes the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020].
-Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu [2020].
-Implementations of models used in our work were obtained from the open-source Neuralforecast [Olivares et al., 2022] and StatsForecast [Garza et al., 2022] libraries.
-We implemented our method using a copy of the Neuralforecast repository to leverage their model training framework.
+The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository (Xie, 2018; Visentin et al., 2016).
+The real-world data includes the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020).
+Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu (2020).
+Implementations of models used in our work were obtained from the open-source Neuralforecast (Olivares et al., 2022b) and StatsForecast (Garza et al., 2022) libraries.
+We implemented our method by adopting the Neuralforecast repository to leverage their model training framework.
 Our research code, along with the Neuralforecast models used in our research, is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
+Appendix A Supplemental Methods
 A.1 Data
 Simulated data
-An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) [Xie, 2018, Visentin et al., 2016] provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes.
+An open-source python implementation of the FDA-approved UVa/Padova Simulator (2008 version) (Xie, 2018; Visentin et al., 2016) provides clinical and biological parameters for 30 patients (10 adults, 10 adolescents, and 10 children) with Type-I diabetes.
 Clinical parameters include information on age and weight, and biological parameters include information on insulin-glucose kinetics, such as absorption constants.
 We generate 54 days of data for each patient to match the median training set of the OhioT1DM, and approximately 9 days for the test set.
-In a similar approach to [Rubin-Falcone et al., 2022], the meal schedule used to generate simulated data was based on the Harrison-Benedict equation [Arthur Harris and Gano Benedict, 21919] with approximately 3 meals per day and no additional snacks.
+In a similar approach to (Rubin-Falcone et al., 2022), the meal schedule used to generate simulated data was based on the Harrison-Benedict equation (Arthur Harris and Gano Benedict, 21919) with approximately 3 meals per day and no additional snacks.
 Height data was not provided in the patient parameters, and so was estimated at 140cm, 170cm, and 175cm for adults, adolescents, and children, respectively.
@@
 OhioT1DM
-12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets [Marling and Bunescu, 2020].
+12 de-identified individuals with Type 1 diabetes are included in the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020).
 Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes.
@@
 Figs.
-9 and 10 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively.
+12 and 13 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively.
 Table 4 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
@@
 The baseline models used in this work have been extensively used in previous blood glucose forecasting work.
-For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies [Qian et al., 2021, Li et al., 2019a, Fox et al., 2018, Rabby et al., 2021, Rubin-Falcone et al., 2020, 2022, Miller et al., 2020].
-Similarly, CNNs [Xie and Wang, 2020, Li et al., 2019a, b] and MLPs [McShinsky and Marshall, 2020, Miller et al., 2020] have been utilized in proposed models and as baselines.
-NHITS achieves state-of-the-art performance, surpassing RNN-based, transformer-based, and other deep learning time series decomposition models (such as N-BEATS [Oreshkin et al., 2020]) in long-horizon forecasting tasks [Challu et al., 2022].
-Exponential Smoothing (ETS) The ETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. [Hyndman, 2008].
-Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation [Nair and Hinton, 2010, Fukushima, 1975, Rosenblatt, 1958].
+For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020).
+Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020) have been used in proposed models and as baselines.
+Exponential Smoothing (AutoETS) The AutoETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. (Hyndman, 2008).
+DLinear (DLinear) A Linear model that leverages linear layers to model the trend and seasonal components.
+A moving average filter is used to separate the time series into its trend and seasonal components (Zeng et al., 2023).
+Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
 Neural Hierarchical Interpolation for Time Series (NHITS) A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands.
-The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation [Challu et al., 2022].
+The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation (Challu et al., 2022).
+Neural Basis Expansion Analysis with Exogenous (NBEATSx) An MLP-based deep learning model that decomposes the input signal into trend and seasonality components using polynomial and harmonic basis projections.
+The NBEATSx extension enables modeling of exogenous variables in addition to the univariate signal (Oreshkin et al., 2020; Olivares et al., 2022a).
 Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states.
-The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts.[Elman, 1990, Cho et al., 2014]
-Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.[Sak et al., 2014]
+The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts (Elman, 1990; Cho et al., 2014).
+Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.(Sak et al., 2014)
 Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts.
-To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions [Bai et al., 2018, van den Oord et al., 2016].
+To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions (Bai et al., 2018; van den Oord et al., 2016).
+Informer (Informer) (Zhou et al., 2021) A Transformer-based deep learning model with a multi-head attention mechanism that uses autoregressive features from a convolution network.
+A ProbSparse self-attention mechanism and a self-attention distilling process are leveraged to reduce computational complexity.
+Patch time series Transformer (PatchTST) (Nie et al., 2023) A Transformer-based deep learning model with a multi-head attention mechanism.
+Input time series into sub-series level patches as input tokens.
+Temporal Fusion Transformer (TFT) An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies.
+TFT also leverages variable selection networks as well as a series of gating layers to suppress unnecessary parts of the architecture (Lim et al., 2021).
 A.3 Model Training and Hyperparameters
-Models were trained using Adam optimizer [Kingma and Ba, 2017].
-Optimal hyperparameters were selected based on a validation set with cross-validation.
-Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected.
-Huber estimator was selected as a robust loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
-$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)$ =$=\begin{cases}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta\cdot(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{cases}$ (8)
-Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set.
 The deep learning models were trained using an NVIDIA A100 Tensor Core GPU.
@@
 We tune additional specific hyperparameters for each architecture.
-For the MLP  we tune the number of hidden units per layer ([128,256]); for the LSTM  the hidden size ([128,256]); for the TCN the hidden size ([128,256]) and number of dilations ([2,4,8]); for the NHITS we did not tune additional hyperparameters.
+For the MLP model, we tune the number of hidden units per layer ([128, 256]).
+For the LSTM model, we tune the hidden size ([128, 256]).
+For the TCN model, we tune the hidden size ([128, 256]) and number of dilations ([2, 4, 8]).
+For the PatchTST model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and patch length ([8, 16, 32, 64, 96, 128]).
+For the Informer model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and number of decoder layers ([1, 2]).
+For the TFT model, we tune the hidden size ([128, 256, 512]) and number of attention heads ([4, 8, 16]).
+For the NBEATSx and NHITS models, we did not tune additional hyperparameters.
 The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
-A.4 Metrics
-We use the following metrics to evaluate model performance. Here, $H$ refers to the forecast horizon.
-Mean Absolute Error (MAE)
-$MAE(y_{t+1,t+H},\hat{y}_{t+1,t+H})=\frac{1}{H}\sum_{i=t+1}^{t+H}|y_{i}-\hat{y}_{i}|$
-Root Mean Square Error (RMSE)
-$RMSE(y_{t+1,t+H},\hat{y}_{t+1,t+H})=\sqrt{\frac{1}{H}\sum_{i=t+1}^{t+H}(y_{i}-\hat{y}_{i})^{2}}$
-A.5 Propositions
+A.4 Propositions
 Proposition A.1.
@@
 Here, $T$ represents the total duration of drug exposure or the time interval over which drug concentrations are being measured and $t$ represents a specific point in time within that interval.
-A.6 Algorithm
+A.5 Algorithm
 We outline the algorithm for training our proposed PK encoder that transforms sparse exogenous variables into time-dependent concentration profiles.
@@
 The algorithm initializes k as a network embeding weight using a value $\textbf{k}^{*}$ from the specified range $[0,1]$ at the beginning of each epoch.
-This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library [Liaw et al., 2018].
+This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library (Liaw et al., 2018).
 For each training step, the algorithm iterates over batches of data.
@@
                $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}_{t},\textbf{k})$
-$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
-$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
+$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
+$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
 $\textbf{k}\leftarrow\textbf{k}-\eta\cdot\nabla_{\textbf{k}}\ell(\textbf{k})$triangleright Update k using SGD
@@
                      $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}_{t},\textbf{k})$
-$\hat{\textbf{y}}_{t:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
-$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t:t+H},\hat{\textbf{y}}_{t:t+H})$
+$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
+$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
              end for
@@
 Concentration-time Profiles
-From PK literature, exponential decay functions are commonly used to model drug concentration over time for intravenous injections, where the drug is directly injected into the bloodstream, leading to instantaneous effects [Rosenbaum, 2017, Angelika et al., 2005].
+From PK literature, exponential decay functions are commonly used to model drug concentration over time for intravenous injections, where the drug is directly injected into the bloodstream, leading to instantaneous effects (Rosenbaum, 2017; Angelika et al., 2005).
 Alternatively, subcutaneous injections involve absorption and elimination processes where the peak onset time may not be instantaneous depending on the drug profile.
@@
 Fig.
-2 is adapted from PK literature and depicts PK modeling of plasma insulin concentration over time for different insulin analogs [Angelika et al., 2005].
+2 is adapted from PK literature and depicts PK modeling of plasma insulin concentration over time for different insulin analogs (Angelika et al., 2005).
 Generally, bolus doses are quick-acting and, thus, often taken before meals.
@@
 Bioavailability varies based on many factors, including type of medication, dosage, and patient-specific characteristics.
-For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent [Søeborg et al., 2012].
+For instance, the bioavailability of an intravenous dose of any drug is close to 100 percent (Søeborg et al., 2012).
 However, for subcutaneous injections, bioavailability is typically lower due to unintended drug absorption or elimination.
@@
 Linear Pharmacokinetics
-Linear pharmacokinetics refers to a situation where the relationship between the dose of a drug and its concentration in the body is linear [Rosenbaum, 2017].
+Linear pharmacokinetics refers to a situation where the relationship between the dose of a drug and its concentration in the body is linear (Rosenbaum, 2017).
 In linear pharmacokinetics, the area under the concentration curve is directly proportional to the dose of the drug.
@@
 Linearity results from the assumption that the drug processes exhibit first-order kinetics and that the PK parameters remain constant with dose, indicating no significant changes in the body’s ability to eliminate the drug with higher doses.
-The majority of drugs used in clinical practice are assumed to follow linear pharmacokinetics [Rosenbaum, 2017].
+The majority of drugs used in clinical practice are assumed to follow linear pharmacokinetics (Rosenbaum, 2017).
 Figure 14: When doses of insulin are administered at close intervals, the subsequent dose may start to act before the previous dose has fully finished its effect.
@@
 6 and Fig.
+6.
+Furthermore, the global NHITS-PK model consistently outperforms the global NHITS-Exogenous model, on average, for OhioT1DM subjects as shown in Fig.
 5.
-Furthermore, the global NHITS-PK model consistently outperforms the global NHITS-Exogenous model, on average, for OhioT1DM subjects as shown in Fig.
-12.
-Table 4: MAE computed across forecast horizon values using local and global models.
-ID Exogenous PK
+Table 6: MAE computed across forecast horizon values using local and global models for the OhioT1DM dataset. Best results are in bold.
+ID w/ Exog. (Sparse) w/ Exog. (PK)
 Local Global Local Global
@@
 (0.401) (0.064) (0.659) (0.080)
-Figure 12: The NHITS-PK hybrid global-local model (Hybrid) trained on all OhioT1DM subject data (blue, solid) has a lower MAE, on average, than NHITS-PK models trained on individual subjects, or local models (purple, dashed) .
-C.2 Forecast Comparison
-Figure 13: Blood glucose forecasts of the final horizon point during 5 randomly selected hypoglycemic threshold periods (blood glucose $\geq$ 180) for 4 randomly selected OhioT1dm patients.
-The NHITS-PK hybrid global-local model (orange, dashed) produces improved forecasts of the ground truth (blue, solid) compared to the NHITS baseline model with sparse exogenous features (yellow, dash-dotted).
-C.3 Forecasting Response to Treatment
-Predicting treatment effects using accurate forecasting algorithms can assist individuals with diabetes and their care providers in making informed decisions about the ideal medication type and dosage, leading to improved glycemic control and reduced risk of both short- and long-term complications.
-Assessing a forecasting model’s ability to effectively capture the impact of variables on the target signal is essential for ensuring the reliability of forecasts.
-We examine whether the NHITS-PK model can accurately capture the treatment effects of insulin on blood glucose by generating forecasts during time periods for three scenarios: with the original bolus insulin doses, with bolus insulin doses removed, and with bolus insulin doses increased by a factor of 10.
-When insulin doses are removed, blood glucose predictions increase, and when insulin doses are augmented, blood glucose predictions decrease as shown in Fig.
-14.
-The NHITS-PK model thus effectively learns the inhibitory treatment effects of insulin on blood glucose.
-Figure 14: Treatment effect of insulin on the NHITS-PK hybrid global-local model.
-Each point in the plot corresponds to the final horizon point.
-Forecasts with original bolus doses (orange, solid), removing bolus insulin doses results in higher forecasts (blue, dashed), and augmenting doses results in lower blood glucose forecasts (yellow, dotted).
-C.4 Forecasting with Dose-Independent Parameters
-[…]
-We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values.
-However, this approach led to worse performance on the OhioT1DM dataset.
-The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters [Xie, 2018, Visentin et al., 2016].
-In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3.
-Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data.
-These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
-Table 5: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180)
+C.2 Forecasting with Dose-Independent Parameters
+Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180).
+Result for models with (w/) and without (w/o) exogenous variables are shown.
+Best results are in bold.
 Simulated Dataset
 Model All Values Critical Values
-Baseline Exog. Baseline Exog.
+w/o Exog. w/ Exog. w/o Exog. w/ Exog.
 NHITS 8.268 7.304 11.114 9.094
@@
 nonlinear (0.080) (0.089)
-C.5 Model MAE and RMSE Results
+[…]
+We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values.
+However, this approach led to worse performance on the OhioT1DM dataset.
+The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters (Xie, 2018; Visentin et al., 2016).
+In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3.
+Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data.
+These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
+C.3 Model MAE and RMSE Results
 This appendix presents the complete results for all models for both simulated data and OhioT1DM.
 Standard deviation across 8 trials are provided in parenthesis.
-Table 6 provides complete results for the simulated data and Table 7 provides complete results for the OhioT1DM dataset.
-Table 6: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated data.
+Tables 8 and 9 provide complete MAE and RMSE results, respectively.
+Table 8: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated dataset and OhioT1DM dataset.
 Average results are computed across 8 trials.
-Model All Values Critical Values
-Baseline Exog. Baseline Exog.
-MAE
-ETS 9.450 13.457
-(0.00) (0.00)
-MLP 10.589 8.433 14.699 10.600
+Models are separated with horizontal lines into statistical, linear, RNN-based, CNN-based, Transformer-based, and MLP-based, and PK model groups, respectively.
+Result for models with (w/) and without (w/o) exogenous (exog.) variables are shown where applicable.
+Best results are highlighted in bold, second best results are underlined.
+PK models with improved forecasts over their respective exogenous baseline mdoel are highlighted in blue.
+Simulated Dataset OhioT1DM Dataset
+Model All Values Critical Values All Values Critical Values
+w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
+AutoETS 9.450 13.457 9.022 10.028
 (0.00) (0.00) (0.00) (0.00)
-RNN 13.703 20.190 21.433 35.955
+DLinear 7.802 7.989 10.575 10.790 11.514 13.011 12.512 14.580
+(0.054) (0.141) (0.126) (0.169) (1.595) (3.590) (1.228) (4.080)
+RNN 13.703 20.190 21.433 35.955 11.325 11.795 13.627 13.552
+(2.645) (9.966) (6.123) (23.235) (1.286) (1.985) (1.573) (2.480)
+LSTM 13.557 10.321 20.537 13.550 9.811 9.467 11.252 11.178
+(9.112) (2.074) (18.191) (2.817) (0.570) (0.230) (0.512) (0.396)
+TCN 12.096 10.109 17.709 12.660 11.525 10.760 13.647 12.678
+(1.681) (0.538) (2.995) (0.826) (2.008) (0.925) (3.349) (1.459)
+Informer 8.599 12.074 13.714 14.867
 (1.060) (2.083) (3.753) (3.620)
-LSTM 13.557 10.321 20.537 13.550
+PatchTST 7.284 9.842 9.922 10.999
 (0.554) (0.939) (0.949) (1.054)
-TCN 12.096 10.109 17.709 12.660
+TFT 6.587 5.826 8.518 7.132 9.402 9.197 10.577 10.643
+(0.115) (0.192) (0.178) (0.243) (0.811) (0.455) (1.038) (0.792)
+MLP 10.589 8.433 14.699 10.600 11.065 11.120 12.630 12.698
+(2.085) (0.490) (3.180) (0.464) (1.461) (0.381) (1.727) (0.821)
+NBEATSx 9.187 6.886 12.757 8.145 9.868 10.280 11.037 11.359
+(0.236) (0.031) (0.547) (0.072) (0.528) (0.214) (0.548) (0.188)
+NHITS 8.268 7.304 11.114 9.094 9.060 8.873 10.337 10.137
+(0.039) (0.055) (0.237) (0.236) (0.142) (0.064) (0.137) (0.108)
+TFT-PK 5.743 7.042 9.529 10.461
 (0.159) (0.270) (0.597) (0.597)
-NHITS 8.268 7.304 11.114 9.094
+NBEATSx-PK 6.806 7.998 9.983 10.927
 (0.060) (0.076) (0.264) (0.271)
-NHITS-ST 7.055 8.648
-(0.038) (0.068)
-NHITS-PK 7.037 8.492
-(0.093) (0.220)
-RMSE
-ETS 14.751 21.454
-(0.00) (0.00)
-MLP 15.561 11.461 22.380 14.449
+NHITS-PK 7.037 8.492 8.758 9.965
 (0.093) (0.220) (0.080) (0.095)
-RNN 19.033 27.116 28.834 43.256
+Table 9: Root Mean Square Error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated dataset and OhioT1DM dataset.
+Average results are computed across 8 trials.
+Models are separated with horizontal lines into statistical, linear, RNN-based, CNN-based, Transformer-based, and MLP-based groups, and PK model groups, respectively.
+Result for models with (w/) and without (w/o) exogenous (exog.) variables are shown where applicable.
+Best results are highlighted in bold, second best results are underlined.
+PK models with improved forecasts over their respective exogenous baseline mdoel are highlighted in blue.
+Simulated Dataset OhioT1DM Dataset
+Model All Values Critical Values All Values Critical Values
+w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
+AutoETS 14.751 21.454 14.569 16.225
 (0.00) (0.00) (0.00) (0.00)
-LSTM 18.567 13.723 26.863 18.028
+DLinear 12.545 12.673 17.606 17.707 17.138 18.609 18.679 20.743
+(0.085) (0.119) (0.340) (0.216) (1.462) (3.511) (0.970) (3.869)
+RNN 19.033 27.116 28.834 43.256 16.498 17.066 19.446 19.423
+(3.479) (12.484) (6.974) (23.190) (1.252) (2.037) (1.709) (2.532)
+LSTM 18.567 13.723 26.863 18.028 14.915 14.597 16.894 16.851
+(11.191) (2.373) (18.642) (3.661) (0.494) (0.279) (0.490) (0.394)
+TCN 17.094 13.551 25.163 16.908 16.690 15.971 19.213 18.418
+(1.719) (0.628) (3.092) (1.074) (1.891) (0.710) (2.706) (1.295)
+Informer 12.348 17.400 18.536 19.904
 (0.918) (1.865) (3.459) (3.220)
-TCN 17.094 13.551 25.163 16.908
+PatchTST 11.467 15.755 15.375 17.065
 (0.603) (1.000) (0.838) (0.928)
-NHITS 12.918 10.546 18.436 13.371
+TFT 10.691 8.898 14.188 10.979 14.832 14.571 16.523 16.577
+(0.139) (0.314) (0.329) (0.446) (0.791) (0.345) (0.959) (0.684)
+MLP 15.561 11.461 22.380 14.449 16.534 16.948 18.651 19.289
+(2.318) (0.553) (3.568) (0.502) (1.280) (0.481) (1.403) (0.943)
+NBEATSx 14.105 9.570 20.526 11.397 15.934 16.224 17.540 17.742
+(0.348) (0.053) (0.756) (0.104) (0.548) (0.675) (0.599) (0.303)
+NHITS 12.918 10.546 18.436 13.371 14.844 14.505 16.789 16.453
+(0.064) (0.106) (0.369) (0.387) (0.192) (0.096) (0.189) (0.132)
+TFT-PK 8.620 10.517 14.725 16.265
 (0.188) (0.367) (0.414) (0.437)
-NHITS-ST 10.082 12.679
-(0.054) (0.135)
-NHITS-PK 10.039 12.339
-(0.188) (0.435)
-Table 7: Mean absolute error (MAE) (top) and Root Mean Square Error (RMSE) (bottom) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for OhioT1DM data.
-Average results are computed across 8 trials.
-Model All Values Critical Values
-Baseline Exog. Baseline Exog.
-MAE
-ETS 9.022 10.028
-(0.00) (0.00)
-MLP 11.065 11.120 12.630 12.698
+NBEATSx-PK 9.480 11.244 16.945 17.659
 (0.084) (0.117) (1.467) (0.658)
-RNN 11.325 11.795 13.627 13.552
+NHITS-PK 10.039 12.339 14.385 16.251
 (0.188) (0.435) (0.092) (0.089)
-LSTM 9.811 9.467 11.252 11.178
-(0.570) (0.230) (0.512) (0.396)
-TCN 11.525 10.760 13.647 12.678
-(2.008) (0.925) (3.349) (1.459)
-NHITS 9.060 8.873 10.337 10.137
-(0.142) (0.064) (0.137) (0.108)
-NHITS-ST 8.958 10.131
-(0.087 ) (0.100)
-NHITS-PK 8.758 9.965
-(0.080) (0.095)
-RMSE
-ETS 14.569 16.225
-(0.00) (0.00)
-MLP 16.534 16.948 18.651 19.289
-(1.280) (0.481) (1.403) (0.943)
-RNN 16.498 17.066 19.446 19.423
-(1.252) (2.037) (1.709) (2.532)
-LSTM 14.915 14.597 16.894 16.851
-(0.494) (0.279) (0.490) (0.394)
-TCN 16.690 15.971 19.213 18.418
-(1.891) (0.710) (2.706) (1.295)
-NHITS 14.844 14.505 16.789 16.453
-(0.192) (0.096) (0.189) (0.132)
-NHITS-ST 14.640 16.492
-(0.205) (0.203)
-NHITS-PK 14.385 16.251
-(0.092) (0.089)

### v9

@@
 Furthermore, our proposed hybrid global-local architecture outperforms patient-specific PK models by 15.8%, on average.
-Our code is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code
+Data and Code Availability
+We use open-source simulated and real-world data that is available to other researchers.
+The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository (Xie, 2018; Visentin et al., 2016).
+The real-world data includes the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020).
+Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu (2020).
+Implementations of models used in our work were obtained from the open-source Neuralforecast (Olivares et al., 2022b) and StatsForecast (Garza et al., 2022) libraries.
+We implemented our method by adopting the Neuralforecast repository to leverage their model training framework.
+Our research code, along with the Neuralforecast models used in our research, is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
+Institutional Review Board (IRB)
+Our research does not require IRB approval.
 1 Introduction
+Figure 1: Insulin dose is recorded, but plasma insulin concentration is not directly measurable and typically requires invasive procedures and laboratory tests.
 Leveraging the predictive power of exogenous variables presents unique challenges, especially in healthcare, where the frequency of patient signals and exogenous factors, such as medication doses or treatments, often exhibit a mismatch in temporal resolution.
@@
 1.
-Consequently, the recorded drug dose alone cannot capture time-dependent drug kinetics, including processes such as absorption and elimination.
-Figure 1: Insulin dose is recorded, but plasma insulin concentration is not directly measurable and typically requires invasive procedures and laboratory tests.
-Consider the task of blood glucose monitoring for patients with diabetes who must regularly monitor and regulate their blood glucose level.
-Careful and prompt administration of insulin and intake of carbohydrates is required to prevent hyper- and hypoglycemic events, which can cause irreversible organ damage or, in severe cases, be fatal.
-Hyperglycemic and hypoglycemic episodes are responsible for numerous hospital admissions each year.
-A 2009 report on hospital admissions indicates over 500,000 patient-days and costs of \$2.4 billion each year for hyperglycemia alone (Kitabchi et al., 2009).
-Another report on hypoglycemic hospitalizations suggests a stable rate of 5.6-7.3 admissions per 1,000 patients annually between 2016-2019 (Pilla et al., 2021).
-With over 38 million patients with Diabetes Mellitus in the United States alone, this equates to a conservative estimate of 212,800 patients hospitalized for hypoglycemic episodes each year.
+Consequently, the recorded drug dose alone cannot capture time-dependent pharmacokinetics, including processes such as absorption and elimination.
+Consider the task of blood glucose monitoring for patients with Type-1 Diabetes Mellitus (T1DM) or insulin-dependent T2DM patients who must regularly monitor and regulate their blood glucose levels.
+Careful and prompt administration of insulin and intake of carbohydrates is required to prevent hyper- and hypoglycemic events, which account for numerous hospital admissions annually, and can cause irreversible organ damage or, in severe cases, be fatal.
+A 2016 report estimated that hypoglycemia causes approximately 100,000 emergency department (ED) visits per year due to insulin-related errors, with nearly one-third resulting in hospitalization (Geller et al., 2014).
+Another study reports that in 2014, there were 184,255 ED visits and in-patient U.S.
+hospital admissions for diabetic ketoacidosis (DKA), often associated with hyperglycemia, with 70.6% occurring in T1DM patients (Benoit et al., 2020).
+Another study reported a trend toward increasing hospitalizations for DKA in T1DM patients from 2008 to 2018 (Kichloo et al., 2021).
 For healthcare challenges like this, advanced forecasting technologies can help patients and clinicians monitor, anticipate, and proactively address emerging abnormalities to prevent adverse events.
 However, accurate forecasting requires models that can leverage exogenous information, such as treatments, while capturing latent treatment kinetics and their effects to generate reliable forecast trajectories.
-To address this, we propose a hybrid model that integrates cohort-level information with patient-specific pharmacokinetic characteristics to improve the predictability of blood glucose responses to treatment.
+To address this, we propose a hybrid model that integrates cohort-level information with patient-specific pharmacokinetics to improve the prediction of blood glucose responses to treatment.
 The main contributions of this paper are:
   1. (i)
-    Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable patient-specific encoder.
+    Hybrid Global-Local Architecture. We bolster global architectures (with parameters shared across patients) with a learnable, patient-specific encoder.
   2. (ii)
@@
 1.1 Pharmacokinetics Background
-PK modeling is a natural solution to the aforementioned problem, as it can reveal the effects of the medication over time through knowledge of drug movement within the body.
-For example, PK provides a well-established understanding of the changes in plasma insulin concentration over time based on drug absorption and elimination characteristics  (Binder et al., 1984).
+Pharmacokinetic (PK) modeling is a natural solution to the aforementioned problem, as it can reveal medication effects over time through well-established knowledge of changes in plasma insulin concentration driven by absorption and elimination characteristics  (Binder et al., 1984).
 Empirical and invasive studies show that concentration-time profiles for subcutaneous injections follow right-skewed curves.
@@
 A complete reference for supporting PK information regarding concentration-time profiles, insulin analogs, bioavailability, first-order kinetics, and linear PK can be found in Appendix B.
-Figure 2: Pharmacokinetic modeling of plasma insulin concentration over time for various insulin types.
+Figure 2: Pharmacokinetic models of plasma insulin concentration over time for various insulin types.
 2 Related Work
@@
 Modeling Blood Glucose Treatment Effects.
-Prior works have looked at modeling treatment effects on blood glucose levels.
+Prior work has looked at modeling treatment effects on blood glucose levels.
 Miller et al.
@@
 Odnoblyudova et al.
-improved the prediction of blood glucose using multi-output Gaussian processes (MOGP) to model the composite effects of meal components, such as carbohydrates and fats (Odnoblyudova et al., 2023).
-Unlike meals, insulin is a homogeneous protein hormone, making component-based modeling inapplicable to our task (Odnoblyudova et al., 2023).
+(2023) improved the prediction of blood glucose using multi-output Gaussian processes (MOGP) to model the composite effects of meal components, such as carbohydrates and fats.
+Unlike meals, insulin is a homogeneous protein hormone, making component-based modeling inapplicable to our task.
 Instead, our objective is to capture the holistic patient-specific effects of the drug and the composite effects of multiple doses.
@@
 The next paragraph outlines these limitations and how our proposed contributions address them.
-Previous work has developed individualized networks that overlook the benefits of developing global models that learn from various, diverse data series.
-Such cross-learning on diverse datasets can enhance generalization abilities, as demonstrated with Time Series Foundation Models (TSFMs) (Goswami et al., 2024).
+Previous work has focused on developing individualized networks, often overlooking the advantages of global models that learn from multiple, diverse time series—a technique known as cross-learning (Smyl, 2019; Semenoglou et al., 2021), and employed by Time Series Foundation Models (TFSMs) (Goswami et al., 2024).
 Our hybrid global-local architecture enables cohort-level training to develop a global model applicable to multiple patients while also preserving patient-specific information for personalized forecasts.
@@
 3 Methods
-We aim to forecast blood glucose levels 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10 hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values.
+We aim to forecast blood glucose values 30 minutes into the future consistent with prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020; Li et al., 2019b), based on a 10-hour history of blood glucose, carbohydrate (CHO), insulin basal, and insulin bolus values.
 We choose a 10-hour duration to balance sufficient historical information with our GPU memory constraints.
@@
 Thus, a 10-hour time span fully captures multiple dose durations, enabling the model to learn treatment effects across various dosages and durations.
+We also demonstrate that the significant performance improvement of the proposed methods holds across other various lag times in Appendix C.3.
 3.1 PK Encoder
-Medication administrations are represented as instantaneous sparse variables in time series data.
-We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentration to enable more accurate forecasting of treatment outcomes as shown in Fig.
+Medication administrations are recorded as instantaneous dosing events, leading to sparse observations in time series data.
+We propose a novel PK encoder that informs deep learning models of time-dependent plasma drug concentrations to enable more accurate forecasting of treatment outcomes as shown in Fig.
 3.
-The encoder consists of a function, $C$, that takes in sparse exogenous variables and generates a concentration-time profile for a specified dose, $\textbf{x}^{(i)}$, using the following equation:
-$C(t,\textbf{x}^{(i)},\textbf{k}^{(i)})=\frac{\textbf{x}^{(i)}}{\textbf{k}^{(i)}\sqrt{2\pi}t}\exp\{-\frac{1}{2{(\textbf{k}^{(i)})}^{2}}(\log(t)-1)^{2}\}$ (1)
-where $t$ is the time and $\textbf{k}^{(i)}$ is a PK parameter affecting elimination rate specific to the patient, $i$.
-Here, $C$ is equivalent to the log-normal probability density function (pdf) with $\sigma=\textbf{k}$ and $\mu=1$ and a scaling factor, x.
+The encoder consists of a function $C$ that generates a concentration-time profile for a single specified dose $x$ at time $t$ using a PK parameter $k$ affecting elimination rate:
+$C(t,x,k)=\frac{x}{tk\sqrt{2\pi}}\exp\{-\frac{1}{2k^{2}}(\log(t)-1)^{2}\}.$ (1)
+Here, $C$ is equivalent to the log-normal probability density function with $\sigma=k$ and $\mu=1$ and a scaling factor $x$.
 The intuition for the log-normal shape of the concentration curves stems from PK literature, as outlined in Section 1.1 and Appendix B.
@@
 Given this assumption, we show mathematically that the area under the concentration curve is bounded by the dose in Proposition A.1 in Appendix A.5.
-To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose x, as a scaling factor in $C$.
-This ensures that the area under the curve integrates to the dose value, under the simplifying assumption that all insulin administrations adhere to 100% bioavailability, which we show is upper bounded by the dose.
-Scaling the concentration curve changes its magnitude, but does not alter the PK parameters or shape of the curve.
-Let $N$ be the number of subjects in the data.
-The PK parameter $\textbf{k}\in[0,1]^{N\times 1}$ that controls the shape of the concentration curve, is represented as network embedding weight that is initialized between values of 1 and 2 prior to model training.
-The initial k values are tuned as model hyperparameters.
-The encoder is trained end-to-end with the model.
-During training, sparse insulin features are passed into the PK encoder along with their respective $\textbf{k}^{(i)}$ parameters, as shown in Fig.
+To model bioavailability and incorporate dose information directly into the concentration curves, we use the dose $x$ as a scaling factor in $C$.
+This ensures that, under the simplifying assumption of 100% bioavailability, the area under the curve integrates to the dose.
+Scaling the concentration curve in this way affects its magnitude but does not alter PK parameters that determine its shape.
+Let $N$ be the number of subjects in the dataset.
+The PK parameter vector $\mathbf{k}\in\mathbb{R}^{N\times 1}$ represents patient-specific embedding weights that control the shape of the patient-specific concentration curves, where $\mathbf{k}^{(i)}$ corresponds to the PK parameter $k$ for the $i$-th patient.
+The vector $\mathbf{k}$ is initialized with values between 1 and 2 prior to model training, with the initial value treated as a tunable hyperparameter.
+The encoder is trained end to end with the model.
+During training, a sparse treatment feature $\mathbf{x}^{(i)}\in\mathbb{R}^{1\times L}$, representing dose administrations within an input window of length $L$ up to and including the current time $T$, is passed to the PK encoder along with the corresponding parameter $\mathbf{k}^{(i)}$ to generate a concentration curve feature $\widecheck{\mathbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$, as shown in Fig.
 4.
-The encoder then replaces sparse medication features in the training window with the corresponding concentration curves.
-For a single dose, x, of insulin administered at time, $t$, we obtain the concentration, $\widecheck{\textbf{x}}^{(i)}$, at that time for patient, $i$:
-$\widecheck{\textbf{x}}^{(i)}$ $=C(t,\textbf{x}^{(i)},\textbf{k}^{(i)}).$ (2)
-For each training step, $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
-For each time step within the forecast horizon, $H$, forecasts, $\hat{\textbf{y}}$, are generated by the model, $f_{\theta}$, as a function of static features, s, treatment concentration curves, $\widecheck{\textbf{x}}$, and historical glucose values y recorded up to and including time $t$ based on the lag time, $L$,
-$\hat{\textbf{y}}^{(i)}_{t+1:t+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{t-L:t},\widecheck{\textbf{x}}^{(i)}_{t-L:t},\textbf{s}^{(i)}).$ (3)
+In the single-dose setting, where a dose is recorded at time step $t_{d}\in\{T-L+1,\dots,T\}$, the encoder replaces the sparse treatment feature $\mathbf{x}_{T-L+1:T}^{(i)}$ with the corresponding concentration feature $\widecheck{\mathbf{x}}_{T-L+1:T}^{(i)}$, computed as:
+$\widecheck{\mathbf{x}}^{(i)}_{t}$ $=\begin{cases}0,&\!\text{for }t\in[T-L+1,t_{d})\\ C\left(\nu(t-t_{d}),\mathbf{x}^{(i)}_{t_{d}},\mathbf{k}^{(i)}\right),&\!\text{for }t\in[t_{d},T]\end{cases}$ (2)
+where $t$ indexes time steps within the window, $\nu$ is the sampling interval of the data, and $C(\cdot)$ computes the concentration at elapsed time $\nu(t-t_{d})$ since the dose was administered.
+During training, the parameter $\mathbf{k}^{(i)}$ is updated for each patient $i$ via gradient descent.
+Forecasts $\hat{\textbf{y}}$ are generated by the model $f_{\theta}$ for each time step within the forecast horizon $H$ as a function of historical glucose values y, treatment concentration curves $\widecheck{\textbf{x}}$, and static features s, recorded within the input window up to and including time $T$,
+$\hat{\textbf{y}}^{(i)}_{T+1:T+H}$ $=f_{\theta}(\textbf{y}^{(i)}_{T-L+1:T},\widecheck{\textbf{x}}^{(i)}_{T-L+1:T},\textbf{s}^{(i)}).$ (3)
 In the proposed context, s includes patient age, weight, one-hot encoded subject identification number, and insulin pump type.
-$\widecheck{\textbf{x}}_{t-L:t}$ includes concentration curves for basal insulin, $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin, $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO, $\widecheck{\textbf{x}}_{\text{CHO}}$.
+$\widecheck{\textbf{x}}_{T-L+1:T}$ includes concentration curve values for basal insulin $\widecheck{\textbf{x}}_{\text{basal}}$, bolus insulin $\widecheck{\textbf{x}}_{\text{bolus}}$, and CHO $\widecheck{\textbf{x}}_{\text{CHO}}$.
 We include treatment effects of carbohydrates because, similar to drugs, they undergo absorption and metabolism processes, resulting in glucose and other sugars entering the bloodstream.
@@
 3 and Fig.
-14.
-To model PK effects for multiple doses in an input window, we generate a concentration curve for each dose event and sum concentration curves across the windows.
-Under the assumption of linear pharmacokinetics (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.4.
-To perform vectorized operations that enable generating concentration curves for multiple doses in an input window, the PK encoder leverages a matrix, $\textbf{w}\in\mathbb{R}^{L\times L}$, where $L$ refers to the lag time.
-This approach allows for efficient modeling of multiple effects in $\mathcal{O}(1)$ time.
-Each row of w represents a time count up to the input time where the count shifts one time step for each subsequent row.
-Concentration curves are generated for each row of w using the dose at the corresponding time index.
-If no dose is administered at a particular time, $t$, the corresponding row of w indexed by $t$ will be zeroed out.
-The final concentration curve feature, which can reflect the impact of multiple doses, is generated by aggregating the individual concentration curves across time steps:
-w $=\begin{bmatrix}0&1&2&3&\cdots&L-1\\ 0&0&1&2&\cdots&L-2\\ \ldots&\ldots&\ldots&\ldots&\ldots&\ldots\\ 0&0&0&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{L\times L}$ (4)
-$\widecheck{\textbf{x}}^{(i)}$ $=\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}^{(i)}_{t},\textbf{k}^{(i)})$ (5)
-Here, $f$ is the frequency of the data, $\textbf{x}^{(i)}\in\mathbb{R}^{L\times 1}$ a sparse exogenous time series of treatment, and $\widecheck{\textbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$ is the new concentration curve feature for patient, $i$.
-The PK encoder training algorithm is described in Appendix A.5, which can be integrated into any deep learning architecture capable of handling exogenous variables in addition to modeling the target variable.
+15.
+To model PK effects for multiple doses within an input window, we generate a concentration curve for each dose event and sum the concentration curves within the window.
+Under the assumption of linear PK (outlined in 1.1 and Appendix B), we show that the overall effect, or concentration, of $n$ doses of the same drug is equal to the sum of the individual effects which is verified mathematically in Proposition A.2 in Appendix A.5.
+To efficiently generate concentration curves for multiple doses within an input window, the PK encoder uses a matrix $\textbf{w}\in\mathbb{R}^{L\times L}$ to support vectorized computation at each timestamp containing a dose, within the lag time $L$.
+This approach enables efficient modeling of multiple effects in $\mathcal{O}(1)$ time.
+Each row $t$ of w represents the elapsed time steps until the end of the window, with the time count shifting one step to the right in each subsequent row.
+A concentration curve is generated for each row of w using the dose at the corresponding time index.
+If no dose is administered at a particular time $t$, the corresponding row of w is zeroed out.
+The final concentration curve feature, which captures the combined impact of multiple doses, is obtained by aggregating the individual curves across the rows:
+$\mathbf{w}$ $=\begin{bmatrix}0&1&2&3&\cdots&L{-}1\\ 0&0&1&2&\cdots&L{-}2\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&0\end{bmatrix}\in\mathbb{R}^{L\times L}$ (4)
+$\widecheck{\textbf{x}}^{(i)}$ $=\sum_{t=0}^{L-1}C(\nu\textbf{w}_{t,:},\textbf{x}^{(i)}_{t},\textbf{k}^{(i)})$ (5)
+Here, $\nu$ is the sampling interval of the data, $\textbf{x}^{(i)}\in\mathbb{R}^{1\times L}$ is a sparse exogenous time series treatment feature, and $\widecheck{\textbf{x}}^{(i)}\in\mathbb{R}^{1\times L}$ is the new concentration curve feature for patient $i$.
+The PK encoder training algorithm 1 is described in Appendix A.6, which can be integrated into any deep learning architecture capable of handling exogenous variables in addition to modeling the target variable.
 Figure 4: Our hybrid global-local architecture combines global model parameters shared across patients with patient-specific pharmacokinetic (PK) parameters.
-Sparse treatment time series features, such as medication doses, are input into the PK encoder to generate patient-specific treatment effect curves, $\check{\textbf{x}}$.
-PK encoder output, $\check{\textbf{x}}$, and any other exogenous features, x, are used as deep learning model inputs.
-During each training step, the model learns relevant hidden states and output parameters, $\hat{\theta}$, across patients, while $\textbf{k}^{(i)}$ values are updated for each patient, $i$, using gradient descent optimization.
+Sparse treatment time series features, such as medication doses, are input into the PK encoder to generate patient-specific treatment concentration curves $\check{\textbf{x}}$.
+PK encoder output $\check{\textbf{x}}$ and any other exogenous features x are used as deep learning model inputs.
+During each training step, the model learns relevant hidden states and output parameters $\hat{\theta}$ across patients, while $\textbf{k}^{(i)}$ values are updated for each patient $i$ using gradient descent optimization.
 3.3 Modeling Dose-Dependent Effects
-A majority of drugs are assumed to adhere to linear PK, where it is assumed that there is a linear relationship between drug dose and concentration in the body (Rosenbaum, 2017).
-A characteristic of this assumption is that PK parameters are constant and do not change with dose.
-In practice, a drug’s PK parameters can vary depending on factors such as the drug’s properties, the type of injection, and the injection site.
-Nonlinear PK occurs when the relationship between drug dose and body concentration is not linear, leading to changes in PK parameters with dose.
+A majority of drugs are assumed to adhere to linear pharmacokinetics (PK), where a linear relationship is expected between drug dose and its concentration in the body (Rosenbaum, 2017).
+In practice, however, whether a drug’s PK parameters follow linear or nonlinear behavior can vary based on factors such as the drug’s physicochemical properties, the route of administration, and the injection site.
+A key characteristic of the linear PK assumption is that PK parameters remain constant and do not change with variables such as dose or time.
+Dose- or time-dependent nonlinearity occurs when PK parameters change nonproportionally with dose or time, respectively.
+In dose-dependent nonlinear PK, the relationship between drug dose and body concentration is no longer linear, resulting in dose-related changes in PK parameters.
+While insulin is not expected to exhibit classic time-dependent nonlinearity, it may demonstrate dose-dependent nonlinearity.
 For instance, subcutaneous injections are administered into the subcutaneous tissue, where higher doses can saturate elimination pathways, causing slower absorption rates (Søeborg et al., 2009; Gradel et al., 2018).
-This variability in PK parameters can result in dose-dependent plasma drug concentration profiles.
+When PK parameters are dose-dependent, the resulting plasma drug concentration profiles also vary with dose.
 Insulin pumps administer the same type of insulin (rapid-acting) for the basal and bolus doses.
 However, bolus insulin doses are generally much larger than basal doses as highlighted in Figs.
-12 and 13 in Appendix A.1.
-To account for potential dose-dependent effects, we treat each dose range ($\textbf{k}_{\text{bolus}}$ and $\textbf{k}_{\text{basal}}$) as a separate entity, so that we can capture potential PK variations without resorting to a fully nonlinear PK model.
-This method simplifies the modeling process by maintaining linear assumptions within each dose category.
-It avoids the complexity while still accounting for potential differences between the two categories of low-dose (basal) and high-dose (bolus) insulin kinetics.
-Specifically, we separately estimate $\textbf{k}_{\text{basal}}$ and $\textbf{k}_{\text{bolus}}$ values.
-For a single dose of basal insulin, $\textbf{x}_{\text{basal}}$, or bolus insulin, $\textbf{x}_{\text{bolus}}$, administered at time, $t$, we obtain the plasma insulin concentration at that time for patient, $i$:
-$\widecheck{\textbf{x}}^{(i)}_{\text{basal}}$ $=C(t,\textbf{x}^{(i)}_{\text{basal}},\textbf{k}_{\text{basal}}^{(i)}),$ (6)
-$\widecheck{\textbf{x}}^{(i)}_{\text{bolus}}$ $=C(t,\textbf{x}^{(i)}_{\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)}).$ (7)
-Treating each dose category as a distinct entity is an approximation and may not capture all nuances of dose-dependent kinetics.
-However, all physiological models approximate complex biological processes.
-Our goal is to develop models validated under established linear PK assumptions from the pharmacology literature while still accounting for dose-dependent effects, balancing simplicity and precision.
+13 and 14 in Appendix A.1.
+To account for potential dose-dependent effects, we model basal and bolus doses separately by learning distinct PK parameters, $\mathbf{k}_{\text{basal}}$ and $\mathbf{k}_{\text{bolus}}$, allowing us to capture PK variations without resorting to a fully nonlinear model.
+This approach maintains linear assumptions within each dose type, simplifying the modeling process while still capturing differences between low-dose (basal) and high-dose (bolus) PK.
+For a single dose of basal insulin $\mathbf{x}_{\text{basal}}$ or bolus insulin $\mathbf{x}_{\text{bolus}}$, administered at time step $t_{d}$, we compute the plasma insulin concentration at time $t\geq t_{d}$ for patient $i$:
+$\widecheck{\textbf{x}}^{(i)}_{\text{basal}}$ $=C\left(\nu(t-t_{d}),\textbf{x}^{(i)}_{t_{d},\text{basal}},\textbf{k}_{\text{basal}}^{(i)}\right),$ (6)
+$\widecheck{\textbf{x}}^{(i)}_{\text{bolus}}$ $=C\left(\nu(t-t_{d}),\textbf{x}^{(i)}_{t_{d},\text{bolus}},\textbf{k}_{\text{bolus}}^{(i)}\right).$ (7)
+Treating each dose category as a distinct entity is an approximation and may not capture all nuances of dose-dependent nonlinear PK.
+All physiological models are approximations of complex biology.
+Our goal is to evaluate a novel deep learning-based PK modeling approach based on established linear PK knowledge, while accounting for potential dose-dependent effects, balancing simplicity and precision.
 This framework also ensures broader applicability across various medications, as most drugs are assumed to follow linear PK (Rosenbaum, 2017).
@@
 Global models are trained on all patient data to create a single model.
-Global model baselines include: (1) univariate models trained with only the blood glucose signal and static features, (2) models trained with univariate blood glucose signals, static features, and sparse historical exogenous features (CHO and insulin dose values), and (3) models trained with the univariate blood glucose signal, static features, and sparse historical exogenous features transformed using the “Sum Total” (ST) approach, which converts features into cumulative values across an input window (Rubin-Falcone et al., 2022).
-Univariate models are included as baselines to highlight forecast improvements from training on historical exogenous variables, motivating advanced transformations for sparse, continuous exogenous features using methods like our PK encoder.
+Global model baselines include: (1) models trained with only the blood glucose signal and static features, (2) models trained with blood glucose signals, static features, and sparse historical exogenous features (CHO and insulin dose values), and (3) models trained with blood glucose signals, static features, and sparse historical exogenous features transformed using the “Sum Total” (ST) approach, which converts features into a cumulative sum of values across an input window (Rubin-Falcone et al., 2022).
+We include models trained without exogenous variables as baselines to show the impact of incorporating sparse exogenous information for each model.
+Together with models trained on sparse and Sum Total features, these baselines demonstrate the importance of effective transformations—such as our PK encoder—in reducing forecast error.
 With the exception of TFT, most Transformer-based models, including Informer and PatchTST, are not designed to model covariates.
@@
 Local Model Baselines.
-Patient-specific sparse exogenous models as well as patient-specific exogenous models using the PK encoder were compared against our hybrid global-local PK encoder model.
+Patient-specific models using sparse exogenous inputs, as well as those incorporating the PK encoder, were compared against our hybrid global-local PK encoder model.
+These baselines highlight the effectiveness of hybrid global-local architectures that leverage cross-learning from multiple time series.
 3.5 Model Training
 Models were trained using Adam optimizer (Kingma and Ba, 2017).
-Model hyperparameters that minimized Huber loss across all time points in the prediction horizon of the validation set were selected.
+Optimal hyperparameters were selected via cross-validation on a validation set by minimizing Huber loss across all prediction time points.
+Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set.
 Huber estimator was selected as the loss function, $\mathcal{L}$, given its ability to reduce the impact of outliers and offer more stable results compared to other functions:
-$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)=\left\{\begin{array}[]{ll}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{array}\right.$ (8)
-Given the use of forward-filling to account for missing values in the OhioT1DM dataset, timestamps with forward-filled values were omitted from the validation set.
-Optimal hyperparameters were selected based on a validation set with cross-validation.
+$\mathcal{L}(\textbf{y},\hat{\textbf{y}},\delta)=\left\{\begin{array}[]{ll}\frac{1}{2}(\textbf{y}-\hat{\textbf{y}})^{2},&\text{if }|\textbf{y}-\hat{\textbf{y}}|\leq\delta,\\ \delta(|\textbf{y}-\hat{\textbf{y}}|-\frac{1}{2}\delta),&\text{otherwise}.\end{array}\right$.
+(8)
 More information on model training and hyperparameters in provided in Appendix A.3.
@@
 Trained models were used to generate rolling window forecasts at one-step intervals.
-Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE):
-$\mathrm{MAE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\frac{1}{H}\sum_{{\tau}=t+1}^{t+H}|\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau}|\quad\text{and}$ (9)
-$\mathrm{RMSE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\sqrt{\frac{1}{H}\sum_{\tau=t+1}^{t+H}(\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau})^{2}}.$ (10)
+Similar to prior work (Rubin-Falcone et al., 2022; Xie and Wang, 2020), we evaluate forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE).
 We compute results across all points within the forecast horizon, and present the average results over 8 trials of individually trained models.
@@
 For the OhioT1DM dataset, timestamps with forward-filled values were omitted from evaluation.
-To evaluate whether hybrid global-local PK models significantly outperform baselines, we use paired t-test statistics computed on the patients’ mean MAE results.
+To evaluate whether hybrid global-local PK models significantly outperform baselines, we use paired t-tests computed on the patients’ mean MAE results.
+Metric formulations and statistical tests are detailed in Appendix A.4.
 4 Cohort
 We use simulated and real-world datasets containing patient insulin, CHO time series, and blood glucose measurements.
-These open-source datasets ensure reproducibility and are widely used in blood glucose forecasting research, featuring some of the largest patient cohorts (Xie and Wang, 2020; Rubin-Falcone et al., 2020, 2022; Rabby et al., 2021; Li et al., 2019b).
+These open-source datasets ensure reproducibility and are among the largest publicly available cohorts for blood glucose forecasting, making them widely used in current research within this domain (Xie and Wang, 2020; Rubin-Falcone et al., 2020, 2022; Rabby et al., 2021; Li et al., 2019b).
 Simulated data.
@@
 Appendix A.1 contains examples of blood glucose time series for 6 individuals, shown in Fig.
-10, as well as basal and bolus dose distributions for 12 individuals, shown in Fig.
-12.
-Figure 5: The hybrid global-local NHITS-PK model trained on all OhioT1DM individuals (orange, solid median) has a lower average MAE than NHITS-PK local models (blue, dashed median).
+11, as well as basal and bolus dose distributions for 12 individuals, shown in Fig.
+13.
+Figure 5: The hybrid global-local NHITS-PK model trained on all OhioT1DM individuals (orange, solid median) has a lower MAE than NHITS-PK local models (blue, dashed median).
 Figure 6: The NHITS-PK hybrid global-local model (orange, solid median) consistently outperforms the local NHITS-PK model (blue, dashed median).
-[…]
+Table 1: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose $\leq 70~|\geq 180$).
+Models are separated with horizontal lines into statistical, linear, RNN-based, CNN-based, Transformer-based, and MLP-based, and PK model groups, respectively.
+Result for models with (w/) and without (w/o) exogenous (exog.) variables are shown where applicable.
+Best results are highlighted in bold, second best results are underlined.
+PK models with improved forecasts over their respective exogenous baseline model are highlighted in blue.
 Simulated Dataset OhioT1DM Dataset
 Model All Values Critical Values All Values Critical Values
-w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
 AutoETS 9.450 13.457 9.022 10.028
 DLinear 7.802 7.989 10.575 10.790 11.514 13.011 12.512 14.580
-RNN 13.703 20.190 21.433 35.955 11.325 11.795 13.627 13.552
-LSTM 13.557 10.321 20.537 13.550 9.811 9.467 11.252 11.178
-TCN 12.096 10.109 17.709 12.660 11.525 10.760 13.647 12.678
+RNN 12.728 12.930 20.308 20.945 11.535 10.907 12.641 12.989
+LSTM 9.250 9.152 12.034 12.000 10.217 10.615 12.102 12.983
+TCN 14.340 11.380 23.370 14.817 10.769 11.516 12.812 13.203
 Informer 8.599 12.074 13.714 14.867
@@
 TFT 6.587 5.826 8.518 7.132 9.402 9.197 10.577 10.643
-MLP 10.589 8.433 14.699 10.600 11.065 11.120 12.630 12.698
+MLP 10.424 8.407 13.932 10.297 10.300 13.254 11.550 14.997
 NBEATSx 9.187 6.886 12.757 8.145 9.868 10.280 11.037 11.359
@@
 Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a continuous glucose monitor (CGM) at a frequency of 5-minutes, basal rate information, bolus insulin (units), and CHO values.
-The maximum number of test samples consistent across all patients (2691 samples) was used to obtain equal test sets of approximately 9 days.
+The maximum number of test samples consistent across all patients (2691 timestamps) was used to obtain equal test sets of approximately 9 days.
 Missing values were forward filled to prevent data leakage.
-The sample count and percentages of missing data for each subject are included in Table 4 in Appendix A.1.
+Sample counts and missing data percentages for each subject are included in Table 4 in Appendix A.1.
 Appendix A.1 also contains examples of blood glucose time series plots for 6 subjects, shown in Fig.
-11, as well as basal and bolus dose distributions for the 12 patients, shown in Fig.
-13.
+12, as well as basal and bolus dose distributions for the 12 patients, shown in Fig.
+14.
 5 Results
@@
 For simulated data, the TFT-PK model achieves the lowest MAE, on average, among all models for both evaluations across all models and critical values.
-The PK models, including TFT-PK, NBEATSx-PK, and NHITS-PK outperform their sparse exogenous model counterparts by up to 5.4%, 3.7%, 9.2%, respectively, on evaluations across all values and 7.7%, 6.6%, 16.4% on evaluations across critical values for individual subjects.
+The PK models, including TFT-PK, NBEATSx-PK, and NHITS-PK outperform their sparse exogenous model counterparts by up to 5.4%, 3.7%, 9.2%, respectively, on evaluations across all values and 7.7%, 6.6%, 16.4% on evaluations across critical values for individual patients.
 For the OhioT1DM dataset, the NHITS-PK model achieves the lowest MAE, on average, among all models for both evaluations across all values and critical values.
 The PK models, including, TFT-PK, NBEATSx-PK, and NHITS-PK, outperform their sparse exogenous model counterparts by up to 5.2%, 8.5%, 4.9% on evaluations across critical values for individual patients.
-Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Appendix C.3 Tables 8 and 9, respectively.
+Mean MAE and RMSE results and their standard deviations computed across 8 trials are shown in Appendix C.7 Tables 11 and 12.
 Hybrid global-local models significantly improve forecasts over global models.
-For the simulated dataset, the TFT-PK, NBEATSx-PK, and NHITS-PK models achieve significantly lower forecast errors for subjects for evaluations on all values (p-values: 4.6e-4 [TFT-PK]; 7.9e-5 [NBEATSx-PK]; 2.0e-7 [NHITS-PK]) and critical values (p-values: 5.7e-4 [NBEATSx-PK]; 3.0e-5 [NHITS-PK]) compared to their sparse exogenous global model counterparts.
-For the OhioT1DM dataset, the NBEATSx-PK and NHITS-PK models also achieve significantly lower forecast errors for subjects for both evaluations on all values (p-values: 3.4e-3 [NBEATSx-PK]; 1.3e-3 [NHITS-PK]) and critical values (p-values: 4.4e-4 [NBEATSx-PK]; 7.0e-3 [NHITS-PK]) compared to their sparse exogenous global model counterparts.
+For the simulated dataset, the TFT-PK, NBEATSx-PK, and NHITS-PK models achieve significantly lower forecast errors for patients for evaluations on all values (p-values: 4.6e-4 [TFT-PK]; 7.9e-5 [NBEATSx-PK]; 2.0e-7 [NHITS-PK]) and critical values (p-values: 5.7e-4 [NBEATSx-PK]; 3.0e-5 [NHITS-PK]) compared to their sparse exogenous global model counterparts.
+For the OhioT1DM dataset, the NBEATSx-PK and NHITS-PK models also achieve significantly lower forecast errors for patients for both evaluations on all values (p-values: 3.4e-3 [NBEATSx-PK]; 1.3e-3 [NHITS-PK]) and critical values (p-values: 4.4e-4 [NBEATSx-PK]; 7.0e-3 [NHITS-PK]) compared to their sparse exogenous global model counterparts.
+The TFT-PK model does not outperform its sparse exogenous counterpart for L=120L\!=\!120.
+However, for L=180L\!=\!180, it achieves significantly lower error (p-value: 4.8e-2) than its sparse exogenous variant and records the lowest error among all TFT model variants, as shown in Table 8.
 Additionally, Table 2 shows that TFT-PK and NHITS-PK models achieve significantly lower errors, on average—2.4% (p-value: 3.6e-12) and 2.3% (p-value: 4.1e-5), respectively—for subjects in evaluations across all values, compared to their “Sum Total” model counterparts.
@@
 Our hybrid global-local NHITS-PK model also significantly outperforms local sparse exogenous baseline models by 14.8%, on average, for individual subjects (p-value: 1.7e-6).
-Table 2: PK models outperform Sum Total (ST) models for both Simulated and OhioT1DM datasets. Best result is in bold.
+Table 2: PK models outperform Sum Total (ST) models for both Simulated and OhioT1DM datasets. The best results are in bold.
 Simulated Dataset OhioT1DM Dataset
@@
 All Values 5.743 5.885 8.758 8.958
-(0.159) (0.187) (0.080) ((0.087)
+(0.159) (0.187) (0.080) (0.087)
 Critical Values 7.042 7.233 9.965 10.131
 (0.270) (0.282) (0.095) (0.100)
-Table 3: MAE computed across forecast horizon values for local NHITS, global NHITS, local NHITS-PK, and hybrid global-local NHITS-PK on the 12 subjects from the OhioT1DM dataset. Best results in bold.
+Table 3: MAE computed across forecast horizon values for local NHITS, global NHITS, local NHITS-PK, and hybrid global-local NHITS-PK on the 12 subjects from the OhioT1DM dataset.
+The best results are in bold.
 ID w/ Exog. (Sparse) w/ Exog. (PK)
@@
 Forecasts for the NHITS-PK model (orange, dashed) are more aligned with ground truth (black, solid) compared to the NHITS sparse exogenous model blue, dash-dotted).
-Figure 8: The NHITS-PK model captures insulin treatment effects: forecasts with removed bolus doses (yellow, dash-dotted) and augmented doses (red, dotted), result in higher and lower glucose predictions, respectively, compared to the original doses (orange, dashed).
-Each point represents the final forecasted timestamp from a rolling window forecast.
 6 Discussion
 The proposed PK encoder can substantially reduce forecasting errors in deep learning models for both simulated and real-world data.
-These improvements are not only statistically significant, but they have potential for positive real-world clinical impact.
 In addition to providing more accurate forecasts during hyperglycemic events, our PK encoder predicts these events earlier, up to 3.6 minutes compared to the NHITS sparse exogenous baseline model, for individual OhioT1DM patients, as shown in Fig.
-9.
-This work uses data from T1DM patients but incorporates key variables relevant to both T1DM and T2DM, such as continuous glucose measurements (CGM), carbohydrate intake, and insulin doses, making our method applicable to both populations.
-Improving the prediction of critical events for 9 out of 12 (75%) of the 38 million Americans with Diabetes Mellitus could benefit 28.5 million individuals (Kitabchi et al., 2009).
-Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing.
-While we demonstrate performance improvements across the top three models, our approach can be easily integrated into other deep learning architectures as new tools emerge.
-It can also be easily adapted for other types of medication administration, such as oral or intravenous injections.
-In addition to improving performance over global model counterparts, our hybrid global-local model architecture produces forecasts with significantly smaller errors than patient-specific models.
-This result is of substantial importance in demonstrating that global models can integrate cohort-level information with patient-specific parameters to produce more accurate personalized forecasts.
+10.
+The TFT-PK model achieves a true positive rate of 0.96 on the simulated dataset, while the NHITS-PK model reaches 0.89 on the OhioT1DM dataset, showcasing their effectiveness in predicting critical events while leaving room for further improvement.
+A complete analysis of hyper- and hypoglycemic event prediction is provided in Appendix C.4.
+Figure 8: The NHITS-PK model captures insulin treatment effects: forecasts with removed bolus doses (yellow, dash-dotted) and augmented doses (red, dotted), result in higher and lower glucose predictions, respectively, compared to the original doses (orange, dashed).
+Each point represents the final forecasted time step from a rolling window forecast.
+In addition to improving performance over global model baselines, our hybrid global-local model architecture produces forecasts with significantly smaller errors than patient-specific models.
+This result is important in demonstrating that global models can integrate cohort-level information with patient-specific parameters to produce more accurate personalized forecasts.
+Evaluations on the OhioT1DM dataset also demonstrate that our hybrid global-local model remains effective even with 20% of missing timestamps in the training set for individual patients, as shown in Table 3, with missing data detailed in Table 4.
 Moreover, our proposed hybrid global-local model offers practical benefits beyond improved forecasting performance as it requires less computational resources and time to train than individual patient models.
 Training patient-specific NHITS-PK models took approximately seven times longer than training a single NHITS-PK hybrid model, despite using a common hyperparameter space.
-The hybrid model also has the potential to be fine-tuned for new patients with limited or no historical data, where training effective patient-specific models is infeasible.
-Our study demonstrates that leveraging domain knowledge to integrate exogenous variables in models offers performance benefits over using sparse features or the “Sum Total” feature engineering approach, both of which are outperformed by our proposed deep PK encoding method.
-Furthermore, dose-dependent PK encoding strategies, such as those proposed in Section 3.3, can lead to more realistic models, as highlighted by our ablation study in Appendix C.2 where we demonstrate that using our PK encoder to infer the same k value for both bolus and basal insulin instead of separate values results in worse performance for the OhioT1DM dataset.
+Additionally, the PK models are as computationally efficient as or more efficient than their sparse exogenous counterparts in terms of floating-point operations per second (FLOPs) and the number of trainable parameters, as shown in Table 9 in Appendix C.5.
+(a) Simulated Dataset
+(b) OhioT1DM Dataset
+(c) Simulated Dataset
+(d) OhioT1DM Dataset
+Figure 9: True positive rate (TPR) for the Simulated and OhioT1DM datasets at various glucose thresholds for (a, b) hyperglycemia (blood glucose $\geq 180$) and (c, d) hypoglyemia (blood glucose $\leq 70$) event warnings.
+TPR represents the proportion of forecast horizon windows in which the model correctly predicts an event when one actually occurs.
+The best performing TFT-PK model for the Simulated dataset predicts hyperglycemic and hypoglycemic events in forecast windows at an approximate TPR of 0.96 and 0.93, respectively.
+The best performing NHITS-PK model for the OhioT1DM dataset predicts hyperglycemic and hypoglycemic events in forecast windows at an approximate TPR of 0.89 and 0.69, respectively.
+Figure 10: The PK encoder enables improved NHITS model prediction lead times for hyperglycemic events (CGM $\geq$ 180) for 9 of 12 OhioT1DM patients.
+Our study demonstrates that leveraging PK domain knowledge to integrate exogenous variables in models offers performance improvements over using sparse features or the “Sum Total” feature engineering approach.
+Furthermore, dose-dependent PK encoding strategies, such as those proposed in Section 3.3, can lead to more realistic models, as highlighted by our ablation study in Appendix C.2 where we demonstrate that using our PK encoder to infer the same k value for both bolus and basal insulin, rather than separate values, results in worse performance on the OhioT1DM dataset.
+Our PK encoder can have multiple beneficial applications in clinical practice, such as issuing early warnings of unexpected treatment responses and characterizing patient-specific effects for personalized treatment dosing.
+We discuss the feasibility of integrating such deep learning models into clinical decision-support systems in section C.6.
+While we demonstrate performance improvements across the top three models, our approach can be easily integrated into other deep learning architectures as new tools emerge.
+It can also be adapted for other types of medication administration, such as oral or intravenous injections.
 Limitations and Future Work.
-This study has two limitations related to the PK encoder: we assume that the concentration curves exhibit 100% bioavailability and maintain a consistent peak concentration across patients.
-These simplifying assumptions are intended to focus model learning on the parameter $k$, rather than on the area or peak of the concentration curve, to emphasize the spread or duration of the concentration curve and to reduce potential issues with parameter identification.
-Future work will explore expanding concentration curve parameters to estimate area and peak parameters to potentially enable more flexible concentration curves.
-We will also aim to apply our approach to additional cohorts of patients, such as those with T2DM, as more data become available, to further validate our findings.
-Future work will also explore modeling dose-dependent effects using more dynamic nonlinear PK frameworks.
-Figure 9: The PK encoder enables improved NHITS model prediction lead times for hyperglycemic events (CGM $\geq$ 180) for 9 of 12 OhioT1DM patients.
+This study has two limitations related to the PK encoder: we assume 100% bioavailability and a consistent peak concentration across patients.
+These assumptions focus model learning on the parameter $k$, rather than the area or peak of the concentration curve, to help mitigate potential issues with parameter identification.
+Future work will explore expanding concentration curve parameters as well as using nonlinear PK frameworks to potentially enable more flexible concentration curves.
+The proposed method is demonstrated on T1DM patient data but has potential to be adapted to other forecasting tasks and populations, such as insulin-dependent T2DM patients, with appropriate cohort data, PK model adjustments, and model training.
+Future work will apply our approach to additional cohorts, including T2DM, as more data becomes available to further validate our findings.
+Future work will also explore methods for estimating patient-specific parameters in a global model for new patients with limited or no historical data.
 […]
-This work was partially supported by the NIH award R01HL141916. The authors wish to thank Dr. Frank Guyette, Dr. Leonard Weiss and Dr. Gilles Clermont for their guidance and informative conversations.
-Data and Code Availability
-We use open-source simulated and real-world data that is available to other researchers.
-The simulated data includes an open-source Python implementation of the FDA-approved UVa/Padova Simulator (2008 version), which is contained in the ‘simglucose’ GitHub repository (Xie, 2018; Visentin et al., 2016).
-The real-world data includes the OhioT1DM 2018 and 2020 datasets (Marling and Bunescu, 2020).
-Information on accessing this dataset, including a link to the required Data Use Agreement (DUA), is provided in the paper by Marling and Bunescu (2020).
-Implementations of models used in our work were obtained from the open-source Neuralforecast (Olivares et al., 2022b) and StatsForecast (Garza et al., 2022) libraries.
-We implemented our method by adopting the Neuralforecast repository to leverage their model training framework.
-Our research code, along with the Neuralforecast models used in our research, is available at https://github.com/PotosnakW/neuralforecast/tree/pk_paper_code.
+This work has been partially supported by the NIH awards R01NR013912 and R01DK131586, and NSF awards 2406231 and 2427948.
+The authors wish to thank Dr.
+Frank Guyette, Dr.
+Leonard Weiss and Dr.
+Gilles Clermont for their guidance and informative conversations.
 Appendix A Supplemental Methods
@@
 Each patient has approximately 8 weeks of data that includes blood glucose (mg/dL) measurements recorded with a CGM at a frequency of 5-minutes.
-The dataset consists of both female (n=5) and male (n=8) subjects with unspecified ages within the ranges of 20-40, 40-60, and 60-80.
+The dataset consists of both female (n=5) and male (n=8) patients with unspecified ages within the ranges of 20-40, 40-60, and 60-80.
 Information on insulin pump model and type of insulin is also provided.
@@
 14 in Appendix A.1.
-Fig.
-10 shows the first 10 days of data for 6 subjects from the simulated dataset.
-Fig.
-11 shows the first 10 days of data for 6 subjects from the OhioT1DM dataset.
-Figs.
-12 and 13 show density histogram plots of nonzero basal and bolus insulin doses for 12 simulated data patients and 12 OhioT1DM patients, respectively.
-Table 4 shows sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set.
-Table 4: Sample count (missing samples) for the original OhioT1DM datasets and preprocessed test set
+Table 4: Sample count (% missing samples) for the original OhioT1DM datasets and preprocessed test set
 ID Original Preprocessed
@@
 (5.42%) (4.69%) (4.91%)
-575 13103 2719 269
+575 13103 2719 2691
 (9.44%) (4.74%) (4.79%)
@@
 A.2 Models
-The baseline models used in this work have been extensively used in previous blood glucose forecasting work.
-For instance, RNNs/LSTMs have been employed both as models and baselines in numerous studies (Qian et al., 2021; Li et al., 2019a; Fox et al., 2018; Rabby et al., 2021; Rubin-Falcone et al., 2020, 2022; Miller et al., 2020).
-Similarly, CNNs (Xie and Wang, 2020; Li et al., 2019a, b) and MLPs (McShinsky and Marshall, 2020; Miller et al., 2020) have been used in proposed models and as baselines.
-Exponential Smoothing (AutoETS) The AutoETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past. (Hyndman, 2008).
-DLinear (DLinear) A Linear model that leverages linear layers to model the trend and seasonal components.
+Exponential Smoothing (AutoETS) - The AutoETS model applies weighted average of past observations and exponentially decreases weights for observations further into the past (Hyndman, 2008).
+DLinear (DLinear) - A linear model that leverages linear layers to model the trend and seasonal components.
 A moving average filter is used to separate the time series into its trend and seasonal components (Zeng et al., 2023).
-Multi Layer Perceptrons (MLP) A neural network architecture composed of stacked Fully Connected Neural Networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
-Neural Hierarchical Interpolation for Time Series (NHITS) A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands.
+Multi Layer Perceptrons (MLP) - A neural network architecture composed of stacked fully connected neural networks trained with backpropagation (Nair and Hinton, 2010; Fukushima, 1975; Rosenblatt, 1958).
+Neural Hierarchical Interpolation for Time Series (NHITS) - A deep learning model that applies multi-rate input pooling, hierarchical interpolation, and backcast residual connections together to generate additive predictions with different signal bands.
 The hierarchical interpolation technique promotes efficient approximations of arbitrarily long horizons and reduced computation (Challu et al., 2022).
-Neural Basis Expansion Analysis with Exogenous (NBEATSx) An MLP-based deep learning model that decomposes the input signal into trend and seasonality components using polynomial and harmonic basis projections.
+Neural Basis Expansion Analysis with Exogenous (NBEATSx) - An MLP-based deep learning model that decomposes the input signal into trend and seasonality components using polynomial and harmonic basis projections.
 The NBEATSx extension enables modeling of exogenous variables in addition to the univariate signal (Oreshkin et al., 2020; Olivares et al., 2022a).
-Recurrent Neural Network (RNN) A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states.
+Recurrent Neural Network (RNN) - A recurrent neural network (RNN) architecture that applies recurrent transformations to obtain hidden states.
 The hidden states are transformed into contexts which are used as inputs to MLPs to generate forecasts (Elman, 1990; Cho et al., 2014).
-Long Short-Term Memory Recurrent Neural Network (LSTM) A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts.(Sak et al., 2014)
-Temporal Convolution Network (TCN) A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts.
+Long Short-Term Memory Recurrent Neural Network (LSTM) - A recurrent neural network (RNN) architecture that transforms hidden states from a multilayer LSTM encoder into contexts which are used as inputs to MLPs to generate forecasts (Sak et al., 2014).
+Temporal Convolution Network (TCN) - A 1D causal-convolutional network architecture that transforms hidden states into contexts which are used as inputs to MLP decoders to generate forecasts.
 To generate contexts, the prediction at time t is convolved only with elements from time t and earlier in the previous layer in what is referred to as causal convolutions (Bai et al., 2018; van den Oord et al., 2016).
-Informer (Informer) (Zhou et al., 2021) A Transformer-based deep learning model with a multi-head attention mechanism that uses autoregressive features from a convolution network.
-A ProbSparse self-attention mechanism and a self-attention distilling process are leveraged to reduce computational complexity.
-Patch time series Transformer (PatchTST) (Nie et al., 2023) A Transformer-based deep learning model with a multi-head attention mechanism.
-Input time series into sub-series level patches as input tokens.
-Temporal Fusion Transformer (TFT) An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies.
+Informer (Informer) - A Transformer-based deep learning model with a multi-head attention mechanism that uses autoregressive features from a convolution network.
+A ProbSparse self-attention mechanism and a self-attention distilling process are leveraged to reduce computational complexity (Zhou et al., 2021).
+Patch time series Transformer (PatchTST) - A Transformer-based deep learning model with a multi-head attention mechanism.
+The input time series is segmented into fixed-length sub-series-level patches, which serve as input tokens (Nie et al., 2023).
+Temporal Fusion Transformer (TFT) - An attention-based deep learning architecture that learns temporal relationships at different scales using LSTMs for local processing and self-attention layers to model long-term dependencies.
 TFT also leverages variable selection networks as well as a series of gating layers to suppress unnecessary parts of the architecture (Lim et al., 2021).
@@
 The deep learning models were trained using an NVIDIA A100 Tensor Core GPU.
-For all baselines and our approach, we tune the learning rate, number of training iterations, batch size, dropout rate, and random seed, with values shown in table 5.
-Additionally, for all models with the Pharmacokinetic encoder, we tune the initial $k_{a}$ values as shown in the second section of table 5.
+For all models, we use the hyperparameter values shown in Table 5 and tune the learning rate and random seed, in addition to architecture-specific hyperparameters discussed below.
+Additionally, for all models with the PK encoder, we tune the initial k values as shown in the second section of table 5.
 We tuned the hyperparameters of our model using ‘HyperOptSearch’ with the Tree-structured Parzen Estimator (TPE) algorithm and 20 samples.
@@
 Table 5: Common hyperparameter search space
-We tune additional specific hyperparameters for each architecture.
-For the MLP model, we tune the number of hidden units per layer ([128, 256]).
-For the LSTM model, we tune the hidden size ([128, 256]).
-For the TCN model, we tune the hidden size ([128, 256]) and number of dilations ([2, 4, 8]).
+In addition to tuning the learning rate and random seed for each model, we also tune architecture-specific hyperparameters.
+For the DLinear model, we tune the moving average window size ([5, 15, 25, 35]).
+For the MLP model, we tune the number of hidden units per layer ([128, 256, 512]).
+For the LSTM and RNN models, we tune the hidden size ([128, 256, 512]).
+For the TCN model, we tune the hidden size ([128, 256, 512]) and number of dilations ([2, 4, 8]).
 For the PatchTST model, we tune the hidden size ([128, 256, 512]), number of attention heads ([4, 8, 16]), number of encoder layers ([3, 4]), and patch length ([8, 16, 32, 64, 96, 128]).
@@
 For the NBEATSx and NHITS models, we did not tune additional hyperparameters.
-The replicable experiments with the full set of hyperparameters for each architecture are included in the paper’s code.
-A.4 Propositions
+The code to train models and tune hyperparameters for each architecture is included in the paper’s code repository.
+A.4 Model Evaluation
+Metrics.
+We evaluate forecasts for each test set in the two datasets using two metrics: mean absolute error (MAE) and root mean squared error (RMSE),
+$\mathrm{MAE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\frac{1}{H}\sum_{{\tau}=t+1}^{t+H}|\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau}|\quad\text{and}$ (9)
+$\mathrm{RMSE}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $=\sqrt{\frac{1}{H}\sum_{\tau=t+1}^{t+H}(\textbf{y}_{\tau}-\hat{\textbf{y}}_{\tau})^{2}}.$ (10)
+Here, $\mathbf{y}_{\tau}$ are the blood glucose values, and $\hat{\mathbf{y}}_{\tau}$ are the model predictions for the forecast horizon $H$, with $\tau=\{T+1,\dots,T+H\}$.
+We compute the metrics across all points within the forecast horizon and present the average results over 8 trials of independently trained models.
+Statistical Significance Tests.
+To evaluate the performance differences between our hybrid global-local PK model and the baseline models, we conducted statistical significance testing on the forecasting MAE (Mean Absolute Error) results.
+For each statistical comparison, patient results were averaged over eight trials of individually trained models, and the paired t-test was computed by pairing patient results between the hybrid global-local PK model and the baseline model.
+The null hypothesis for each test was that there is no difference in performance between our hybrid global-local PK model and the compared baseline across patients.
+We considered results statistically significant at the conventional threshold of $\text{p-value}<0.05$, indicating strong evidence against the null hypothesis.
+A.5 Propositions
 Proposition A.1.
@@
 Given some mathematical model, $C(t,\cdot)_{x}$, of concentration for a drug dose, $x$, that is a function of time, $t$, and pharmacokinetic parameters, $\cdot$, we show that the overall effect (concentration) of $n$ drug doses of the same drug is equal to the sum of the individual effects under linear pharmacokinetic assumptions:
-[…]
-[…]
-[…]
-[…]
+$x$ $\propto\int_{0}^{T}C(t,\cdot)_{x}dt$ (15)
+$n\cdot x$ $=n\int_{0}^{T}C(t,\cdot)_{x}dt$ (16)
+$=\int_{0}^{T}nC(t,\cdot)_{x}dt$ (17)
+$=\int_{0}^{T}\underbrace{C(t,\cdot)_{x}+C(t,\cdot)_{x}+\cdots+C(t,\cdot)_{x}}_{n\text{ times}}dt$ (18)
 $=\int_{0}^{T}\sum_{i=1}^{n}C(t,\cdot)_{x}dt$ (19)
 Here, $T$ represents the total duration of drug exposure or the time interval over which drug concentrations are being measured and $t$ represents a specific point in time within that interval.
-A.5 Algorithm
-We outline the algorithm for training our proposed PK encoder that transforms sparse exogenous variables into time-dependent concentration profiles.
-This framework can be implemented in any deep learning architecture.
-It can also be adapted to utilize other concentration-time profile functions pertinent to a specific healthcare context.
-The algorithm initializes k as a network embeding weight using a value $\textbf{k}^{*}$ from the specified range $[0,1]$ at the beginning of each epoch.
-This $\textbf{k}^{*}$ is a hyperparameter that is tuned using the Hyperopt package from the RayTune library (Liaw et al., 2018).
-For each training step, the algorithm iterates over batches of data.
-It computes a context vector c based on the input data x, k, and other parameters.
-It then updates the hyperparameters k for each subject using stochastic gradient descent (SGD) based on the loss function $\mathcal{L}$ between the predicted $\hat{\textbf{y}}$ and actual y values.
-After a defined number of training steps (determined by a validation step parameter), the algorithm performs a validation step.
-It calculates the total loss $\ell(\textbf{k})_{\text{total}}$ over the validation set by summing the loss $\ell(\textbf{k})$ and the current batch loss $\mathcal{L}(\textbf{y},\hat{\textbf{y}})$.
-It then updates the hyperparameter vector $\textbf{k}^{*}$ to minimize the total loss.
-Finally, the algorithm returns the final k values.
-Algorithm 1 PK Encoder Training.
-$E$ is the number of trainings steps, $\eta$ is the learning rate, $N$ is the number of subjects, $H$ is the forecast horizon, and $L$ is the lag time.
-Data: $N>0,f\in\mathbb{R}$, $L\in\mathbb{R}$, $H\in\mathbb{R}$
-Result: k
-$\mathcal{B}_{train}\leftarrow$ (split training data into batches)
-$\mathcal{B}_{val}\leftarrow$ (split validation data into batches)
-$\textbf{k}^{*}\leftarrow$ (Randomly select $\textbf{k}^{*}$ from defined hyperparmeter options, $\textbf{k}^{*}\in\{1,\ldots,2\}$)
+A.6 Algorithm
+We present the training procedure for the hybrid global-local PK model in Algorithm 1.
+This framework is compatible with any deep learning architecture capable of modeling exogenous variables alongside the endogenous target forecast signal.
+The algorithm initializes k as a vector of network embedding weights with values between 1 and 2 before model training.
+The initial value of k is treated as a hyperparameter and tuned using the Hyperopt package from the RayTune library (Liaw et al., 2018).
+At each training step, the algorithm iterates over batches of data.
+The model $f_{\theta}$ generates forecasts $\hat{\textbf{y}}_{\tau}$ based on the input data.
+Network parameters, including k, are updated using stochastic gradient descent (SGD) based on the loss of the model $\ell(f_{\theta})$, computed between the predicted values $\hat{\textbf{y}}$ and the true values y using the loss function $\mathcal{L}$.
+After a pre-specified number of training steps, the algorithm performs a validation step by calculating the total validation loss $\ell(f_{\theta})$.
+The final hyperparameters are selected based on the configuration that minimizes validation loss across all runs and are used to train the final model.
+The algorithm returns the final trained model $f_{\theta}$.
+Algorithm 1 Hybrid Global-Local PK Model Training.
+$E$ is the number of training steps, $\eta$ is the learning rate, $N$ is the number of subjects, $\nu$ is the sampling interval of the data, $L$ is the lag time, and $H$ is the forecast horizon.
+The model $f_{\theta}$ hyperparameters include the learning rate $\eta$, the hidden size h, and the initial (init) k values for bolus, basal, and CHO concentration curves, among other hyperparameters.
+For brevity, we define $\textbf{k}_{\text{init}}=(\textbf{k}_{\text{init, bolus}},\textbf{k}_{\text{init, basal}},\textbf{k}_{\text{init, CHO}})$ to represent the initial concentration curve parameters.
+$\mathcal{L}$ is the loss function (Huber Loss).
+$V$ is the specified validation step, or the number of training steps between each validation loss check.
+Forecasts $\hat{\textbf{y}}$ are generated by $f_{\theta}$ as a function of historical glucose values y, treatment concentration curves $\widecheck{\textbf{x}}$, and static features s.
+Input: $N>0$, $\nu>0$, $L>0$, $H>0$, $\eta>0$, $\textbf{h}\in\{128,256,512,1024\}$, $\textbf{k}_{\text{init}}\in[1,2]^{N\times 1}$
+Output: $f_{\theta}$
+$\mathcal{B}_{\text{train}}\leftarrow$ (split training data into batches)
+$\mathcal{B}_{\text{val}}\leftarrow$ (split validation data into batches)
+for each run $o$ in hyperparameter tuning samples (runs) $O$ do
+    $f_{\theta}\leftarrow$ (initialize the model)
+$\textbf{k}\leftarrow\textbf{k}_{\text{init}}$ $\triangleright$ Each entry in k is initialized to the same scalar value
+$\ell(f_{\theta})_{\text{val}}^{(o)}\leftarrow 0$ $\triangleright$ Initialize validation loss for run $o$
 for each step $i$ from $1$ to $E$ do
-       $\textbf{k}\leftarrow\textbf{k}^{*}$triangleright Initialize k
-for batch $b\in\mathcal{B}_{train}$ do
-               $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{x}_{t},\textbf{k})$
-$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
-$\ell(k)\leftarrow\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
-$\textbf{k}\leftarrow\textbf{k}-\eta\cdot\nabla_{\textbf{k}}\ell(\textbf{k})$triangleright Update k using SGD
+    for batch $b\in\mathcal{B}_{\text{train}}$ do
+          $\widecheck{\textbf{x}}\leftarrow\sum_{t=0}^{L}C(\nu\textbf{w}_{t,:},\textbf{x}_{t},\textbf{k})$
+$\hat{\textbf{y}}_{\tau}\leftarrow f_{\theta}(\textbf{y},\widecheck{\textbf{x}},\textbf{s})$
+$\ell_{\text{batch}}(f_{\theta})\leftarrow\mathcal{L}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$ $\triangleright$ $\tau$ represents the forecast horizon
+$\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\ell_{\text{batch}}(f_{\theta})$ $\triangleright$ update network parameters using SGD
       end for
-      if i % validation step == 0 then
-               $\ell(\textbf{k})\leftarrow 0$
-for batch $b\in\mathcal{B}_{val}$ do
-                     $\textbf{x}_{t-L:t}\leftarrow\sum_{t=0}^{L}C(\textbf{w}_{t,j}\cdot f,\textbf{d}_{t},\textbf{k})$
-$\hat{\textbf{y}}_{t+1:t+H}\leftarrow f_{\theta}(\textbf{y}_{t-L:t},\textbf{x}_{t-L:t})$
-$\ell(\textbf{k})_{total}\leftarrow\ell(\textbf{k})+\mathcal{L}(\textbf{y}_{t+1:t+H},\hat{\textbf{y}}_{t+1:t+H})$
+    if $i\mod V==0$ then
+          $\ell(f_{\theta})_{\text{total}}\leftarrow 0$
+for batch $b\in\mathcal{B}_{\text{val}}$ do
+             $\widecheck{\textbf{x}}\leftarrow\sum_{t=0}^{L}C(\nu\textbf{w}_{t,:},\textbf{x}_{t},\textbf{k})$
+$\hat{\textbf{y}}_{\tau}\leftarrow f_{\theta}(\textbf{y},\widecheck{\textbf{x}},\textbf{s})$
+$\ell_{\text{batch}}(f_{\theta})\leftarrow\mathcal{L}(\textbf{y}_{\tau},\hat{\textbf{y}}_{\tau})$
+$\ell(f_{\theta})_{\text{total}}\leftarrow\ell(f_{\theta})_{\text{total}}+\ell_{\text{batch}}(f_{\theta})$
          end for
-            $\textbf{k}^{*}\leftarrow\arg\min_{\textbf{k}^{*}\in\{1,\ldots,2\}}\ell(\textbf{k})_{total}$triangleright Select the hyperparameter that minimizes loss
+       $\ell(f_{\theta})_{\text{val}}^{(o)}\leftarrow\frac{\ell(f_{\theta})_{\text{total}}}{|\mathcal{B}_{\text{val}}|}$ $\triangleright$ Average loss over entire validation set and update loss
       end if
    end for
-return k
-Figure 10: Time series for the first 10 days of 6 subjects from the simulated dataset.
-Figure 11: Time series for the first 10 days of 6 subjects from the OhioT1DM dataset.
-Figure 12: Density plots of bolus and basal insulin for 12 simulated data subjects. Bolus doses are general larger than basal doses.
-Figure 13: Density plots of bolus and basal insulin for OhioT1DM subjects. Bolus doses are general larger than basal doses.
+end for
+$\{\eta,\textbf{h},\textbf{k}_{\text{init}},\ldots\}\leftarrow\arg\min_{o\in O}\ell(f_{\theta})_{\text{val}}$ $\triangleright$ select best hyperparameters; retrain final model
+return $f_{\theta}$
+Figure 11: Time series for the first 10 days of 6 patients from the simulated dataset.
+Figure 12: Time series for the first 10 days of 6 patients from the OhioT1DM dataset.
+Figure 13: Density plots of bolus and basal insulin for 12 simulated data patients. Bolus doses are general larger than basal doses.
+Figure 14: Density plots of bolus and basal insulin for OhioT1DM patients. Bolus doses are general larger than basal doses.
 Appendix B Pharmacokinetic Background
@@
 We showcase the effectiveness of global models, with parameters shared across subjects, in generating more accuracy forecasts compared to local models trained for individual subjects.
-Global models consistently outperform local models, on average across 8 trials, for each OhioT1DM subject as shown in Table.
-6 and Fig.
+Global PK models consistently outperform local PK models, on average across 8 trials, for each OhioT1DM patient as shown in Table 6 and Fig.
 6.
-Furthermore, the global NHITS-PK model consistently outperforms the global NHITS-Exogenous model, on average, for OhioT1DM subjects as shown in Fig.
-5.
-Table 6: MAE computed across forecast horizon values using local and global models for the OhioT1DM dataset. Best results are in bold.
+Furthermore, the global NHITS-PK model outperforms the global NHITSsparse exogenous model for 11 out of 12 OhioT1DM patients, on average across 8 trials, as shown in Table 6.
+Table 6: MAE computed across forecast horizon values using local and global models for the OhioT1DM dataset. The best results are in bold.
 ID w/ Exog. (Sparse) w/ Exog. (PK)
@@
 C.2 Forecasting with Dose-Independent Parameters
-Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180).
+Table 7: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and across only critical values in the forecast horizon (blood glucose $\leq 70|\geq 180$).
 Result for models with (w/) and without (w/o) exogenous variables are shown.
-Best results are in bold.
+The best results are in bold.
 Simulated Dataset
 Model All Values Critical Values
-w/o Exog. w/ Exog. w/o Exog. w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
 NHITS 8.268 7.304 11.114 9.094
@@
 […]
-We found that learning a single k value for both bolus and basal insulin resulted in improved performance on the simulated dataset for forecasts evaluated at critical values.
-However, this approach led to worse performance on the OhioT1DM dataset.
-The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static insulin-glucose kinetics parameters (Xie, 2018; Visentin et al., 2016).
-In contrast, insulin may exhibit dose-dependent or nonlinear kinetics in real-world scenarios, leading to changes in PK parameters with dose, as mentioned in section 3.3.
-Our results demonstrate a future potential application for our PK encoder to be applied for the purpose of assessing whether linear or nonlinear kinetics are present in data.
-These results also highlight the limitations of current simulated data to assess beneficial modeling choices for real-world data.
-C.3 Model MAE and RMSE Results
-This appendix presents the complete results for all models for both simulated data and OhioT1DM.
-Standard deviation across 8 trials are provided in parenthesis.
-Tables 8 and 9 provide complete MAE and RMSE results, respectively.
-Table 8: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated dataset and OhioT1DM dataset.
+We found that learning a single k value for both bolus and basal insulin improved performance on the simulated dataset only when forecasts were evaluated at critical values.
+However, this approach resulted in worse performance when evaluated across all values in the simulated dataset, as well as on the OhioT1DM dataset.
+The improved performance on the simulated dataset may be due to the simulated data generator relying on linear PK by treating bolus and basal doses in aggregate and using static parameters for insulin-glucose kinetics (Xie, 2018; Visentin et al., 2016).
+In contrast, insulin may exhibit nonlinear dose-dependent PK in real-world scenarios, as mentioned in section 3.3.
+Table 8: Comparing lag time $L$ effect on model performance in terms of mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose $\leq 70|\geq 180$) for simulated dataset and OhioT1DM dataset.
+Average results are computed across 8 trials.
+Best results are highlighted in bold.
+Simulated Dataset OhioT1DM Dataset
+Lag Time ($L$) Model All Values Critical Values All Values Critical Values
+w/ Exog.
+w/ Exog. (PK)
+w/ Exog.
+w/ Exog. (PK)
+w/ Exog.
+w/ Exog. (PK)
+w/ Exog.
+w/ Exog. (PK)
+$L=30$ timesteps TFT 5.836 6.064 7.318 7.756 9.136 9.145 10.271 10.464
+(0.248) (0.466) (0.364) (0.847) (0.363) (0.480) (0.260) (0.790)
+(2.5 hours) NBEATSx 7.619 7.506 9.230 8.915 9.263 8.789 10.583 9.841
+(0.092) (0.180) (0.126) (0.241) (0.089) (0.146) (0.148) (0.167)
+NHITS 6.807 6.566 8.454 7.784 8.356 8.231 9.585 9.393
+(0.050) (0.076) (0.212) (0.250) (0.020) (0.031) (0.080) (0.057)
+$L=60$ timesteps TFT 6.087 5.686 7.344 6.748 9.012 9.236 10.090 10.207
+(0.466) (0.073) (0.461) (0.177) (0.195) (0.544) (0.151) (0.485)
+(5 hours) NBEATSx 7.241 7.200 8.511 8.395 9.810 9.572 10.995 10.621
+(0.052) (0.103) (0.064) (0.117) (0.122) (0.710) (0.145) (0.689)
+NHITS 6.934 6.662 8.543 7.931 8.528 8.462 9.731 9.691
+(0.050) (0.092) (0.151) (0.216) (0.046) (0.059) (0.089) (0.079)
+$L=120$ timesteps TFT 5.826 5.743 7.132 7.042 9.197 9.529 10.643 10.461
+(0.192) (0.159) (0.243) (0.270) (0.455) (0.597) (0.792) (0.498)
+(10 hours) NBEATSx 6.886 6.806 8.145 7.998 10.280 9.983 11.359 10.927
+(0.031) (0.060) (0.072) (0.076) (0.214) (0.264) (0.188) (0.271)
+NHITS 7.304 7.037 9.094 8.492 8.873 8.758 10.137 9.965
+(0.055) (0.093) (0.236) (0.220) (0.064) (0.080) (0.108) (0.095)
+$L=180$ timesteps TFT 5.967 5.829 7.243 7.048 8.926 8.852 10.314 10.165
+(0.749) (0.122) (0.884) (0.257) (0.229) (0.195) (0.593) (0.400)
+(15 hours) NBEATSx 6.803 6.731 7.900 7.854 10.356 10.543 11.493 11.630
+(0.035) (0.045) (0.061) (0.073) (0.307) (0.363) (0.366) (0.488)
+NHITS 7.446 7.289 9.283 8.867 8.981 9.087 10.130 10.249
+(0.046) (0.040) (0.238) (0.122) (0.058) (0.131) (0.043) (0.128)
+(a) Simulated Dataset
+(b) OhioT1DM Dataset
+(c) Simulated Dataset
+(d) OhioT1DM Dataset
+Figure 16: (a, b) True positive rate (TPR) and (c, d) false positive rate (FPR) for the simulated and OhioT1DM datasets at various glucose thresholds for hyperglyemia event warnings (blood glucose $\geq 180$).
+TPR represents the proportion of forecast horizon windows in which the model correctly predicts an event when one actually occurs, while FPR represents the proportion of forecast horizon windows in which the model predicts an event when none occurs.
+The best performing TFT-PK model for the simulated dataset predicts key glucose thresholds in forecast windows at an approximate TPR above of 0.96 and an FPR of 0.09.
+The best performing NHITS-PK model for the OhioT1DM dataset predicts key glucose thresholds in forecast windows at an approximate TPR of 0.89 and an FPR of 0.02.
+The TPR increases as blood glucose thresholds decrease toward 160, with lower thresholds potentially serving as preliminary warning triggers to prompt patients to monitor their blood glucose and stay alert for further triggers at higher glucose levels.
+(a) Simulated Dataset
+(b) OhioT1DM Dataset
+(c) Simulated Dataset
+(d) OhioT1DM Dataset
+Figure 17: (a, b) True positive rate (TPR) and (c, d) false positive rate (FPR) for the simulated and OhioT1DM datasets at various glucose thresholds for hypoglyemia event warnings (blood glucose $\leq 70$).
+TPR represents the proportion of forecast horizon windows in which the model correctly predicts an event when one actually occurs, while FPR represents the proportion of forecast horizon windows in which the model predicts an event when none occurs.
+The best performing TFT-PK model for the simulated dataset predicts key glucose thresholds in forecast windows at an approximate TPR of 0.93 and an FPR of 0.11.
+The best performing NHITS-PK model for the OhioT1DM dataset predicts key glucose thresholds in forecast windows at an approximate TPR of 0.69 and an FPR of 0.01.
+The TPR increases as blood glucose thresholds increase toward 90, with higher thresholds potentially serving as preliminary warning triggers to prompt patients to monitor their blood glucose and stay alert for further triggers at lower glucose levels.
+C.3 Lag Time Ablation
+We performed an ablation study on model performance for various lag times $L$, including 2.5 hours (30 time steps), 5 hours (60 time steps), and 10 hours (120 time steps; results presented in the main paper), as well as 15 hours (180 time steps), for the top three performing deep learning models: TFT, NBEATSx, and NHITS.
+The study was conducted at a fixed forecast horizon $H$ of 30 minutes (6 time steps).
+For each lag time, the NBEATSx-PK and NHITS-PK models had a lower MAE than their sparse exogenous model counterparts, except for L=180L\!=\!180 in the OhioT1DM dataset, as shown in Table 8.
+The TFT-PK model had a lower MAE than its sparse exogenous model counterpart for $L\in\{60,120,180\}$ in the simulated dataset and for L=180L\!=\!180 in the OhioT1DM dataset.
+For each lag time, we evaluate whether the PK models have a significantly lower forecasting error, in terms of MAE, than their sparse exogenous model counterparts using paired t-tests, as outlined in Section A.4.
+For the simulated dataset, the NBEATSx-PK and NHITS-PK models showed a statistically significant performance improvement over their sparse exogenous model counterparts ($\text{p-value}<0.05$) for all lag times, while the TFT-PK model showed a statistically significant performance improvement for all lag times except $L=2.5$.
+For the OhioT1DM dataset, the NBEATSx-PK and NHITS-PK models demonstrated a statistically significant performance improvement over their sparse exogenous model counterparts ($\text{p-value}<0.05$) for lag times of 2.5, 5, and 10 hours, while the TFT-PK model showed a statistically significant performance improvement for L=180L\!=\!180.
+Notably, for L=180L\!=\!180, the TFT-PK model was also the best-performing TFT model variant across all lag times for the OhioT1DM dataset.
+C.4 Critical Event Prediction Rate Analysis
+We evaluate the ability of models to predict critical events using two metrics: the true positive rate (TPR), or the proportion of forecast horizon windows in which the model correctly predicts an event when one actually occurs, and the false positive rate (FPR), or the proportion of forecast horizon windows in which the model predicts an event when none occurs.
+From the TPR, one can also obtain the false negative rate ($\text{FNR}=1-\text{TPR}$), or the proportion of forecast horizon windows in which the model does not predict an event when one actually occurs.
+We measure each metric on multiple blood glucose thresholds up to critical event thresholds (blood glucose $\leq 70|\geq 180$).
+The TPR and FPR for the simulated and OhioT1DM datasets at various glucose thresholds for hyperglycemia event warnings are shown in Fig.
+16.
+The best performing TFT-PK model for the simulated dataset predicts hyperglycemic events (blood glucose $\geq 180$) in forecast windows at an approximate TPR of 0.96 and an FPR of 0.09.
+The best performing NHITS-PK model for the OhioT1DM dataset predicts hyperglycemic events in forecast windows at an approximate TPR of 0.89 and an FPR of 0.02.
+The TPR increases as blood glucose thresholds decrease toward 160 mg/dL, with lower thresholds potentially serving as preliminary warnings, prompting patients to monitor their blood glucose and stay alert for further warnings at higher glucose levels.
+The TPR and FPR for the simulated and OhioT1DM datasets at various glucose thresholds for hypoglycemia event warnings are shown in Fig.
+17.
+The best performing TFT-PK model for the simulated dataset predicts hypoglycemic events (blood glucose $\leq 70$) in forecast windows at an approximate TPR of 0.93 and an FPR of 0.11.
+The best performing NHITS-PK model for the OhioT1DM dataset predicts hypoglycemic events in forecast windows at an approximate TPR of 0.69 and an FPR of 0.01.
+The TPR increases as blood glucose thresholds decrease toward 90 mg/dL, with higher thresholds potentially serving as preliminary warnings to prompt patients to monitor their blood glucose and stay alert for further warnings at lower glucose levels.
+These results highlight the effectiveness of the models in predicting critical glucose events, demonstrating high TPRs while maintaining relatively low FPRs.
+The ability to provide early warnings at varying glucose thresholds can be valuable for proactive patient monitoring and timely intervention.
+However, there is still room for improvement, as achieving higher TPRs with even lower FPRs could further enhance reliability and reduce false alarms.
+C.5 Computational Efficiency Analysis
+Model FLOPS # Trainable Parameters Inference Time (ms)
+TFT w/o Exog. 3.66e8 2.53 million 15.9
+TFT w/ Exog. 6.11e8 ($\uparrow 66.94\%$) 2.78 million ($\uparrow 9.88\%$) 17.7
+TFT-PK w/ Exog. 6.07e8 ($\uparrow 65.85\%$) 2.78 million ($\uparrow 9.88\%$) 21.3
+NBEATSx w/o Exog. 8.01e7 10.02 million 5.2
+NBEATSx w/ Exog. 8.89e7 ($\uparrow 10.99\%$) 11.13 million ($\uparrow 11.08\%$) 5.8
+NBEATSx-PK w/ Exog. 8.30e7 ($\uparrow 3.62\%$) 10.39 million ($\uparrow 3.69\%$) 5.7
+NHITS w/o Exog. 8.19e7 10.25 million 5.5
+NHITS w/ Exog. 9.08e7 ($\uparrow 10.87\%$) 11.36 million ($\uparrow 10.83\%$) 5.8
+NHITS-PK w/ Exog. 8.49e7 ($\uparrow 3.66\%$) 10.62 million ($\uparrow 3.61\%$) 8.4
+Table 9: Computational efficiency characteristics including FLOPS (floating-point operations per second), number of trainable parameters, and inference time measured for the OhioT1DM dataset models.
+The percentage increase in FLOPs and the number of trainable parameters over the baseline model without (w/o) exogenous (exog.) variables is shown in parentheses (e.g., $\uparrow\text{x.xx}\%$).
+As expected, models with (w/) exogenous variables exhibit larger FLOPs, more trainable parameters, and longer inference times due to the processing of higher-dimensional input.
+However, the PK models have the same or fewer FLOPs and trainable parameters compared to their sparse exogenous model counterparts, as the exogenous features are leveraged more effectively in the PK model architectures.
+To assess the computational efficiency characteristics of the deep learning models, we calculate three key metrics: FLOPs (floating-point operations per second), the number of trainable parameters, and the inference time for the OhioT1DM dataset models.
+These metrics provide insights into the computational complexity, memory usage, and real-time efficiency of the model.
+We calculate FLOPs using the FlopCounterMode class from Pytorch, which tracks operations during model execution.
+The resulting FLOPs represent the computational load of the model for each forward pass.
+The number of trainable parameters in a model reflects its size.
+We calculate this by iterating through all the model’s parameters and summing the number of elements for those that require gradients.
+Inference time refers to the time it takes for the model to generate predictions from input data.
+To measure this, we execute the model multiple times (100 runs) and record the time taken for each forward pass.
+The inference time is then calculated as the average time taken across these runs, providing an estimate of the model’s efficiency during real-time deployment.
+This metric is essential for applications like clinical decision support or real-time monitoring devices, where timely predictions are crucial.
+The metrics for the three best performing deep learning models, TFT-PK, NBEATSx-PK, and NHITS-PK are presented in Table 9.
+In addition to providing metrics for the PK models, the table also includes metrics for models with (w/) and without (w/o) exogenous (exog.) features for reference.
+As expected, models with exogenous variables exhibit larger FLOPs, more trainable parameters, and longer inference times due to the processing of higher-dimensional input.
+However, the PK models have the same or fewer FLOPs and trainable parameters compared to their sparse exogenous model counterparts, as the exogenous features are leveraged more effectively in the PK model architectures.
+For NBEATSx and NHITS models, this efficiency stems from our proposed approach, where exogenous variables are processed in only one of the deep stacks of fully connected layers rather than across all stacks.
+This design improves processing efficiency and reduces model size.
+Without the PK encoder, the approach of isolating sparse exogenous variables to a single stack generally results in worse performance, as shown in Table 10.
+The TFT-PK, NBEATSx-PK, and NHITS-PK models with sparse exogenous features show increases in FLOP of 65.30%, 3.62%, and 3.66%, respectively, compared to the baseline models without exogenous variables.
+In contrast, the models with sparse exogenous features exhibit higher FLOP increases of 66.94%, 10.99%, and 10.87%, respectively, compared to the baseline models without exogenous variables.
+This result highlights the utility of the PK models, as they not only provide lower forecasting errors but also offer greater computational efficiency.
+Additionally, the models have fast inference speeds, on the order of magnitudes smaller than 1 second.
+Table 10: Processing covariates in 1 versus 3 of the deep stacks of fully connected layers for NBEATSx and NHITS.
+Without the PK encoder, the approach of isolating sparse exogenous variables to a single stack generally results in worse performance.
+Simulated Dataset OhioT1DM Dataset
+NBEATSx NHITS NBEATSx NHITS
+1 Stack 6.908 7.429 10.020 8.897
+(0.101) (0.042) (0.569) (0.246)
+3 Stacks 6.886 7.304 10.280 8.873
+(0.031) (0.055) (0.214) (0.064)
+C.6 Feasibility of Model Integration into Clinical Systems
+Integrating deep learning forecasting models into clinical decision-support systems or real-time monitoring devices can provide timely, data-driven alerts for critical events, aiding decision-making for proactive interventions.
+To ensure smooth integration into a clinical decision-support system or real-time monitoring device, deep learning models must be both data-efficient and capable of delivering reasonable inference times to facilitate updated forecasts at appropriate time intervals.
+The OhioT1DM patients use Medtronic 530G or 630G insulin pumps and Medtronic Enlite Continuous Glucose Monitor (CGM) sensors.
+The MiniMed 630G Insulin Pump is capable of storing 90 days of pump history and glucose sensor data.
+Given that such devices are equipped to store data on the scale of many days, the feasibility of integrating the proposed hybrid global-local PK models into such systems is high, as our PK models only require a maximum of 10 hours of historical data to generate forecasts.
+The PK models are also memory-efficient.
+For example, the NHITS-PK model requires less than 20MB of memory during inference on a CPU in a cold-start scenario.
+A cold start refers to the initial model inference, where no preloaded state or cached data is used, providing an upper bound on memory usage.
+Furthermore, the NHITS-PK model demonstrates an inference time of approximately 8 milliseconds when executed on a CPU, ensuring rapid predictions that support real-time operation.
+This combination of low storage requirements and fast inference times makes the integration of such models into clinical systems feasible to help monitor and provide early warnings for critical events.
+While a graphics processing unit (GPU) is recommended for efficient model training, a central processing unit (CPU) is sufficient for inference with the trained model, offering space, cost, and energy efficiency benefits over dedicated graphics processing devices.
+In addition to CGMs, deep learning models could be integrated into other mobile devices, such as a phone app.
+The model could provide continuous forecast windows that update every 5 minutes (the frequency of the data) and trigger a beep or vibration if the patient is at risk of a critical event occurring within the forecast horizon.
+The device could alert the patient with customizable blood glucose thresholds, accounting for warnings of varying severity.
+For example, preliminary warnings could be set at less severe blood glucose levels to prompt patients to monitor their blood glucose and remain alert for further triggers at more severe thresholds.
+Furthermore, integrating the model into a phone app could facilitate sharing alert information, notifying patients and allowing the inclusion of other monitoring users, such as parents, school nurses, or caregivers.
+Table 11: Mean absolute error (MAE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose $\leq 70|\geq 180$) for simulated dataset and OhioT1DM dataset.
 Average results are computed across 8 trials.
@@
 Best results are highlighted in bold, second best results are underlined.
-PK models with improved forecasts over their respective exogenous baseline mdoel are highlighted in blue.
+PK models with improved forecasts over their respective exogenous baseline model are highlighted in blue.
 Simulated Dataset OhioT1DM Dataset
 Model All Values Critical Values All Values Critical Values
-w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
 AutoETS 9.450 13.457 9.022 10.028
@@
 (0.054) (0.141) (0.126) (0.169) (1.595) (3.590) (1.228) (4.080)
-RNN 13.703 20.190 21.433 35.955 11.325 11.795 13.627 13.552
+RNN 12.728 12.930 20.308 20.945 11.535 10.907 12.641 12.989
 (8.829) (8.819) (19.533) (19.925) (3.056) (0.809) (2.342) (1.822)
-LSTM 13.557 10.321 20.537 13.550 9.811 9.467 11.252 11.178
+LSTM 9.250 9.152 12.034 12.000 10.217 10.615 12.102 12.983
 (2.041) (2.097) (1.830) (3.983) (0.930) (0.807) (1.218) (1.464)
-TCN 12.096 10.109 17.709 12.660 11.525 10.760 13.647 12.678
+TCN 14.340 11.380 23.370 14.817 10.769 11.516 12.812 13.203
 (8.271) (2.049) (20.857) (3.258) (1.093) (1.846) (0.901) (1.882)
@@
 (0.115) (0.192) (0.178) (0.243) (0.811) (0.455) (1.038) (0.792)
-MLP 10.589 8.433 14.699 10.600 11.065 11.120 12.630 12.698
+MLP 10.424 8.407 13.932 10.297 10.300 13.254 11.550 14.997
 (1.478) (0.387) (0.861) (0.329) (0.329) (0.245) (0.359) (0.353)
@@
 (0.093) (0.220) (0.080) (0.095)
-Table 9: Root Mean Square Error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose 70|180$\leq$ 70$\geq$ 180≤ 70 | ≥ 180) for simulated dataset and OhioT1DM dataset.
+Table 12: Root mean square error (RMSE) computed for model predictions across all values in the forecast horizon and only across critical values in the forecast horizon (blood glucose $\leq 70|\geq 180$) for simulated dataset and OhioT1DM dataset.
 Average results are computed across 8 trials.
@@
 Best results are highlighted in bold, second best results are underlined.
-PK models with improved forecasts over their respective exogenous baseline mdoel are highlighted in blue.
+PK models with improved forecasts over their respective exogenous baseline model are highlighted in blue.
 Simulated Dataset OhioT1DM Dataset
 Model All Values Critical Values All Values Critical Values
-w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog. w/o Exog. w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
+w/o Exog.
+w/ Exog.
 AutoETS 14.751 21.454 14.569 16.225
@@
 (0.085) (0.119) (0.340) (0.216) (1.462) (3.511) (0.970) (3.869)
-RNN 19.033 27.116 28.834 43.256 16.498 17.066 19.446 19.423
+RNN 17.679 18.017 26.312 26.849 16.433 15.762 18.274 18.375
 (10.978) (10.935) (19.591) (19.935) (2.734) (0.657) (2.199) (1.654)
-LSTM 18.567 13.723 26.863 18.028 14.915 14.597 16.894 16.851
+LSTM 12.993 12.526 17.027 16.222 15.337 15.571 17.759 18.399
 (1.896) (2.389) (1.599) (4.528) (0.750) (0.684) (0.949) (1.295)
-TCN 17.094 13.551 25.163 16.908 16.690 15.971 19.213 18.418
+TCN 20.040 15.181 30.812 19.573 16.073 16.483 18.642 18.677
 (10.281) (2.571) (20.017) (4.054) (0.855) (1.544) (0.601) (1.451)
@@
 (0.139) (0.314) (0.329) (0.446) (0.791) (0.345) (0.959) (0.684)
-MLP 15.561 11.461 22.380 14.449 16.534 16.948 18.651 19.289
+MLP 15.196 11.267 21.246 13.832 15.908 18.986 17.762 21.340
 (1.322) (0.405) (0.692) (0.336) (0.276) (0.312) (0.421) (0.440)
@@
 (0.188) (0.435) (0.092) (0.089)
+C.7 Model MAE and RMSE Results
+This appendix presents the complete results for all models for both simulated data and OhioT1DM.
+Standard deviation across 8 trials are provided in parenthesis.
+Tables 11 and 12 provide complete MAE and RMSE results, respectively.